### v1

# BED-LLM: Intelligent Information Gathering with LLMs and Bayesian Experimental Design

Abstract
We propose a general-purpose approach for improving the ability of Large Language Models (LLMs) to intelligently and adaptively gather information from a user or other external source using the framework of sequential Bayesian experimental design (BED). This enables LLMs to act as effective multi-turn conversational agents and interactively interface with external environments. Our approach, which we call BED-LLM (Bayesian Experimental Design with Large Language Models), is based on iteratively choosing questions or queries that maximize the expected information gain (EIG) about the task of interest given the responses gathered previously. We show how this EIG can be formulated in a principled way using a probabilistic model derived from the LLM’s belief distribution and provide detailed insights into key decisions in its construction. Further key to the success of BED-LLM are a number of specific innovations, such as a carefully designed estimator for the EIG, not solely relying on in-context updates for conditioning on previous responses, and a targeted strategy for proposing candidate queries. We find that BED-LLM achieves substantial gains in performance across a wide range of tests based on the 20–questions game and using the LLM to actively infer user preferences, compared to direct prompting of the LLM and other adaptive design strategies.
1 Introduction
Despite their many successes, LLMs currently fall short on a crucial aspect of interactive intelligence: proactively seeking out information from a user or external environment in an intelligent and adaptive manner (Laban et al,, 2025; Li et al, 2025c, ). For example, they have been shown to perform poorly on problems such as multi-step guessing games (Bertolazzi et al,, 2023; Zhang et al,, 2024), task clarification (Chi et al,, 2024), IT task automation (Jha et al,, 2025), and iterative external tool use (Patil et al,, 2025). In particular, while modern LLMs are often capable of producing coherent and insightful questions (or other external queries) in a one-shot manner, they typically struggle to appropriately tailor their questions to previously gathered responses on interactive tasks (Bertolazzi et al,, 2023; Patil et al,, 2025).
There is, therefore, a pressing need to improve the ability of LLMs to adaptively ask questions and gather information in a targeted manner. Such capabilities are essential for a wide variety of problems, such as clarifying user intent, personalizing model behavior to a particular user, or generally acting as effective multi-turn conversational agents. They are also critical if we want to use LLMs in data gathering tasks or as automated agents in decision-making pipelines (Wu et al,, 2025). In turn, these capabilities are essential for a wide variety of downstream applications from medical diagnostic assistants, targeted code generation, and tutoring systems, to conducting automated surveys, AI–driven scientific inquiries, and market research (Liu et al, 2024a, ; Mandal et al,, 2025; Jacobsen et al,, 2025; Lu et al,, 2024). Note that in all these problems it is not enough for the LLM to simply one-shot generate good fixed questions based only on its parametric knowledge. Instead, we need it to be able to adaptively choose questions that are good, conditioned on the already-collected user responses.
We propose to address this challenge using the framework of sequential Bayesian experimental design (BED, Lindley, (1956); MacKay, (1992); Chaloner & Verdinelli, (1995); Sebastiani & Wynn, (2000); Rainforth et al, (2024)), which provides a model-based, information-theoretic mechanism for making adaptive design decisions, given a generative model of the experiment. Specifically, we show how the problem of interactive information gathering with LLMs can be formulated as a sequential experimental design problem, wherein we iterate between choosing queries based on maximizing their expected information gain (EIG) and updating our beliefs with the information from the received response. Here, the underlying generative model from which the sequential BED procedure is formed is derived from the LLM, and we carefully show how this should be constructed, providing extensive insights on key design decisions.
We call our approach BED-LLM (which stands for Bayesian Experimental Design with Large Language Models) and show how its success is critically dependent on careful choices for how the EIG is estimated, how candidate queries are generated, and how the model is updated given previous responses. For example, we show how approximations to the EIG made in previous work (Kobalczyk et al,, 2025; Hu et al,, 2024) can significantly degrade performance compared to using more carefully constructed estimation strategies. We also show how simple in-context learning (Brown et al,, 2020) is typically not sufficient to incorporate all the information from previous query–response pairs, with an additional sample–then–filter strategy introduced to correct this. Finally, we introduce a targeted strategy for proposing candidate queries (whose EIG we then evaluate) that directly utilizes our current belief state and provides important improvements over direct generation by the LLM.
Together, we find that these innovations provide substantial performance benefits over directly generating queries from the LLM and more basic approximations of the sequential BED framework. Specifically, we first find that BED-LLM provides substantial improvements in the success rate for the 20-Questions problem across a variety of LLMs and target quantities. For example, when predicting celebrities with Mistral-Large, we observe an improvement in success rate from 14% to 91%. Second, we demonstrate noticeable improvements in using the LLM for movie recommendations, showing that these benefits hold even when the LLM’s predictive model differs from that of the answerer.
In summary, we illustrate how the adaptive information gathering capabilities of LLMs can be significantly improved by the careful application of the sequential BED framework to their decision making, thereby making an important step towards their more general usage as interactive agents.
2 Problem Formulation and Background
There are two natural ways one might look to improve LLMs’ adaptive information gathering capabilities: we can look to change the LLM itself through finetuning or altering the pretraining process, or we can make changes in how the LLM is used at deployment time. Our focus will be on the latter, as this means we can make improvements to performance without requiring access to any additional data or expensive retraining, noting that for information–gathering problems, we will not usually have task–specific data available upfront (e.g. we cannot expect our LLM to have any knowledge about a particular user’s preferences). However, we emphasize that any improvements that can be made to the LLM itself will be complementary to our deployment–time approach.
To formalize the notion of information gathering, we need a concrete idea of what we wish to learn about. We denote this target quantity as $\theta$, and note that it could, for example, be a user’s preference, the answers to one or more questions, a recommendation, or a piece of content that a user wants to generate. To meaningfully reason about $\theta$, we need to have some belief distribution for it, which we denote $p\left(\theta;h_{t}\right)$, [^1] Note we carefully distinguish between explicit conditioning, i.e. $p(a|b)$, and more general dependency, $p(a;b)$, in our notation. The former explicitly corresponds to the conditional distribution of an associated joint distribution, $p(a,b)$, while the latter may not. So here, for example, $h_{t}$ influences our distribution on $\theta$, but we are not deriving this distribution through conditioning of some explicit joint distribution. where $h_{t}$ represents the history of queries and responses at time $t$. As more queries are made, this history will grow and our beliefs on $\theta$ will update accordingly. We note that $p\left(\theta;h_{t}\right)$ will typically not be defined explicitly. Even in the LLM setting, there is significant flexibility to how it is constructed, as we discuss in Sections 3.3 and 4.
2.1 Information Gathering with LLMs
Our focus is on how to use LLMs to gather information about our target quantity $\theta$. The precise set of problems this can relate to is quite broad and includes, to name but a few, task clarification (Zhang et al, 2025a, ), preference learning (Handa et al,, 2024; Chakraborty et al,, 2024; Ouyang et al,, 2022), diagnosis and troubleshooting (Hirosawa et al,, 2024; Jha et al,, 2025), automating external simulations or real-world experiments (Mandal et al,, 2025), conducting surveys or collecting feedback (Aher et al,, 2023; Lee et al,, 2024), and AI tutoring agents (Kestin et al,, 2025; Liu et al, 2024a, ).
What all of these have in common is that the LLM starts with incomplete information about the target $\theta$, but can refine its beliefs by making queries, $x\in\mathcal{X}$, to the user or some other external agent and receiving responses, $y\in\mathcal{Y}$, that are informative about $\theta$. In general, multiple such queries, $x_{1},\dots,x_{n}$, can be made. We can choose them adaptively in a sequential decision-making process where we iteratively choose each $x_{t+1}$ based on the collected history $h_{t}:=(x_{i},y_{i})_{i=1:t}$. For clarity of exposition, we focus on the case where the $x_{i}$ correspond to explicit questions asked to the user, but emphasize that the approach applies more broadly to other forms of external interaction by the LLM, such as retrieving documents or calling external functions.
The simplest and cheapest way to incorporate the interaction history into the LLM to aid our sequential decision making is to include it in the context (Brown et al,, 2020). Namely, if the LLM’s distribution over generated text, $z\in\mathcal{Z}$, is $p_{\mathrm{LLM}}(z;c)$, where $c$ is an initial prompt, then we can consider $p_{\mathrm{LLM}}(z;[c,h_{t}])$ as an updated LLM that uses the previous question–response pairs as in-context examples, where $c,h_{t}]$ represents some appropriate concatenation of the history into the context. From this, we can derive an updated belief distribution over $\theta$. Most simply, this can be done by using $p_{\mathrm{LLM}}(z;[c,h_{t}])$ to directly query about $\theta$ (e.g. if $\theta$ is some preference, we could prompt the LLM to predict this preference). However, as we show later, this approach often fails to appropriately or fully incorporate the information from $h_{t}$, leading to belief states that are inconsistent with past observations. We will introduce a more robust and sophisticated method for deriving $p(\theta;h_{t})$ from $p_{\mathrm{LLM}}(z;[c,h_{t}])$ in Section˜3.3.
In the interest of brevity, for the rest of the paper we will drop the initial prompt $c$ from our notation and also assume that an appropriate final prompt has been added to end of the context to predict/sample our quantity of interest. For example, we will use $p_{\mathrm{LLM}}(\theta;h_{t})$ to denote the distribution when our context is $c,h_{t}]$ followed by an appropriate prompt to sample $\theta$ as its next token(s).
2.2 Information–Theoretic Experimental Design
Our aim of maximizing the information gathered about a target quantity can be formalized using the framework of Bayesian experimental design (BED; Lindley,, 1956; 1972; Chaloner & Verdinelli,, 1995; Rainforth et al,, 2024), which provides a powerful and principled framework for data acquisition.
The core of the BED framework is a joint generative model, $p(\theta,y;x)$ over the target quantity $\theta$ and outcomes, $y$, given designs $x$. Traditionally, this is most commonly specified as a conventional Bayesian model using a prior $p\left(\theta\right)$ and likelihood $p(y|\theta;x)$ (Rainforth et al,, 2024), but it can also be directly defined using a stochastic simulator (Kleinegesse & Gutmann,, 2020; Rainforth,, 2017) or as the marginal of some larger model (e.g. because $\theta$ corresponds to something other than the model parameters (Bickford Smith et al,, 2023; MacKay,, 1992)).
In the most general case, designs are then chosen to maximize the expectation of some utility function $U(\theta,y,x)$ under this generative model; that is we take $x^{*}=\mathrm{argmax}_{x}\mathbb{E}_{p(\theta,y;x)}\left[U(\theta,y,x)\right]$. The most common choice is to take $U(\theta,y,x)=\log p(\theta,y;x)-\log p\left(\theta\right)\log p(y;x)$ where $p\left(\theta\right)$ and $p(y;x)$ are the marginal distributions on $\theta$ and $y$ implied by our joint model and we have assumed that our current beliefs on $\theta$ are independent of the design. This leads to an objective corresponding to the expected information gain (EIG) in $\theta$ (Lindley,, 1956; 1972),
$\mathrm{EIG}_{\theta}(x)$ $=\mathbb{E}_{p(\theta,y;x)}\left[\log\frac{p(\theta,y;x)}{p\left(\theta\right)p(y;x)}\right],$ (1)
which using Bayes’ rule can also be represented in the following equivalent forms
$\text{EIG}_{\theta}(x)$ $=\mathrm{H}[p\left(\theta\right)]-\mathbb{E}_{p(y;x)}[\mathrm{H}[p(\theta|y;x)]]$ (2)
$=\mathrm{H}[p(y;x)]-\mathbb{E}_{p\left(\theta\right)}[\mathrm{H}[p(y|\theta;x)]],$ (3)
where $\mathrm{H}$ denotes the Shannon entropy (i.e., $\mathrm{H}[p\left(\theta\right)]=-\mathbb{E}_{p\left(\theta\right)}[\log p\left(\theta\right)]$). We can thus equivalently think of this EIG as the mutual information between $\theta$ and $y$, the expected information gain over possible data simulated from our model (where the information gain is defined as the reduction in entropy from our prior on $\theta$ to the posterior), and the expected reduction in entropy over data from observing $\theta$ simulated from our prior (Sebastiani & Wynn,, 2000). By maximizing this objective with respect to $x$, we are thus maximizing the expected amount of information we gain in $\theta$, subject to the assumptions that our generative model can both accurately estimate the true data-generating process and encapsulate our beliefs about $\theta$.
Working with the EIG is highly suited to a sequential or adaptive design approach, wherein it is generally referred to as sequential BED or Bayesian Adaptive Design (Rainforth et al,, 2024). Because the EIG is only a function of our underlying model, when we update the model as new data becomes available, our EIG design objective will naturally update as well. In particular, if $p(\theta,y_{t+1};h_{t},x_{t+1})$ represents our generative model after already observing the history $h_{t}$, then we can iteratively optimize the incremental EIG (MacKay,, 1992; Cavagnaro et al,, 2010),
$\mathrm{EIG}_{\theta}(x_{t+1};h_{t})$ $=\mathbb{E}_{p(\theta,y_{t+1};h_{t},x_{t+1})}\left[\log\frac{p(\theta,y_{t+1};h_{t},x_{t+1})}{p\left(\theta;h_{t}\right)p(y_{t+1};h_{t},x_{t+1})}\right].$ (4)
Here our updated model, $p(\theta,y_{t+1};h_{t},x_{t+1})$, conventionally comes from a Bayesian update of the original model, noting that the Savage axioms dictate that a rational agent should update its beliefs in a Bayesian manner (Savage,, 1954). However, in many cases, this is not practical and other non-Bayesian updates are performed instead. In particular, in many Bayesian active learning problems, the update actually corresponds to retraining the model with an updated dataset (Gal et al,, 2017; Bickford Smith et al,, 2023).
It has recently been noted that this traditional sequential BED approach can sometimes be suboptimal in practice, as it only optimizes the EIG of the next observation, without planning ahead for the fact that design decisions taken at a given step can also influence the achievable EIGs from future steps (Foster,, 2021). A variety of policy–based BED approaches have subsequently been proposed to address this (Foster et al,, 2021; Ivanova et al,, 2021; Blau et al,, 2022; Huan & Marzouk,, 2016; Hedman et al,, 2025), while also removing the need to make model updates and conduct optimizations during the experiment itself. However, training these policies typically requires a very large number of simulations from the generative model, which makes them impractical for our setting, due to the expense of simulating full experimental rollouts from the LLM.
3 Sequential Bayesian Experimental Design for LLMs
The sequential BED framework described in Section˜2.2 requires two core components to be specified by the user: an initial joint model $p(\theta,y;x)$ over hypotheses $\theta$ and outcomes $y$, given chosen experiment(s) $x$; and a procedure to derive an updated model $p(\theta,y_{t+1};h_{t},x_{t+1})$ after observing an experimental history $h_{t}$. In the LLM setting, there is significant flexibility in these critical design decisions. In particular, there are various ways that one can derive a suitable joint distribution from the LLM and the ability to learn in context provides opportunities for update methods that go beyond standard Bayesian model updates.
Model Construction
A major challenge in the LLM setting is that unlike conventional probabilistic models, in general, $p_{\mathrm{LLM}}(\theta)\,p_{\mathrm{LLM}}(y;[\theta,x])\neq p_{\mathrm{LLM}}(y;x)\,p_{\mathrm{LLM}}(\theta;[x,y])$. That is, we induce a different joint distribution if we first sample $\theta$ then sample $y$ with $\theta$ in context (which we refer to as the prior–likelihood pairing), than if we first sample $y$ then sample $\theta$ with $y$ in context (which we refer to as the data–estimation pairing). Moreover, in each pairing we may wish to deviate from the LLM’s directly induced distribution on one or both components. The success of using sequential BED with LLMs turns out to be critically dependent on which pairing we choose and the deviations we make from LLM’s directly induced distribution.
We delay proper discussion of this complex issue until Section˜4, where we will see that the preferable setup can depend on the precise problem setting and, in particular, the relative complexity of spaces of $\theta$ and $y$. For now, we will focus on using the prior–likelihood pairing, as this transpires to generally be the advantageous setup when the space of $\theta$ is more complex than that of $y$, which we believe to be the more commonly occurring setting and the one we consider in our experiments. While we will generally use the LLM’s directly induced distribution for the likelihood, we allow the prior to deviate from this in a problem–specific manner. As such, our initial joint model will be $p(\theta,y;x)=p(\theta)p_{\mathrm{LLM}}(y;[\theta,x])$, where the specification of the prior $p(\theta)$ can vary between problems.
Model Updating
Optimally updating the joint model in this setting requires incorporating new observations in a way that both fully captures the available information from new data and is computationally tractable. At one extreme, we could perform full Bayesian updates via approximate inference, as in classical sequential BED. This approach, however, demands a prohibitively large number of LLM evaluations to accurately approximate the posterior, and it does not exploit the power of the LLM as a probabilistic generative model, where autoregressive sequential rollouts often lead to more nuanced and diverse behavior than repeated static likelihood queries. The other obvious choice is to rely on simple in context updates, such that we use $p(\theta;h_{t})=p_{\mathrm{LLM}}(\theta;h_{t})$. However, as we show later, this fails to sufficiently capture the information provided by the new data and thus undermines the sequential BED approach.
As we will discuss in Section˜3.3, we therefore employ a strategy that is somewhere between the two, wherein we set up a rejection sampling procedure that draws samples from the in–context distribution, $p_{\mathrm{LLM}}(\theta;h_{t})$, but then filters out samples that are actually not compatible with previous observations. We refer to the resulting distribution as $p_{f}(\theta;h_{t})$.
In our experiments and subsequent exposition of our approach, we will not update the likelihood model as more data is gathered, that is, our likelihood in the sequential setting will be $p_{\mathrm{LLM}}\left(y_{t+1};\left[\theta,x_{t+1}\right]\right)$ instead of $p_{\mathrm{LLM}}(y_{t+1};[h_{t},\theta,x_{t+1}])$. The main rationale of this choice is that for many problems our beliefs on $\theta$ capture all the required information to predict $y|x$, hence including the history is adding unnecessary context that could influence the LLM’s behavior in undesirable ways. However, it is important to note that $p_{\mathrm{LLM}}(y_{t+1};[h_{t},\theta,x_{t+1}])$ should be used instead for problems where $\theta$ will not capture all information from previous data, e.g. if $\theta$ is a binary value corresponding to whether we reject a null hypothesis, or is the answer to a particular other question of interest.
BED-LLM
We now introduce our specific principled algorithm for intelligent and adaptive information gathering in LLMs using sequential BED. We call our approach BED-LLM, which stands for Bayesian Experimental Design with Large Language Models. Here, the questions posed to the user [^2] More broadly, these could be more general queries such as document retrieval or function calling; however in our experiments we only consider questions asked of a user. will correspond to our designs, $x$, and the responses received will correspond to our outcomes, $y$. Using the LLM to derive a joint model over these outcomes and the target variables $\theta$ as described above, we can derive a precise form of the expected information gain, $\text{EIG}_{\theta}(x)$. We can then interleave choosing informative questions by optimising the incremental EIG defined in Equation˜4, and updating our underlying model based on the received question–response pairs. Specifically, BED-LLM iterates over the following key steps, where $t$ indexes the current turn.
  1. 1.
    Generate Candidate Questions: Propose a candidate set of $M$ diverse, multiple-choice questions, $\mathcal{X}^{\text{cand}}$, by appropriate sampling of the LLM based on the current conversational context (Section˜3.1).
  2. 2.
    Compute EIG Estimator: For each candidate question $x\in\mathcal{X}^{\text{cand}}$, estimate $\text{EIG}_{\theta}(x)$ as described in in Section˜3.2.
  3. 3.
    Select and Ask Optimal Question: Choose the question $x_{t}\in\mathcal{X}^{\text{cand}}$ that yields the highest estimated EIG. Pose $x_{t}$ to the user, observe response $y_{t}$, and update the history, $h_{t}=(h_{t-1},(x_{t},y_{t}))$.
  4. 4.
    Construct updated joint, $p(\theta,y_{t+1};h_{t},x_{t+1})=p_{f}(\theta;h_{t})p_{\mathrm{LLM}}\left(y_{t+1};\left[\theta,x_{t+1}\right]\right)$, using the new history (Section˜3.3) and return to Step 1 (unless a termination criterion like maximum number of questions has been achieved).
Because BED-LLM uses the prior–likelihood pairing, our belief state on $\theta$ is simply given by $p_{f}(\theta;h_{t})$ at any given point in the process; our underlying aim is that the uncertainty of this distribution contracts as quickly as possible.
In the remainder of this section, we provide details on these key individual steps. We note that some of the choices made are closely linked to our joint model construction; we will later provide extensive discussion on why this model construction was chosen and possible alternatives in Section˜4.
3.1 Candidate Question Generation
The first step of each BED-LLM iteration is to generate a finite pool of candidate questions, $\mathcal{X}^{\text{cand}}$, from which our chosen question will be selected. To ensure we ultimately choose good questions, we must generate a candidate set $\mathcal{X}^{\text{cand}}$ that is rich and diverse enough to include sufficiently informative questions in the first place.
We consider two approaches for generating candidate questions:
  • Unconstrained generation. Given the accumulated history, the LLM is simply asked to propose new questions by sampling from $p_{\mathrm{LLM}}(x_{t+1};h_{t})$ with appropriate prompting.
  • Conditional generation. The LLM is given both the accumulated history and a generated set of hypotheses $\Theta^{\text{cand}}=\{\theta^{(n)}\}_{n=1}^{N}$, such that we sample from $p_{\mathrm{LLM}}(x_{t+1};[h_{t},\Theta^{\text{cand}}])$. Specifically, the LLM is prompted to propose questions whose answers would “slice” the hypothesis pool into roughly balanced subsets.
For both strategies, we sample $M$ questions independently with a relatively high temperature (we use $T=1.3$ in our experiments), in order to encourage diversity. Conditional generation allows us to “guide” the LLM to propose highly informative questions. However, it risks over-fitting to the sampled set of hypotheses. In practice, we found that conditional generation works well for discrete, easily divisible hypothesis spaces such as those found in the 20–Questions setting, but tends to overfit in the case of more complex, overlapping hypotheses such as persona identification. We can avoid this overfitting by generating two hypothesis sets, one for generating questions and one for estimating EIG. However, this adds significant computational cost, leading us to prefer the unconstrained generation approach in such cases.
We restrict our questions to be multiple-choice questions. This is because having a small space of single-token responses $y$ (e.g., yes/no, A/B/C/D) ensures that $p_{\mathrm{LLM}}\left(y_{t+1};\left[\theta,x_{t+1}\right]\right)$ provides a meaningful uncertainty over outcomes and that we can reliably calculate/estimate its entropy. We discuss this restriction at length in Section˜4.2, where we will see that limitations in LLM uncertainty quantification over unrestricted spaces generally require us to use a restricted space for $y$ (as done here) or $\theta$ (which can be tackled by instead using the data–estimation pairing).
3.2 Estimating EIG for each question
We can estimate the EIG for a given question $x_{t+1}$ using the following Rao-Blackwellized estimator of Equation 3, independently introduced by Gal et al, (2017) and Rainforth, (2017):
$\begin{split}\text{EIG}_{\theta}(x_{t+1};h_{t})\approx&\frac{1}{N}\sum_{n=1}^{N}\sum_{y_{t+1}\in\mathcal{Y}}p_{\mathrm{LLM}}(y_{t+1};[\theta^{(n)},x_{t+1}])\log p_{\mathrm{LLM}}(y_{t+1};[\theta^{(n)},x_{t+1}])\\ &-\sum_{y_{t+1}\in\mathcal{Y}}\hat{p}(y_{t+1};[h_{t},x_{t+1}])\log\hat{p}(y_{t+1};[h_{t},x_{t+1}]),\end{split}$ (5)
where $\hat{p}(y_{t+1};[h_{t},x_{t+1}])\!:=\!\frac{1}{N}\sum_{n=1}^{N}p_{\mathrm{LLM}}(y_{t+1};[\theta^{(n)},x_{t+1}])$ and $\theta^{(n)}\sim p_{f}(\theta;h_{t})$ (see Section˜3.3). While this estimator is biased for finite NN, [^3] In general it is not possible to derive unbiased estimates for the EIG without using MLMC approaches (Goda et al,, 2022) that significantly increase the cost-normalized variance. it is still a singly intractable estimator that converges as $N\to\infty$ at the standard Monte Carlo rate $O(1/\sqrt{N})$ (Rainforth et al,, 2018). Note that the samples $\theta^{(n)}$ do not need to be independent for this to hold, provided that they can be shown to satisfy some appropriate form of ergodicity or decaying correlation (see, e.g., Billingsley, (2013)).
Note that when constructing this estimator, one should directly utilize the LLM’s logits (whenever this is possible) to calculate the $p_{\mathrm{LLM}}(y_{t+1};[\theta^{(n)},x_{t+1}])$ terms. By the Rao-Blackwell theorem (Rao et al,, 1945), this will always produce lower variance estimators than purely sample–based estimators, like those employed in Hu et al, (2024) and Kobalczyk et al, (2025).
3.2.1 EIG cannot be well approximated using predictive entropy
Previous attempts to apply BED to the LLM setting with a prior–likelihood based model have all, either explicitly or implicitly, assumed that the expected likelihood entropy, $\mathbb{E}_{p(\theta;h_{t})}[\mathrm{H}[p(y_{t+1}|\theta;x_{t+1},h_{t})]]$, is constant over designs (Kobalczyk et al,, 2025; Mazzaccara et al,, 2024; Cooper et al,, 2025); for BED-LLM this would equate to ignoring the first line of the estimator in Equation˜5. Under this assumption, the EIG breaks down to simply the marginal predictive entropy, $\mathrm{H}[\mathbb{E}_{p(\theta;h_{t})}[p(y_{t+1}|\theta;x_{t+1},h_{t})]]$. Therefore, approaches that make this assumption are no longer information–based, and are instead making decisions simply based on predictive uncertainty.
This is problematic as, in practice, this expected conditional uncertainty will vary between questions. For example, if $\theta=$“dog”, then the response to the question “Is it an animal?” should be close to deterministic, but the expected response to the question “Does it have black fur?” clearly is not. In general, $\mathbb{E}_{p(\theta;h_{t})}[\mathrm{H}[p(y_{t+1}|\theta;x_{t+1},h_{t})]]$ measures how easily the question can be answered once $\theta$ is known, and intuitively this will not be constant across questions for the vast majority of problems.
This variability in the expected entropy of the likelihood can be crucial in determining good questions and omitting it provides no meaningful computational saving (as the same LLM evaluations are used for the top and bottom lines of Equation˜5, doing the full estimate of the EIG requires no additional LLM calls to be made). We, therefore, strongly advocate against making this assumption: including the expected entropy of the likelihood is essential in avoiding questions that are irrelevant, ambiguous, unclear, or simply unhelpful in our quest to learn about $\theta$. This recommendation will be strongly backed up by our experiments, where we will see that approximating the EIG with the marginal predictive entropy significantly degrades performance of the sequential BED approach. A further example of how predictive entropy can be detrimental is given in Figure˜1.
One might instead be tempted to replace $\hat{p}(y_{t+1};[h_{t},x_{t+1}])$ with $p_{\mathrm{LLM}}(y_{t+1};[h_{t},x_{t+1}])$ in the EIG estimation, as the two essentially offer alternative predictive distributions for the outcome. We also advise against this though, noting that it again provides no meaningful computational benefits (unless one also assumes a deterministic likelihood, but this would then mean we no longer consider $\theta$ at all). A key reason for avoiding this substitution is that it would mean we are no longer estimating a true EIG: the inconsistency between the likelihood and the marginal data distribution means there is no longer a joint model where we are minimising our expected uncertainty in $\theta$. We also find that the LLM process of sampling $\theta$ from $p_{f}(\theta;h_{t})$ followed by $y$ from $p_{\mathrm{LLM}}\left(y_{t+1};\left[\theta,x_{t+1}\right]\right)$ tends to give a better uncertainty over responses than sampling $y$ directly from $p_{\mathrm{LLM}}(y_{t+1};[h_{t},x_{t+1}])$.
3.3 Prior Construction and Belief Updating
The Savage axioms (Savage,, 1954) tell us that a rational agent should update its beliefs in a Bayesian manner. However, doing full Bayesian updates to our model as the history grows is generally impractical for computational reasons in the LLM setting, as it requires approximate inference and this, in turn, typically requires large numbers of expensive likelihood evaluations. Furthermore, the Savage axioms only hold if our (implied) prior truly represents our beliefs, but we find that $p_{\mathrm{LLM}}(\theta)$ is typically heavily over–collapsed to a small number of possible hypotheses and can struggle to convey the full range of possibilities even with careful prompting and a high temperature (c.f. Figure˜8).
Question 1 Question:
Which ice cream flavor feels like the best match for this user?
Choose one option: A. Vanilla B. Dark Chocolate C. Strawberry Swirl D. Mint Chocolate Chip
Predictive Entropy: Very High
EIG: 0
Question 2 Question:
Which film genre does the user most prefer?
Choose one option: A. Action B. Sci-Fi C. Comedy D. Horror
Predictive Entropy: High
EIG: High
Figure 1: Predictive entropy vs. expected information gain (EIG) in a film‐preferences elicitation task. Left: very high predictive entropy (answer is completely unknown) but EIG =0=0 because the answer provides no insight into the user’s film preferences. Right: both predictive entropy and EIG are high as the answer is uncertain, but different answers would lead to markedly different posterior updates, making it informative for learning film preferences. This thus demonstrates how the two criteria can select different questions.
A natural tractable alternative is to simply derive our beliefs through in-context updates, that is, use $p_{\mathrm{LLM}}(\theta;h_{t})$, noting that this has been shown to behave differently to Bayesian updating in the LLM context (Falck et al,, 2024; Kossen et al,, 2024). Unfortunately, we find that the ability of current LLMs to fully incorporate the information from $h_{t}$, even when using large state-of-the-art LLM variants such as GPT-4o (OpenAI,, 2024) or Mistral Large (Mistral AI Team,, 2024), is surprisingly lacking. In particular, we find that all LLMs we evaluate regularly sample hypotheses $\theta$ that are incompatible with previous observations in $h_{t}$, especially as the length of the history grows. We also, again, find problems with over–collapsing, with LLMs often jumping to conclusions about the value of $\theta$ without appropriate evidence for this certainty in $h_{t}$.
The shortfalls of in-context learning in such settings have also previously be noted by, for example, (Liu et al, 2024b, ; Zhang et al, 2025b, ; Zhang et al,, 2024). We posit two reasons why they likely struggle in such settings. First, the information from the different examples in the history are generally highly distinct in these information-gathering settings (indeed, this is part of our aim in adaptively design informative questions), making it harder for the LLM to appropriately reconcile all the provided information than in many other uses of in–context learning. Second, $\theta$ will often represent a user–specific variable that cannot easily be predicted from any data other than the user’s responses to questions: it has been argued that much of the success of in–context learning in LLMs is down to improving problem specification and linking the requested task to data it has seen in its training, rather than truly “learning” from the provided examples (Min et al,, 2022; Kossen et al,, 2024), but the history in our setting is rarely helpful for this due to its user–specific nature.
Addressing these shortfalls is critical to the success of BED-LLM. Although we will still use $p_{\mathrm{LLM}}(\theta;h_{t})$ as the basis for deriving our belief state over $\theta$ (i.e. our intermediate prior), we make various alterations to effectively incorporate historical information and ensure diversity. We refer to our derived distribution as $p_{f}(\theta;h_{t})$ and its key component is to filter the generated hypotheses according to whether they are compatible with the history $h_{t}$. We do this by using the LLM to zero-shot check the compatibility of each sampled $\theta$ with all the previous question–answer pairs in $h_{t}$ (using $p_{\mathrm{LLM}}(y;[\theta,x])$) and then rejecting that sample if an incompatibility is found. Specifically, a sample is rejected if the likelihood of an observed answer falls below a predefined threshold, chosen to balance robustness to model uncertainty against the need to enforce strict historical coherence. To reduce the computational cost of generating and evaluating hypotheses, we further include a hypothesis-retention mechanism: any hypotheses from the previous step t1t\!-\!1 which remain consistent with the most recent question and observation are retained in the hypothesis set for step $t$ without regeneration.
We also make some further variations from $p_{\mathrm{LLM}}(\theta;h_{t})$ to try and combat the over–collapsing issue and encourage diversity (noting that our initial sampling acts like a proposal in a rejection sampler and invalid samples should still be removed by the filtering). First, we sample at a high temperature ($T=1.3$). Second, rather than generating candidates independently, we prompt the LLM to generate batches of candidates using a prompt that requests a “diverse and representative set of candidates”. We can sample multiple batches, with the previous batches in context to force the LLM to generate different samples. Finally, we assume a uniform distribution over the $\theta$ that make it through the rejection, such that we remove any duplicated samples and do not make any direct use of the LLM’s logits on $\theta$. We note that the precise prompting strategy used in $p_{\mathrm{LLM}}(\theta;h_{t})$ can also be quite important in achieving effective performance, with our prompts set up to try and maximize the diversity in the generated $\theta$ samples. Details on our exact setup for $p_{f}(\theta;h_{t})$ are in Appendix˜B.
Under the aforementioned setup for $p_{f}(\theta;h_{t})$, the original prior $p(\theta)$ only affects the EIG at the first step as it has no direct bearing on $p_{f}(\theta;h_{t})$ for $t>0$. We, therefore, define our original prior to simply be equal to $p_{f}(\theta;h_{t})$ with $h_{t}=\emptyset$, such that no filtering is performed. We note though that in some cases it may be preferable to specify some customized $p(\theta)$, potentially then adjusting the filtering process and/or weighting the accepted samples in a non–uniform manner to ensure this prior information is not lost as more data is gathered.
4 On the Specification of $p(\theta,y_{t+1};h_{t},x_{t+1})$, and its Implications
As we described in Section˜3, adapting BED to the LLM setting hinges upon how we specify, and update, the joint distribution $p(\theta,y_{t+1};h_{t},x_{t+1})$. As we discuss in this section, this choice is critical to the success of BED-LLM and impacts how we model the current uncertainty about $\theta$; how easily and accurately we can estimate EIG; and how aligned the model’s actual information gain under a question is predicted by the EIG, which ultimately governs whether the questions we choose are actually good at reducing uncertainty about $\theta$.
As discussed at the start of Section˜3, there are two distinct ways to derive the joint model from our LLM: using a prior–likelihood pairing, $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}p(\theta;h_{t})p\left(y_{t+1};\left[\theta,x_{t+1}\right]\right)$}, or a data–estimation pairing, $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}p\left(y_{t+1};[h_{t},x_{t+1}]\right)p\left(\theta;[h_{t},x_{t+1},y_{t+1}]\right)$}, [^4] One could in principle also consider prompting the LLM to generate $\theta$ and $y$ jointly as well, but we will omit these from our discussions on the basis that this will generally be impractical from an estimation perspective, while LLMs’ autoregressive nature means that this should be equivalent to one of the aforementioned joints, up to effects in prompting strategy and API handling etc. where each of the $p$ terms either corresponds directly to the LLM’s distribution over outputs given the provided context and suitable prompting (i.e. the corresponding $p_{\mathrm{LLM}}(\cdot;\cdot)$), some valid distribution derived from this, or an externally specified distribution. The first construction mirrors deriving our beliefs about $\theta$ from a conventional Bayesian posterior with a concrete prior and likelihood derived (at least partially) from the LLM, whereas the second is analogous to a one-sample martingale posterior that directly samples hypothetical data and draws inferences on $\theta$ given hypothetical data using in-context learning (Fong et al,, 2023; Falck et al,, 2024).
In our outlined BED-LLM approach (c.f. Section˜3), we adopted a prior–likelihood pairing. Below, we will expand on some reasons to prefer this formulation and discuss certain settings where the data–estimation might be preferable.
4.1 Modeling flexibility
The most obvious relative merits of the prior–likelihood and data–estimation pairings are in the flexibility in how each term is chosen. The prior–likelihood pairing gives us greater flexibility to construct a prior set of beliefs over $\theta$ that is distinct to the LLM’s internal beliefs, as it allows us to directly control this prior by changing $p(\theta;h_{t})$, whereas the prior is only implicitly defined in the data–estimation pairing. In Section˜3.3 we exploited this flexibility through our definition of $p_{f}(\theta;h_{t})$, wherein we made various adjustments to $p_{\mathrm{LLM}}(\theta;h_{t})$ to encourage diversity and better incorporate information from the history.
On the other hand, the data–estimation pairing could provide some beneficial flexibility in specifying how the data itself is simulated through changing $p\left(y_{t+1};[h_{t},x_{t+1}]\right)$. Most notably, in some applications like scientific discovery, we may have access to external simulators that can be run to predict our data, with the LLM used instead primarily to automate how we draw conclusions from such hypothetical data (Ghareeb et al,, 2025).
In the cases considered in this paper, we do not have such external, validated simulators, and the flexibility in prior construction is more important for our needs.
4.2 Faithfulness of conditional distributions
In Section˜4.1, we considered flexibility in specifying the marginal distributions $p(\theta;h_{t})$ and $p\left(y_{t+1};[h_{t},x_{t+1}]\right)$. While deviations from relying on direct LLM predictions are also in principle possible for the conditional models $p\left(y_{t+1};\left[\theta,x_{t+1}\right]\right)$ and $p\left(\theta;[h_{t},x_{t+1},y_{t+1}]\right)$, in practice, these will typically be more difficult and expensive to implement. This is first because these conditionals need to be instantiated for each sampled instance of the conditional variable ($\theta$ and $y_{t+1}$ respectively), rather than just needing us to set up one marginal distribution that can then be repeatedly sampled from. Second, to construct estimators for Equations (2) and (3), we require access to concrete probabilities for the conditional distributions (in order to calculate entropies), whereas we only needed to draw samples for the marginal distributions (in order to approximate expectations). As such, the conditionals need to be explicit distributions, or at least distributions where the probability can be cheaply estimated (noting that such estimation will be nested and thus more expensive than in conventional settings (Rainforth et al,, 2018)), so they are more difficult to define through the output of some algorithmic procedure, especially if the number of possible realizations is large.
When considering the conditional distributions, the decisive question on the relative merit of the two formulations is which conditional factor we are willing to trust the LLM to supply as a full probability distribution. Critically, we rely on how the LLM captures uncertainty in this full distribution—including, for example, tail behavior—not merely the fidelity of typical samples; the marginal factors, by contrast, only need to be sampled from. If we accept the LLM’s direct predictive distribution for $p\left(y_{t+1};\left[\theta,x_{t+1}\right]\right)$, then we are basing our notion of uncertainty around (and will need to calculate) $\mathrm{H}[p_{\mathrm{LLM}}\left(y_{t+1};\left[\theta,x_{t+1}\right]\right)]$, and if instead we place more faith in the LLM’s internal distribution for $p\left(\theta;[h_{t},x_{t+1},y_{t+1}]\right)$, then we are basing our uncertainty around $\mathrm{H}[p_{\mathrm{LLM}}(\theta;[h_{t},x_{t+1},y_{t+1}])]$. In essence, the choice between prior–likelihood and data–estimation pairings thus comes down to whether we believe the LLM will produce a more appropriate conditional uncertainty over $\theta$ or $y$, along with our ability to numerically estimate this uncertainty cheaply (noting the nesting means we need to repeatedly carry out this estimation for different $\theta$ and $y$ samples respectively).
This difference becomes particularly noticeable when the complexities of the spaces of $\theta$ and $y$ differ significantly. Our ability to draw sensible samples of either will generally be quite robust to these spaces being complex or high–dimensional; this is where LLMs tend to thrive, effectively generating highly complex outputs in an autoregressive manner. However, evaluating the entropy of a distribution becomes dramatically harder as the dimensionality or complexity increases (Acharya et al,, 2019; Paninski,, 2003), and the entropy of the predictive distribution of an LLM in such cases will not typically provide a sensible measure of uncertainty (Kadavath et al,, 2022; Desai & Durrett,, 2020). As such, the decision between joint formulations should predominantly be based on the complexity of the space of $\theta$ versus that of $y$: we should generally favor the prior–likelihood formulation if $\theta$ is more complex and the data–estimation formulation if $y$ is more complex. For example, suppose $\theta$ represents the space of Python code a user wants to generate and $y$ is the answer to a simple yes/no question about that code’s desired behavior. Here, the LLM’s entropy over yes vs. no when given the code will be a much more meaningful uncertainty measure than its entropy over code when told the answer to the question.
For the problems that we consider, the space of $y$ is less complex than that of $\theta$, indicating we should, in general, use the prior–likelihood formulation. However, in cases where this is not true, the data–estimation formulation may be preferable instead. We provide additional discussion on the impact of the chosen pairing on our entropy estimate, plus discussion on the choice of $\theta$, in Appendix˜A.
4.3 Extracting the belief state
A further advantage of the prior–likelihood construction is that our belief state on $\theta$ can be extracted directly, as it is simply $p_{f}(\theta;h_{t})$. With the data–estimation construction, we would have to estimate a marginal distribution from $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}p\left(y_{t+1};[h_{t},x_{t+1}]\right)p\left(\theta;[h_{t},x_{t+1},y_{t+1}]\right)$} by integrating out the synthetic response $p(\theta;h_{t})=\int{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}p\left(y_{t+1};[h_{t},x_{t+1}]\right)p\left(\theta;[h_{t},x_{t+1},y_{t+1}]\right)}\;\mathrm{d}y_{t+1}$. Direct access to $p(\theta;h_{t})$ is also important to ensure that our current belief state is independent of the next question $x_{t+1}$, as the marginal obtained from the integral above is unlikely to remain the same for all $x_{t+1}$ in the way we intuitively want it to.
4.4 Alignment between EIG and belief updating procedure
Our ultimate goal is to minimize uncertainty in $\theta$, as measured by its entropy. With this in mind, we can use the expected uncertainty reduction framework of Bickford Smith et al, (2025) to provide insights into how well our EIG formulation and belief updating procedures align.
To simplify discussions, for now we consider the setting where we choose a single question $x$ and obtain a response $y$. Following Bickford Smith et al, (2025), we can think of the “true” optimal design as selecting
$x^{*}_{\mathrm{true}}=\mathrm{argmin}_{x}\mathbb{E}_{p_{\mathrm{true}}(y;x)}\left[\mathrm{H}[p(\theta;x,y)]\right],$ (6)
where $p_{\mathrm{true}}(y;x)$ is the true response distribution and $p(\theta;x,y)$ is our belief state after the experiment.
Note here that true optimal design has no direct dependency on our current beliefs about $\theta$; it only depends on $p_{\mathrm{true}}(y;x)$ and the hypothetical beliefs we produce for given observed data, $p(\theta;x,y)$. Thus, we can now see that our choice of joint model corresponds to different choices for approximating these quantities. Assuming that the LLM distribution is used directly for the conditional as per Section˜4.2, we thus have that:
  • The prior-likelihood pairing equates to the approximations $p_{\mathrm{true}}(y;x)\approx\int p(\theta)p_{\mathrm{LLM}}(y;[\theta,x])d\theta$ and $p(\theta;x,y)\approx p(\theta)p_{\mathrm{LLM}}(y;[\theta,x])/\int p(\theta)p_{\mathrm{LLM}}(y;[\theta,x])d\theta$;
  • The data–estimation pairing equates to directly specifying a model for $p_{\mathrm{true}}(y;x)$ and then using the approximation $p(\theta;x,y)\approx p_{\mathrm{LLM}}(\theta;[x,y])$.
The appropriateness of each of these options, therefore, comes down to how faithful these approximations are respectively to the true data distribution, $p_{\mathrm{true}}(y;x)$, and how we actually derive our belief distribution on $\theta$ in practice once we have seen the new data.
The former of these considerations is difficult to control for as we simply do not know the true response distribution and it is hard to say which approach will thus estimate it best (though we can refer to the discussion in Section˜4.1 to determine which best matches our beliefs about the true response distribution). However, we do know upfront how we plan to derive our belief distribution on $\theta$ in practice, so we can use this to guide which joint model we formulate our EIG from. Namely, we see that using
  • The prior–likelihood EIG pairing equates to assuming we will make a Bayesian update to our beliefs on $\theta$ using the likelihood $p_{\mathrm{LLM}}(y;[\theta,x])$;
  • The data–estimation EIG pairing equates to assuming we will make an in-context update to our beliefs on $\theta$, as we are treating $p(\theta;x,y)$ as $p_{\mathrm{LLM}}(\theta;[x,y])$.
Our preference between the pairings should therefore be guided in part by how we plan to update the model in practice. In particular, if we plan to make pure Bayesian updates, then the prior-likelihood formulation will tend to yield an EIG that is more faithful to our updating procedure, while if we only make simple in-context updates, the data–estimation formulation will tend to yield a more faithful EIG instead.
The update we use in practice, namely taking $p(\theta;x,y)=p_{f}(\theta;[x,y])$ as outlined in Section˜3.3, can be seen as being somewhere between the in-context and Bayesian updating: we initially sample from $p_{\mathrm{LLM}}(\theta;[x,y])$, but then perform filtering and other steps. The relative extent to which it resembles each will be problem–dependent and again be linked to how much we trust the LLM to capture uncertainty in the space of $\theta$ vs. $y$.
For the settings we consider, we expect $p_{f}(\theta;[x,y])$ to generally be better approximated by a Bayesian update than an in–context update, aligning with our decision to use the prior–likelihood formulation. The reasons for this are that a) the filtering often removes a large proportion of the generated samples, especially at later experiment turns, with $p_{\mathrm{LLM}}(\theta;h_{t})$ not fully incorporating information from the history; b) the maintaining of the set of one consistent hypotheses from one turn to the next encourages a more Bayesian behavior, with samples persisting unless contradicted by a new likelihood term; and c) the typical premature over–collapse of $p_{\mathrm{LLM}}(\theta;h_{t})$ to a small number of hypotheses means it is typically unrepresentative of our beliefs.
These theoretical benefits are perhaps secondary to the more practical benefits from the ease of constructing an appropriate model in the prior–likelihood formulation and avoiding direct uncertainty estimation in the space of $\theta$. Nonetheless, they help confirm that our choices have not induced unnecessary mismatch between the EIG formulation and our updating procedure.
The analysis here can get a little more complicated when we consider the fact that “one-step optimal” decisions like Equation˜6 may no longer be optimal in the sequential decision–making setting. We provide more discussion on this in Section˜A.3.
5 Related Work
In this work, we consider the broad problem of sequential information gathering using LLMs—that is, adaptively asking natural-language questions to allow the LLM to rapidly learn about a parameter of interest $\theta$. Several works have explored the baseline ability of LLMs to ask such questions (Zhang et al,, 2024; Li et al, 2025b, )—effectively, the “naive" baseline in our experiments. While these works demonstrate some ability to adaptively construct information-seeking questions, they often fail to extract important information (Li et al, 2025a, ), leading us to more sophisticated approaches.
Ours is also not the first paper to use a targeted model-based approach to information gathering with LLMs: several existing works have previously attempted to choose questions based on estimates of the expected information gain, marginal predictive entropy, or similar quantities (Piriyakulkij et al,, 2023; Hu et al,, 2024; Kobalczyk et al,, 2025; Wang et al,, 2025; Mazzaccara et al,, 2024; Chan et al,, 2025; Cooper et al,, 2025). However, as explained below, these existing methods all make design choices that significantly limit performance. Further, none of them provide comparable insights into how the joint model should be best formulated and updated from the underlying LLM; most of them even fail to directly link their approach and BED itself.
A number of these previous works, including all that also employ a prior–likelihood pairing in their joint model, actually implicitly select queries using predictive entropy in practice (rather than EIG as claimed) because they assume a constant expected likelihood entropy (Kobalczyk et al,, 2025; Mazzaccara et al,, 2024; Cooper et al,, 2025). As we discussed in Section˜3.2.1 and show empirically in Section˜6, this simplification is not generally reasonable—as predictive entropy is not in general a good approximation to EIG—and leads to worse performance. We believe our work to be the first which uses a prior–likelihood model pairing without simplifying the EIG to just the predictive entropy.
All of the other methods have instead relied on models derived from a data–estimation pairing, then either obtaining a sample-based estimate of $\mathbb{E}_{p\left(y_{t+1};[h_{t},x_{t+1}]\right)}\left[\mathrm{H}[p_{\mathrm{LLM}}\left(\theta;[h_{t},x_{t+1},y_{t+1}]\right)]\right]$ (Piriyakulkij et al,, 2023; Hu et al,, 2024; Wang et al,, 2025), or in the case of Chan et al, (2025), an approximation to this entropy based on the size of conformal sets. As discussed in Section˜4, our prior–likelihood model pairing provides significant benefits, particularly in the large or unbounded hypothesis spaces we explore in this paper. To get an approach which is based on the data–estimation model pairing to work, previous works have thus had to restrict the set of allowable hypotheses (i.e. $\theta$) to a small set, typically also restricting the space of allowing outcomes and further making a deterministic likelihood assumption (thereby again implicitly reducing the EIG to predictive entropy). These restrictions significantly limit applicability and performance.
Rather than augmenting a frozen LLM with the ability to estimate utility functions, some works have instead aimed to post-train an LLM to improve its ability to ask questions (Zhang et al,, 2024; Wu et al,, 2025; Andukuri et al,, 2024). Most do not explicitly consider informativeness of questions: Zhang et al, (2024) and Wu et al, (2025) use reinforcement learning techniques to reward generations that quickly lead to the correct answer, and Andukuri et al, (2024) builds on Li et al, 2025b by fine-tuning on successful traces. Mazzaccara et al, (2024) do indirectly incorporate uncertainty: they use predictive entropy to identify informative questions, and then either fine-tune on the highest-entropy question, or perform DPO comparing the highest-entropy question with a lower-entropy question. We do not address fine-tuning in this work, focusing instead on exploring the correct way to formulate BED using LLMs.
As discussed in Section˜4, a key challenge in adapting BED to the LLM setting is in aligning the expected information gain with the actual uncertainties extracted from the LLM after updating. Handa et al, (2024) take a different approach to this problem by using the LLM to generate features for an external conventional Bayesian joint model (in their case, a linear Bradley Terry model), rather than deriving their joint model more directly from the LLM itself. This can be a good choice when we already have a well-specified Bayesian model form for the problem at hand; however, this may be challenging in arbitrarily large and complex hypothesis spaces. In particular, their specific method is not applicable more widely beyond the preference learning context they consider.
6 Experiments
We now empirically evaluate our proposed BED-LLM method, compare it against relevant baselines, and analyze its properties through targeted ablation studies. We consider two interactive question–answering scenarios that test different types of sequential information acquisition: 1) 20–Questions: a deductive reasoning game requiring identification of a hidden target entity through binary (Yes/No) queries, and 2) Active Preference Elicitation: a task in which the goal is to infer a latent user preference vector via multiple–choice questions.
To evaluate the different questioner methods (i.e., BED-LLM and our baselines), we further need some form of answerer that represents the user or external environment we are trying to learn about. Due to the complications of evaluating with real human participants, we instead conduct our tests with a simulated user based on a separate LLM instance that has been given a ground truth target $\theta^{*}$ (corresponding to the hidden target entity for 20–Questions and a user profile for the active preference elicitation). Note that this answerer LLM does not itself perform any learning—it does not keep previous question–response pairs in context. While we consider both deriving the questioner and answerer from the same LLM and from different LLMs, we emphasize that they are always using separate LLM instances with different prompts and contexts, thus no information is transferred between them except in the questions and responses themselves.
We compare BED-LLM against two baselines:
a) Entropy: selecting questions using the maximum predictive entropy, as done in Kobalczyk et al, (2025), Mazzaccara et al, (2024), and Cooper et al, (2025). For fairness and to ensure a strong baseline, we keep the entropy setup as close to BED-LLM as possible (including prompts, candidate question generation, and our joint model construction with hypothesis filtering), varying only in replacing the EIG estimator with an estimator of the marginal predictive entropy,
$\mathrm{Entropy}_{\theta}(x_{t+1})=-\sum_{y_{t+1}\in\mathcal{Y}}\hat{p}(y_{t+1};[h_{t},x_{t+1}])\log\hat{p}(y_{t+1};[h_{t},x_{t+1}]).$ (7)
We construct an estimator for Equation˜7 analogously to Equation˜5. We also note here that our version of the Entropy baseline produces noticeably stronger performance on the 20–Questions problem than Kobalczyk et al, (2025) report for their own “EIG” approach, demonstrating it efficacy. [^5] Though Kobalczyk et al, (2025) describe their approach as targeting the EIG, the estimator they use (given by their Algorithm 1) actually equates to a sample-based estimator of the marginal predictive entropy instead and is thus comparable with our Entropy baseline. As they do not provide code specifically for the 20–questions experiments or explicit results on their approaches’ success rate, we were unable to provide direct comparisons against their exact setup. However, we believe that it is still clear from the results they do present that their approach is performing worse than the Entropy baseline we do consider. First, the reported ranks they give indicate they are not beating the naive LLM approach with the same regularity as Entropy does. Second, in their rebuttal on OpenReview they explain that they only report ranks and not success rate because “only very few runs would result in a non-zero accuracy”, whereas we see that our Entropy baseline is regularly successful.
b) Naive QA:we prompt the questioner LLM to directly produce an informative next question without explicit hypothesis generation, EIG computation, or filtering. The question is obtained by drawing a single sample at temperature $T=1$ from the questioner LLM. This approach, which was explored in Zhang et al, (2024), relies on the implicit reasoning capabilities of the LLM to select good questions.
6.1 20–Questions
We first evaluate BED-LLM in the classic “20–Questions” game, where the questioner must identify a hidden target $\theta^{*}$ by asking up to 20 Yes/No questions. Here, the hypothesis space of possible $\theta$ is large and not explicitly defined or restricted: it is bounded only by what the LLM can generate. As such, all previous work that has considered using BED-style decision with LLMs either a) cannot be directly applied in our precise problem setup because it requires the questioner to be given access to an explicit, restricted set of possible $\theta$ (Hu et al,, 2024; Piriyakulkij et al,, 2023; Wang et al,, 2025; Chan et al,, 2025), or b) is a variant of our Entropy baseline, due to using marginal predictive entropy in lieu of the EIG (Kobalczyk et al,, 2025; Mazzaccara et al,, 2024; Cooper et al,, 2025).
We evaluate across three distinct datasets each containing a mix of 100 obscure and common targets corresponding to either Animals, Celebrities, or Things. Here, the dataset is just a list of different $\theta^{*}$ that will be individually provided to the answerer to instantiate different problems (e.g. we conduct a trial where $\theta^{*}=$“dog”, then one where $\theta^{*}=$“cat”, etc). The list of targets is never provided to the questioner model to restrict the set of possible hypotheses: the questioner is only prompted that is trying to identify an “animal”, “celebrity”, or “thing”. The datasets are: a) Animals: a set of animal species generated with OpenAI’s o3 model to ensure a diverse mix and balanced taxonomy, see Appendix˜D for details; b) Celebrities: a diverse set of public figures, as used by Zhang et al, (2024); c) Things: a collection of everyday and exotic entities drawn from the web corpus, as used by Zhang et al, (2024); it covers a wide range of categories, from plants and clothing to professions, events, and mythical creatures. We note that the best success rates achieved by Zhang et al, (2024) on Celebrities and Things are substantially below those of both BED-LLM and our Entropy baseline.
Using our sample–then–filter process (c.f. Section˜3.3), we aim to sample at least $N=15$ hypotheses, repeating the cycle up to three times if needed (the exact number of hypotheses can be less than this as it may not be possible to generate sufficient valid hypotheses, especially in later experiment turns). The questioner generates $M=15$ candidate questions to test, $\mathcal{X}^{\text{cand}}$, using the “conditional generation” approach of Section˜3.1 when possible, but falling back on “unconditional generation” if insufficient candidate hypotheses have been generated.
Evaluation
[…]
Table 1: Success rate (%) for 20–Questions at the end of the game. Best result in bold. Success rates are integers because each dataset contains exactly 100 entities. $\pm$ numbers represent the standard error of the mean estimate (this is estimated using $\sqrt{p(1-p)/(n-1)}$ where $p$ is the success percentage and $n$ is the number of datapoints, noting this estimator is positively biased and thus conservative).
Success Rate (%) $\uparrow$
Dataset Model Naive QA Entropy BED-LLM
[…]
Figure 2: Success rate against turn for the 20–Questions game. Error bars represent the standard error of the mean estimate across 100 target entities in each dataset. The same model used as both questioner and answerer.
Figure 3: Success rates for the 20–Questions game with different models: the questioner Q and answerer A.
Results
First, we evaluate in the setting where the same LLM is used as both the questioner and the answerer. Table˜1 shows the success rates for the different datasets and a variety of different LLMs. Across all datasets and all LLMs, BED-LLM achieves the highest success rate, significantly outperforming the Entropy and Naive QA baselines. It is particularly notable how the success rate of our principled information-gain strategy is typically more than double that of Naive QA which relies solely on implicit LLM reasoning.
Per-turn success rates are shown in Figure˜2. Over the first few turns, BED-LLM and Entropy perform similarly. However, after around 7-10 turns we consistently see BED-LLM pull away from Entropy. This may simply be because it takes some time for the effect of choosing good questions to take effect (the large space of possible $\theta$ means that no approach can reliably determine the answer in only a few steps), or it may be that the effect of using proper EIG estimates becomes more pronounced as the set of viable hypotheses decreases and it is harder to find useful questions.
Different questioner and answerer
In the experiments above, the same LLM served as both the questioner and the answerer. This setting implicitly assumes that the questioner model has access to the answerer’s response distribution, i.e. that there is no misspecification in the likelihood used by BED-LLM and Entropy or the internal reasoning of Naive QA. However, in practice, the answerer is typically a human whose responses will follow a different distribution to the questioner’s internal model for reasoning about responses. To test the robustness of the various methods to this model misspecification, we therefore also evaluate scenarios where the questioner and answerer are formulated with entirely different LLMs, instead of just different instances of the same LLM. As shown in Figure˜3, the benefit of BED-LLM persists even under this model misspecification; it consistently achieves higher success rates than both baselines across datasets. In contrast, Entropy exhibits a more severe drop in performance in this more realistic setting, often now performing more in line with Naive QA, which itself appears to be quite robust to the misspecification.
Figure 4: Data–estimation ablation using Qwen2.5-72B-Instruct on Celebrities dataset.
Data–estimation ablation
As described in Section˜3, an important design choice is whether we choose the prior–likelihood pairing or the data–estimation pairing for BED-LLM. As an ablation for this critical decision, we also implement a variant of the alternative Data–Estimation method (see Appendix˜C for details). This gives an approach that is similar to those of Hu et al, (2024); Piriyakulkij et al, (2023); Wang et al, (2025); Chan et al, (2025), but which does not require us to assume a single fixed set of possible hypotheses throughout the sequential BED process. We evaluate this alternative approach on the Celebrities dataset using Qwen2.5-72B as both the questioner and answerer. Results are shown in Figure 4. Consistent with our analysis in Section˜4, BED-LLM substantially outperforms the Data–Estimation alternative. Data–Estimation still outperforms the Naive QA baseline, indicating that even this pairing benefits from explicit EIG–driven question selection over purely heuristic question generation, but interestingly does worse than Entropy, emphasizing the importance of estimating our uncertainty in the $y$ space instead of $\theta$ space. These findings reinforce our claim that the choice of joint–model factorization is a critical design decision for sequential BED in the LLM setting.
6.2 Active Preference Elicitation
In our 20–Questions problems, the generated hypotheses were all single, concrete entities (barring any failures of the LLMs) and most reasonable questions had clear–cut answers (e.g. “is Beyoncé a singer?”). In contrast, many problems, especially those related to eliciting information from a user, involve more abstract targets and/or more subjective and less predictable answers. For example, if we are learning about the characteristics or preferences of a user, then it may be difficult to explicitly define a concrete closed set of possible $\theta$, while it is also quite challenging for the LLM to accurately predict user responses and develop appropriate uncertainty estimates.
To study such a scenario, we evaluate BED-LLM on inferring users’ film preferences. Here the target we are trying to gather information about is somewhat abstract, and we have some flexibility in how we define $\theta$ in our joint model. Our chosen setup is to define $\theta$ to be a natural language paragraph describing a user’s taste in films; we refer to this as a persona. To emulate a human user, our answerer is set up to be an LLM prompted to answer questions as if it were a person with a given persona $\theta^{*}$.
To generate a dataset of ground-truth personas, we take a set of $200$ real user ratings from the MovieLens-100K dataset (Harper & Konstan,, 2015), then use an “oracle” LLM (namely, OpenAI’s o3 model) to produce a natural language persona that is consistent with each distinct set of ratings. As was the case for the 20–Questions problems, this dataset is never provided to the questioner and the set of allowed $\theta$ is not constrained.
Because we need the LLM to be able to meaningfully capture uncertainty in the space of responses as discussed in Section˜4.2, we restrict questions to be multiple-choice. Specifically, the questioner is tasked with producing a question along with five possible responses A/B/C/D/E. We then define each $x_{t}$ to be the question coupled with the possible responses, and each $y_{t}$ to be one of the letters A/B/C/D/E to provide a restricted set of tokens over which we can measure entropy. Option E is further constrained to always be “none of the above” so that the answerer is not committed to choosing one of the directly generated choices if none are suitable.
For BED-LLM and Entropy, we compare $M=8$ candidate questions at each turn and we aim to generate at least $N=5$ candidate hypotheses. We use the “unconstrained generation” approach of candidate question generation (c.f. Section˜3.1) as the user profiles can be quite diffuse and we are only generating a small number of possible hypotheses that can be quite easy to split.
We note that data–estimation setup is not at all viable for this problem because the large number of tokens and varying dimensionality of each $\theta$ sample mean that $\mathrm{H}[p_{\mathrm{LLM}}(\theta;[h_{t},x_{t+1},y_{t+1}])]$ is not only infeasible to estimate, but also is not a meaningful measure of uncertainty.
(a) Qwen2.5-72B
(b) GPT-4o-mini
(c) Llama-3.3-70B
(d) GPT-4o
Figure 5: Results for the active preference elicitation with the same LLM used for the questioner and answerer. Shown is the average ratings of 10 recommendations for each of 200 separate simulated users.
Evaluation
To again allow tracking of progress through the experiment, after each turn of the interaction $t$, the questioner generates ten film recommendations, conditioned on $h_{t}$. These recommendations are checked for consistency with prior questions and answers; if any are judged inconsistent then they are removed and additional recommendations are generated. The quality of the film recommendations is then assessed using an “LLM-as-judge” protocol (Zhu et al,, 2025; Trivedi et al,, 2024). Namely, the answerer evaluates each of the 10 films recommended by the questioner, conditioned on the hidden target user profile $\theta^{*}$. It scores each film on a scale of 1 to 5 (in 0.5 increments), based on how well the recommendation aligns with $\theta^{*}$ — this score is output together with a brief justification to increase reliability. We report the mean rating and standard error across 200 users, over 5 question–answer turns.
Results
In Figure˜5, we show the results using the same model for both questioner and answerer. For all models, BED-LLM establishes an advantage over other baselines by turn $t=3$, although the differences in performance between methods are smaller than for the 20–Questions game. On average, Entropy performs similarly to Naive QA.
(a) Q: Qwen2.5-72B. A: GPT-4o-mini.
(b) Q: GPT-4o-mini. A: Qwen2.5-72B.
Figure 6: Results for Active Preference Elicitation with a different models: questioner Q and answerer A.
Different questioner and answerer
Like in the 20–Questions setting, we now also evaluate performance on Active Preference Elicitation in the misspecified model scenario by using a different LLM for the questioner and answerer. Figure˜6 presents the results. While BED-LLM and Entropy seem to be quite robust to the model misspecification, it is now Naive QA that seems to be brittle, exhibiting a drop in performance when using Qwen2.5-72B questioner and GPT-4o-mini answerer.
7 Conclusion
In this work, we have shown how to effectively apply the framework of sequential Bayesian experimental design (BED) to the problem of interactive information gathering with LLMs. Specifically, we have introduced BED-LLM, which provides a specific, information–theoretic, sequential BED approach that makes a variety of carefully justified design choices in the joint-model factorization, belief updating, and candidate question generation. Particularly central to BED-LLM is the prior–likelihood pairing with filtering of hypotheses for consistency with the history. BED-LLM is notably the first work that uses both this prior–likelihood pairing without resorting to using marginal predictive entropy in lieu of the expected information gain (EIG). Together, these innovations lead to substantial performance improvements compared to previous approaches. The results thus confirm that principled EIG–driven strategies can yield substantial gains for interactive, multi–turn, information gathering problems.
Appendix A Additional discussions on design choices
A.1 An alternative view on the faithfulness of conditional distributions
Another way of viewing the distinction between the prior–likelihood and data–estimation constructions is in which of the EIG forms, Equation˜2 or Equation˜3, we center our reasoning. For a given joint model, the two are, of course, mathematically equivalent. However, they give us different ways of thinking about what it means to maximize the EIG: reducing entropy in $\theta$ from seeing $y$, or reducing entropy in $y$ from seeing $\theta$. This, in turn, gives us a way to reason about how appropriate our joint model is. When we choose to use one of $p\left(y_{t+1};\left[\theta,x_{t+1}\right]\right)$ or $p\left(\theta;[h_{t},x_{t+1},y_{t+1}]\right)$, we are centering our reasoning around the entropy of this quantity making sense, while allowing the other entropy in the other form to be implicitly defined from the resulting joint distribution; because the two forms are equivalent, we know that if our explicit form is suitable/unsuitable, the implicit form will be as well. If, for example, we directly fix the form of $p\left(\theta;[h_{t},x_{t+1},y_{t+1}]\right)$ using our LLM’s predictive distribution, we are also directly relying on its expected entropy being a meaningful measure of design quality. If $\theta$ is high–dimensional and predominantly free–form, the resulting entropy produced by the LLM is unlikely to be meaningful and using the data–estimation pairing is unlikely to produce an effective strategy. However, if $y$ is instead quite constrained, the LLM can produce a meaningful entropy over it, and choosing a model based on the prior–likelihood pairing is likely to implicitly define a meaningful distribution, and thus entropy, on $\theta$. Conversely, if $\theta$ is constrained and $y$ is free form, the opposite will hold instead.
A.2 Choice of $\theta$
An important corollary of this reasoning is that it can be important to be careful in our choice of exactly what we take $\theta$ to be, especially if we are using the data–estimation formulation. In particular, it is essential for entropy in the space of $\theta$ to form a meaningful notion of uncertainty, even if this entropy is not being measured through the LLM’s predictive distribution of $\theta$ directly. Thus, while $\theta$ inherently represents what we are trying to learn about and should always be set up as such, if there is flexibility in how exactly we formulate it, we should be careful to choose a form that yields an appropriate uncertainty measure. For example, if the LLM is trying to clarify what code a user wishes it to generate, we could either choose $\theta$ to be the code itself or, following (Neiswanger et al,, 2021; Bickford Smith et al,, 2023), the output the code produces. Here the entropy over code outputs induced by our distribution on code is likely to be a much better measure of uncertainty than the entropy of the raw code itself, given that there are multiple ways one can code the same operation.
A.3 Expected uncertainty reduction in the sequential setting
In Section˜4.4, we looked at whether the prior–likelihood pairing or the data–estimation pairing is better aligned with how we update our beliefs in practice. The picture here can get a somewhat more complicated once we move into the sequential BED setting. Here, our ultimate aim is actually to minimize $\mathrm{H}[p(\theta;h_{T})]$ at some final future horizon $T$. Now, we only care about intermediary belief states $p(\theta;h_{t})$ through their aid in future decision making toward the goal of minimizing the final entropy. Thus, even if we are working with in-context updates, it might be the case that $p(\theta;h_{t})$ only starts to produce a meaningful entropy once we have seen enough data to sufficiently narrow down the possibilities on $\theta$. The optimal behavior in such settings would be to learn a policy that directly targets this final belief state instead of sequentially targeting the incremental EIGs. However, this will typically not be computationally feasible in practice and we instead need to resort to a myopic decision-making strategy. It might thus still be better to use the prior–likelihood formulation in such myopic decision making settings, even if we are sequentially updating our beliefs on $\theta$ through in-context updates, if this allows us to better guide the sequential decisions towards our final objective. The coherence of Bayesian updating means that the converse is unlikely to be true, so this provides further evidence towards using the prior–likelihood formulation.
Appendix B Generating candidate hypotheses for BED-LLM
A fundamental challenge for BED-LLM and the same method with the Entropy baseline is generating a sufficiently diverse set of candidate hypotheses from the LLM prior. As illustrated in Figure˜8, the raw distribution $p_{\mathrm{LLM}}(\theta)$ is highly over–collapsed, often concentrating mass on only a few high–likelihood hypotheses. Thus, it is not practical to directly use the LLM’s distribution as a prior $p(\theta)$. Instead, we jointly sample candidate $\theta$ and assume a uniform distribution over them. We can view this as sampling thetas from a mixture distribution.
Instead, we employ a joint sampling strategy, in which the LLM is prompted to generate a list of $N$ hypotheses in a single rollout, which corresponds to drawing from the autoregressive list distribution
$p_{\mathrm{LLM}}(\theta^{(1)}_{t},\ldots,\theta^{(N)}_{t};h_{t})\;=\;\prod_{n=1}^{N}p_{\mathrm{LLM}}(\theta_{t}^{(n)};[\theta_{t}^{(1:n-1)},h_{t}])$
Each such joint sample is then filtered for history consistency using $p_{\mathrm{LLM}}(h_{t})$. This may be repeated. The surviving hypotheses from the current rollout are then merged with those retained from previous steps to form a cumulative candidate set $\Theta$. While one could in principle reweight candidates using importance sampling, in practice we choose to not rely on the model’s internal probabilities. Instead approximate the prior as a uniform distribution over this union
$p_{f}(\theta;h_{t})\approx\frac{1}{\Theta}\sum_{\theta\in\Theta}\delta_{\theta}.$
To further improve coverage and diversity, we incorporate several refinements in the generation of $p_{f}(\theta;h_{t})$. Due to the differences between different LLMs, the ways this is encouraged may differ, but we find in general that the following are effective:
  • Repeat sampling and filtering: The sampling, filtering, and merging steps may be repeated to encourage diversity and ensure filtering correctness.
  • Prompts: Prompts are designed to elicit stratified hypotheses by encouraging the LLM to consider different semantic features (e.g. age groups, genres, or categories) and implicitly diversify across them.
  • Context ordering: During generation, we reverse the order of the question–answer pairs in $h_{t}$ to place the most recent question at the top of the context window (while retaining earlier exchanges), ensuring that specific constraints are prioritized and mitigating context drift.
Finally, we note that different LLMs respond differently to strategies aiming to increase diversity: some benefit more from a higher temperature while others benefit from more repetitions of the sampling–filtering cycle.
Appendix C Data–estimation method
Our Data–Estimation method is based on a model derived from a data–estimation pairing (Section˜3).
C.1 EIG Estimation
Suppose our model is given by $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}p\left(y_{t+1};[h_{t},x_{t+1}]\right)p\left(\theta;[h_{t},x_{t+1},y_{t+1}]\right)$}. Here, it will clearly be beneficial to directly use Equation˜2 for estimating the EIG, as here we directly have access to all the required terms, other than $p(\theta)$ which can be simply ignored as it is not a function of $x$ so does not affect optimization of the question. If the possible values for $y$ are enumerable and we can evaluate $p\left(y_{t+1};[h_{t},x_{t+1}]\right)$ in closed–form, we can directly calculate the exact EIG (up to a constant) without requiring any estimation at all:
$\text{EIG}_{\theta}(x)-\mathrm{Const}=-\sum_{y}p\left(y_{t+1};[h_{t},x_{t+1}]\right)\mathrm{H}\left[p\left(\theta;[h_{t},x_{t+1},y_{t+1}]\right)\right],$ (8)
where the entropy $\mathrm{H}\left[p\left(\theta;[h_{t},x_{t+1},y_{t+1}]\right)\right]$ can be evaluated directly from the logits of the LLM, or if these are not available, estimated by sampling. If we cannot enumerate $y$ or evaluate $p\left(y_{t+1};[h_{t},x_{t+1}]\right)$, we can simply instead resort to Monte Carlo and use the estimator:
$\text{EIG}_{\theta}(x)-\mathrm{Const}\approx-\frac{1}{N}\sum_{n=1}^{N}\mathrm{H}\left[p(\theta;[c,x,y_{n}])\right]\quad\mathrm{where}\quad y_{n}\sim p\left(y_{t+1};[h_{t},x_{t+1}]\right).$ (9)
We provide an overview of how to do this in the LLM setting in Algorithm˜1.
Algorithm 1 Data–Estimation Selection at Turn $t$
1:History $h_{t}$; candidate questions $\mathcal{X}_{\text{cand}}^{t}$; answer sets $\{\mathcal{Y}(x)\}$
2:Selected question $x_{t}^{\star}$
3:for each $x\in\mathcal{X}_{\text{cand}}^{t}$ do
4:  Obtain predictive answer distribution $p_{\text{LLM}}(y;x,h_{t})$ for all $y\in\mathcal{Y}(x)$
5:  for each $y\in\mathcal{Y}(x)$ do
6:   Construct $\Theta_{t,y}\sim p_{f}(\theta;h_{t}\cup(x,y))$
7:   Compute entropy $H_{y}\leftarrow\mathrm{H}\!\left[p_{\mathrm{LLM}}(\theta;h_{t},x,y)\right]$
8:  end for
9:  Compute $\text{EIG}(x)\leftarrow-\sum_{y\in\mathcal{Y}(x)}p_{\text{LLM}}(y;x,h_{t})\;H_{y}$
10:end for
11:Select $x_{t}^{\star}\leftarrow\arg\max_{x\in\mathcal{X}_{\text{cand}}^{t}}\text{EIG}(x)$
12:return $x_{t}^{\star}$
C.2 Generating candidate hypotheses
To generate candidate values of $\theta$ for the data–estimation method, we used the prompt in Figure˜7. Figure˜8 shows an example of the distribution of samples obtained following two rounds in the 20–questions game. Note that the samples are highly concentrated on just a handful of answers. This lack of diversity shows that the model’s belief distribution is far more concentrated relative to the variability over valid hypotheses in the ground–truth task distribution, which negatively impacts the performance of the data–estimation method.
Return only the full name of one randomly selected famous person (living or deceased) consistent with the questions and answers above.
To increase randomness:
1. Internally brainstorm a pool of diverse and representative individuals.
2. Avoid defaulting to the most globally ubiquitous celebrities or famous figures.
Output rules:
- Output ONLY the person’s full name (with spaces, capitalization and accents), nothing else.
- No extra words, explanations, numbering, or punctuation beyond what’s in the name itself (hyphens/apostrophes allowed if part of the name).
Figure 7: Prompt for generating hypotheses (and evaluating their probability) for the data–estimation method.
Is this person known for their contributions to science?
No.
Is this person known for their contributions to the arts?
Yes.
{"Vincent van Gogh": 93,
"Georgia O Keeffe": 10,
"Frida Kahlo": 37,
"Salvador Dali": 44,
"August Wilson": 8,
"Auguste Rodin": 8}
Figure 8: An example of the sample distribution generated using the prompt in Figure˜7, conditioned on the two question/answer pairs at the top of this figure. At this stage of the game, we independently sample 200 hypotheses from the LLM and record their frequencies. Note that the distribution exhibits strong mode collapse, with most of the mass highly concentrated on just a few answers, which negatively impacts the performance of the data–estimation method. This summary is for diagnostic purposes: Algorithm˜1 operates on the probabilities of individual samples and never instantiates such an aggregated summary.
Appendix D Experimental details
D.1 Generation of the Animals Dataset
To create the Animals dataset, we prompted OpenAI o3 (OpenAI,, 2025, o3-2025-04-16) to generate a list of animals, using the prompt in Figure˜9. The resulting list is shown in Figure˜10. Alternative names (after $|$ character) were manually added.
You are a zoologist. Please generate a list of 100 living animal species with very high taxonomic diversity, including diversity in phyla, classes, orders, and families. Present each animal on a different line.
Figure 9: Prompt for Animals dataset generation.
African elephant
Bengal tiger
Bald eagle
Blue whale
Red kangaroo
Giant panda
Snow leopard
Green sea turtle
American alligator
Bottlenose dolphin
Emperor penguin
Great white shark
Golden poison frog | Golden poison dart frog
Honey bee
Monarch butterfly
Okapi
Chimpanzee
Arctic fox
Komodo dragon
Giraffe
Cheetah
Hammerhead shark
Axolotl
Orca
Puffin
Red panda
Platypus
Rhinoceros beetle
Tasmanian devil
Wombat
Sloth
Blue-ringed octopus | Blue ringed octopus
Manatee
Narwhal
Sea otter
Coral snake
King cobra
Harpy eagle
Lemur
Koala
Aye-aye | Ayeaye
Snowy owl
Elk
Wolverine
Caracal
Cassowary
Quokka
Pangolin
Saiga antelope
Galápagos tortoise | Galapagos tortoise
Sumatran orangutan
Red-eyed tree frog | Redeyed tree frog
European badger
Moose
African grey parrot
Scarlet macaw
Black mamba
Albatross
Humpback whale
Dugong
Anaconda
Kookaburra
Coyote
Brown bear
Golden jackal
Capybara
Ibex
Japanese macaque
Kiwi
Leafcutter ant
Mantis shrimp
Ocelot
Peregrine falcon
Quetzal
Raccoon
Sand cat
Tarantula
Uakari
Vicuña
Wildebeest
Rock hyrax | dassie
Yak
Zebra
Blue dragon nudibranch | Blue dragon sea slug
Chinchilla
Dhole
Electric eel
Flying fox
Gharial
Horseshoe crab
Indigo bunting
Jerboa
Kakapo
Lionfish
Markhor
Nautilus
Olive baboon
Pika
Quoll
Rosy boa
Figure 10: Animals dataset (generated using OpenAI o3, with manual curation)
D.2 LLM models details
The open-weights LLM models used in Section˜6 were obtained from HuggingFace (HF) (Wolf et al,, 2020) and their repositories are: mistralai/Mistral-Large-Instruct-2411, Qwen/Qwen2.5-72B-Instruct, meta-llama/Llama-3.3-70B-Instruct. Since all models are instruction-finetuned, we used the latest chat templates provided as part of the HF implementation. All of the experiments using mistralai/Mistral-Large-Instruct-2411 were conducted by Oxford researchers on Oxford compute infrastructure.
The OpenAI API models used were: gpt-4o-mini-2024-07-18, gpt-4o-2024-08-06.
Appendix E Additional Personalization Experiments
E.1 EIG estimates of questions chosen by different questioners
Figure 11 demonstrates the progression of EIG estimates throughout the turns for BED-LLM and Naive QA. We see that at turn $t=1$, the questions chosen by both methods have similar EIG estimates, which explains the similar performance of BED-LLM and Naive QA at turn $t=1$.
Figure 11: The average information gains for each strategy at each turn for the Qwen2.5-72B vs Qwen2.5-72B example. We see that the initial questions have similar information gains, which explains why the performance of the Naive QA method outperforms EIG at turn 1. Afterwards, the EIG information gains are much larger while the Baseline information gains decrease sharply.

### v2

@@
 Our approach, which we call BED-LLM (Bayesian experimental design with large language models), is based on iteratively choosing questions or queries that maximize the expected information gain (EIG) about the task of interest given the responses gathered previously.
-We show how this EIG can be formulated in a principled way using a probabilistic model derived from the LLM’s belief distribution and provide detailed insights into key decisions in its construction.
-Further key to the success of BED-LLM are a number of specific innovations, such as a carefully designed estimator for the EIG, not solely relying on in-context updates for conditioning on previous responses, and a targeted strategy for proposing candidate queries.
-We find that BED-LLM achieves substantial gains in performance across a wide range of tests based on the 20–questions game and using the LLM to actively infer user preferences, compared to direct prompting of the LLM and other adaptive design strategies.
+We show how this EIG can be formulated (and then estimated) in a principled way using a probabilistic model derived from the LLM’s predictive distributions and provide detailed insights into key decisions in its construction and updating procedure.
+We find that BED-LLM achieves substantial gains in performance across a wide range of tests based on the 20 Questions game and using the LLM to actively infer user preferences, compared to direct prompting of the LLM and other adaptive design strategies.
 1 Introduction
-Despite their many successes, LLMs currently fall short on a crucial aspect of interactive intelligence: proactively seeking out information from a user or external environment in an intelligent and adaptive manner (Laban et al,, 2025; Li et al, 2025c, ).
-For example, they have been shown to perform poorly on problems such as multi-step guessing games (Bertolazzi et al,, 2023; Zhang et al,, 2024), task clarification (Chi et al,, 2024), IT task automation (Jha et al,, 2025), and iterative external tool use (Patil et al,, 2025).
-In particular, while modern LLMs are often capable of producing coherent and insightful questions (or other external queries) in a one-shot manner, they typically struggle to appropriately tailor their questions to previously gathered responses on interactive tasks (Bertolazzi et al,, 2023; Patil et al,, 2025).
-There is, therefore, a pressing need to improve the ability of LLMs to adaptively ask questions and gather information in a targeted manner.
+Intelligent information gathering—the ability to ask the right questions at the right time—is fundamental to effective AI systems.
+However, despite their many successes, LLMs currently fall short on a crucial aspect of interactive intelligence: proactively seeking out information from a user or external environment in an intelligent and adaptive manner (Laban et al., 2025; Li et al., 2025c).
+For example, they have been shown to perform poorly on problems such as multi-turn guessing games (Bertolazzi et al., 2023; Zhang et al., 2024), task clarification (Chi et al., 2024), IT task automation (Jha et al., 2025), and multi-step tool use (Patil et al., 2025).
+In particular, while modern LLMs are often capable of producing coherent and insightful questions (or other external queries) in a single-turn setting, they typically struggle to appropriately tailor their questions to previously gathered responses on interactive tasks (Bertolazzi et al., 2023; Patil et al., 2025).
+There is, therefore, a pressing need to improve the ability of LLMs to adaptively ask questions based on previous responses, and gather information in a targeted manner.
 Such capabilities are essential for a wide variety of problems, such as clarifying user intent, personalizing model behavior to a particular user, or generally acting as effective multi-turn conversational agents.
-They are also critical if we want to use LLMs in data gathering tasks or as automated agents in decision-making pipelines (Wu et al,, 2025).
-In turn, these capabilities are essential for a wide variety of downstream applications from medical diagnostic assistants, targeted code generation, and tutoring systems, to conducting automated surveys, AI–driven scientific inquiries, and market research (Liu et al, 2024a, ; Mandal et al,, 2025; Jacobsen et al,, 2025; Lu et al,, 2024).
-Note that in all these problems it is not enough for the LLM to simply one-shot generate good fixed questions based only on its parametric knowledge.
-Instead, we need it to be able to adaptively choose questions that are good, conditioned on the already-collected user responses.
-We propose to address this challenge using the framework of sequential Bayesian experimental design (BED, Lindley, (1956); MacKay, (1992); Chaloner & Verdinelli, (1995); Sebastiani & Wynn, (2000); Rainforth et al, (2024)), which provides a model-based, information-theoretic mechanism for making adaptive design decisions, given a generative model of the experiment.
-Specifically, we show how the problem of interactive information gathering with LLMs can be formulated as a sequential experimental design problem, wherein we iterate between choosing queries based on maximizing their expected information gain (EIG) and updating our beliefs with the information from the received response.
-Here, the underlying generative model from which the sequential BED procedure is formed is derived from the LLM, and we carefully show how this should be constructed, providing extensive insights on key design decisions.
-We call our approach BED-LLM (which stands for Bayesian Experimental Design with Large Language Models) and show how its success is critically dependent on careful choices for how the EIG is estimated, how candidate queries are generated, and how the model is updated given previous responses.
-For example, we show how approximations to the EIG made in previous work (Kobalczyk et al,, 2025; Hu et al,, 2024) can significantly degrade performance compared to using more carefully constructed estimation strategies.
-We also show how simple in-context learning (Brown et al,, 2020) is typically not sufficient to incorporate all the information from previous query–response pairs, with an additional sample–then–filter strategy introduced to correct this.
-Finally, we introduce a targeted strategy for proposing candidate queries (whose EIG we then evaluate) that directly utilizes our current belief state and provides important improvements over direct generation by the LLM.
+They are also critical if we want to use LLMs in data gathering tasks or as automated agents in decision-making pipelines (Wu et al., 2025).
+In turn, these capabilities are essential across domains ranging from medical diagnosis (Hirosawa et al., 2024), troubleshooting (Jha et al., 2025), preference learning (Handa et al., 2024; Chakraborty et al., 2024; Ouyang et al., 2022), and tutoring systems (Kestin et al., 2025; Liu et al., 2024a), to conducting automated surveys (Aher et al., 2023; Lee et al., 2024; Jacobsen et al., 2025), and AI–driven scientific inquiries (Lu et al., 2024; Mandal et al., 2025).
+Note that in all these problems it is not enough for the LLM to generate full sets of suitable questions up front, we need it to be able to adaptively choose questions that are tailored to the already-collected user responses.
+We propose to address this challenge using the framework of sequential Bayesian experimental design (BED, Lindley (1956); MacKay (1992); Chaloner & Verdinelli (1995); Sebastiani & Wynn (2000); Rainforth et al.
+(2024)), which provides a model-based, information-theoretic mechanism for making adaptive design decisions, given a generative model of the experiment.
+Specifically, we show how the problem of interactive information gathering with LLMs can be formulated as a sequential experimental design problem with a model derived from the LLM, wherein we iterate between choosing queries based on maximizing their expected information gain (EIG) and updating our beliefs with the information from the received response.
+We call our approach BED-LLM and show how its success is critically dependent on our precise model formulation, belief updating procedure, and EIG estimation strategy.
+In particular, we show that it is essential to formulate the model with a precise distribution pairing that does not solely rely on in-context learning to update beliefs and uses the LLM’s uncertainties in the space of answers rather than the more complicated underlying hypothesis space we are trying to learn in.
 Together, we find that these innovations provide substantial performance benefits over directly generating queries from the LLM and more basic approximations of the sequential BED framework.
-Specifically, we first find that BED-LLM provides substantial improvements in the success rate for the 20-Questions problem across a variety of LLMs and target quantities.
-For example, when predicting celebrities with Mistral-Large, we observe an improvement in success rate from 14% to 91%.
+Specifically, we first find that BED-LLM provides substantial improvements in the success rate for the 20 Questions problem across a variety of LLMs and target quantities.
+For example, when guessing celebrities with a small model, Llama-3.1-8B, we observe a 5.8x gain in success rate.
 Second, we demonstrate noticeable improvements in using the LLM for movie recommendations, showing that these benefits hold even when the LLM’s predictive model differs from that of the answerer.
-In summary, we illustrate how the adaptive information gathering capabilities of LLMs can be significantly improved by the careful application of the sequential BED framework to their decision making, thereby making an important step towards their more general usage as interactive agents.
 2 Problem Formulation and Background
-There are two natural ways one might look to improve LLMs’ adaptive information gathering capabilities: we can look to change the LLM itself through finetuning or altering the pretraining process, or we can make changes in how the LLM is used at deployment time.
-Our focus will be on the latter, as this means we can make improvements to performance without requiring access to any additional data or expensive retraining, noting that for information–gathering problems, we will not usually have task–specific data available upfront (e.g.
-we cannot expect our LLM to have any knowledge about a particular user’s preferences).
-However, we emphasize that any improvements that can be made to the LLM itself will be complementary to our deployment–time approach.
+There are two natural ways to improve LLMs’ ability to gather information: modifying the model itself (e.g.
+via test-time- or post-training) or altering how the model is used at deployment time.
+We focus on the latter, since information–gathering tasks rarely provide task-specific data upfront (e.g.
+a user’s unknown preferences), and deployment–time methods avoid the cost and difficulty of finetuning an LLM altogether and are applicable to any existing LLM.
+However, we emphasize that improvements at the model level (e.g., Zhang et al., 2024) would be complementary to our approach.
 To formalize the notion of information gathering, we need a concrete idea of what we wish to learn about.
-We denote this target quantity as $\theta$, and note that it could, for example, be a user’s preference, the answers to one or more questions, a recommendation, or a piece of content that a user wants to generate.
-To meaningfully reason about $\theta$, we need to have some belief distribution for it, which we denote $p\left(\theta;h_{t}\right)$, [^1] Note we carefully distinguish between explicit conditioning, i.e.
-$p(a|b)$, and more general dependency, $p(a;b)$, in our notation.
-The former explicitly corresponds to the conditional distribution of an associated joint distribution, $p(a,b)$, while the latter may not.
-So here, for example, $h_{t}$ influences our distribution on $\theta$, but we are not deriving this distribution through conditioning of some explicit joint distribution.
-where $h_{t}$ represents the history of queries and responses at time $t$.
-As more queries are made, this history will grow and our beliefs on $\theta$ will update accordingly.
-We note that $p\left(\theta;h_{t}\right)$ will typically not be defined explicitly.
-Even in the LLM setting, there is significant flexibility to how it is constructed, as we discuss in Sections 3.3 and 4.
-2.1 Information Gathering with LLMs
-Our focus is on how to use LLMs to gather information about our target quantity $\theta$.
-The precise set of problems this can relate to is quite broad and includes, to name but a few, task clarification (Zhang et al, 2025a, ), preference learning (Handa et al,, 2024; Chakraborty et al,, 2024; Ouyang et al,, 2022), diagnosis and troubleshooting (Hirosawa et al,, 2024; Jha et al,, 2025), automating external simulations or real-world experiments (Mandal et al,, 2025), conducting surveys or collecting feedback (Aher et al,, 2023; Lee et al,, 2024), and AI tutoring agents (Kestin et al,, 2025; Liu et al, 2024a, ).
-What all of these have in common is that the LLM starts with incomplete information about the target $\theta$, but can refine its beliefs by making queries, $x\in\mathcal{X}$, to the user or some other external agent and receiving responses, $y\in\mathcal{Y}$, that are informative about $\theta$.
-In general, multiple such queries, $x_{1},\dots,x_{n}$, can be made.
-We can choose them adaptively in a sequential decision-making process where we iteratively choose each $x_{t+1}$ based on the collected history $h_{t}:=(x_{i},y_{i})_{i=1:t}$.
-For clarity of exposition, we focus on the case where the $x_{i}$ correspond to explicit questions asked to the user, but emphasize that the approach applies more broadly to other forms of external interaction by the LLM, such as retrieving documents or calling external functions.
-The simplest and cheapest way to incorporate the interaction history into the LLM to aid our sequential decision making is to include it in the context (Brown et al,, 2020).
-Namely, if the LLM’s distribution over generated text, $z\in\mathcal{Z}$, is $p_{\mathrm{LLM}}(z;c)$, where $c$ is an initial prompt, then we can consider $p_{\mathrm{LLM}}(z;[c,h_{t}])$ as an updated LLM that uses the previous question–response pairs as in-context examples, where $c,h_{t}]$ represents some appropriate concatenation of the history into the context.
+We denote the target quantity of interest as $\theta$, which may represent, for example, a user’s preferences, the answer to a question, or a desired piece of content.
+We start with incomplete information about $\theta$, as represented by an initial belief distribution or prior, $p(\theta)$, but can refine these beliefs by making queries, $x\in\mathcal{X}$, to the user or some other external agent and receiving responses, $y\in\mathcal{Y}$, that are informative about $\theta$.
+Multiple such queries, $x_{1},\dots,x_{n}$, can be adaptively selected in a sequential decision-making process where we iteratively choose each $x_{t}$ based on the collected history $h_{t-1}:=(x_{i},y_{i})_{i=1:t-1}$.
+As our history grows, we will update our belief distribution to obtain $p(\theta;h_{t-1})$ via some model updating procedure.
+[^1] We carefully distinguish between explicit probabilistic conditioning, i.e.
+$p(a|b)$, and more general dependency, $p(a;b)$.
+The former corresponds to the conditional distribution of an associated joint distribution, $p(a,b)$, while the latter may not.
+Here, $h_{t-1}$ influences our distribution on $\theta$, but it is not derived via a joint distribution.
+In the LLM setting, there is considerable flexibility in how $p(\theta;h_{t-1})$ is constructed, as discussed in §3.3 and §4.
+While $p(\theta;h_{t-1})$ need not be explicitly defined, it provides the foundation for our information-theoretic method of query selection.
+For clarity of exposition, we focus on the case where the $x_{t}$ correspond to explicit questions asked to the user, but emphasize that the approach applies more broadly to other forms of external interaction by the LLM, such as retrieving documents or calling external functions.
+2.1 In–Context Updating of the belief distribution
+A natural and cheap way to incorporate the interaction history into the LLM is to include it in the context (Brown et al., 2020).
+If the LLM’s distribution over generated text, $z\!\in\!\mathcal{Z}$, is $p_{\mathrm{LLM}}(z)$ given appropriate prompting, then $p_{\mathrm{LLM}}(z;h_{t-1})$ is an updated distribution with the previous question–response pairs in context.
 From this, we can derive an updated belief distribution over $\theta$.
-Most simply, this can be done by using $p_{\mathrm{LLM}}(z;[c,h_{t}])$ to directly query about $\theta$ (e.g.
+Most simply, this can be done by using $p_{\mathrm{LLM}}(z;h_{t-1})$ to directly query about $\theta$ (e.g.
 if $\theta$ is some preference, we could prompt the LLM to predict this preference).
-However, as we show later, this approach often fails to appropriately or fully incorporate the information from $h_{t}$, leading to belief states that are inconsistent with past observations.
-We will introduce a more robust and sophisticated method for deriving $p(\theta;h_{t})$ from $p_{\mathrm{LLM}}(z;[c,h_{t}])$ in Section˜3.3.
-In the interest of brevity, for the rest of the paper we will drop the initial prompt $c$ from our notation and also assume that an appropriate final prompt has been added to end of the context to predict/sample our quantity of interest.
-For example, we will use $p_{\mathrm{LLM}}(\theta;h_{t})$ to denote the distribution when our context is $c,h_{t}]$ followed by an appropriate prompt to sample $\theta$ as its next token(s).
+However, as we show later, this approach often fails to appropriately incorporate the information from $h_{t-1}$, leading to a belief distribution inconsistent with past observations.
+This is consistent with recent work that shows that in context updating does not treat all contextual information equally (Kossen et al., 2024; Liu et al., 2024b; Zhang et al., 2024).
+In §3.3, we introduce a more robust method for deriving $p(\theta;h_{t-1})$.
 2.2 Information–Theoretic Experimental Design
-Our aim of maximizing the information gathered about a target quantity can be formalized using the framework of Bayesian experimental design (BED; Lindley,, 1956; 1972; Chaloner & Verdinelli,, 1995; Rainforth et al,, 2024), which provides a powerful and principled framework for data acquisition.
-The core of the BED framework is a joint generative model, $p(\theta,y;x)$ over the target quantity $\theta$ and outcomes, $y$, given designs $x$.
-Traditionally, this is most commonly specified as a conventional Bayesian model using a prior $p\left(\theta\right)$ and likelihood $p(y|\theta;x)$ (Rainforth et al,, 2024), but it can also be directly defined using a stochastic simulator (Kleinegesse & Gutmann,, 2020; Rainforth,, 2017) or as the marginal of some larger model (e.g.
-because $\theta$ corresponds to something other than the model parameters (Bickford Smith et al,, 2023; MacKay,, 1992)).
-In the most general case, designs are then chosen to maximize the expectation of some utility function $U(\theta,y,x)$ under this generative model; that is we take $x^{*}=\mathrm{argmax}_{x}\mathbb{E}_{p(\theta,y;x)}\left[U(\theta,y,x)\right]$.
-The most common choice is to take $U(\theta,y,x)=\log p(\theta,y;x)-\log p\left(\theta\right)\log p(y;x)$ where $p\left(\theta\right)$ and $p(y;x)$ are the marginal distributions on $\theta$ and $y$ implied by our joint model and we have assumed that our current beliefs on $\theta$ are independent of the design.
-This leads to an objective corresponding to the expected information gain (EIG) in $\theta$ (Lindley,, 1956; 1972),
-$\mathrm{EIG}_{\theta}(x)$ $=\mathbb{E}_{p(\theta,y;x)}\left[\log\frac{p(\theta,y;x)}{p\left(\theta\right)p(y;x)}\right],$ (1)
-which using Bayes’ rule can also be represented in the following equivalent forms
-$\text{EIG}_{\theta}(x)$ $=\mathrm{H}[p\left(\theta\right)]-\mathbb{E}_{p(y;x)}[\mathrm{H}[p(\theta|y;x)]]$ (2)
-$=\mathrm{H}[p(y;x)]-\mathbb{E}_{p\left(\theta\right)}[\mathrm{H}[p(y|\theta;x)]],$ (3)
-where $\mathrm{H}$ denotes the Shannon entropy (i.e., $\mathrm{H}[p\left(\theta\right)]=-\mathbb{E}_{p\left(\theta\right)}[\log p\left(\theta\right)]$).
-We can thus equivalently think of this EIG as the mutual information between $\theta$ and $y$, the expected information gain over possible data simulated from our model (where the information gain is defined as the reduction in entropy from our prior on $\theta$ to the posterior), and the expected reduction in entropy over data from observing $\theta$ simulated from our prior (Sebastiani & Wynn,, 2000).
-By maximizing this objective with respect to $x$, we are thus maximizing the expected amount of information we gain in $\theta$, subject to the assumptions that our generative model can both accurately estimate the true data-generating process and encapsulate our beliefs about $\theta$.
-Working with the EIG is highly suited to a sequential or adaptive design approach, wherein it is generally referred to as sequential BED or Bayesian Adaptive Design (Rainforth et al,, 2024).
+The core of the BED framework is a joint generative model $p(\theta,y;x)$ over the target quantity $\theta$ and outcomes, $y$, given designs $x$.
+Most commonly, this is specified as a Bayesian model using a prior $p(\theta)$ and likelihood $p(y|\theta;x)$.
+In the general case, designs are then chosen to maximize the expectation of some utility function $U(\theta,y,x)$ under this model: we choose $x^{*}=\mathrm{argmax}_{x}\mathbb{E}_{p(\theta,y;x)}\left[U(\theta,y,x)\right]$.
+The most common choice is to take $U(\theta,y,x)=\log p(\theta,y;x)-\log p(\theta)\log p(y;x)$, where $p(\theta)$ and $p(y;x)$ are the marginal distributions on $\theta$ and $y$ implied by our joint model and we have assumed that our current beliefs on $\theta$ are independent of the design $x$.
+This leads to an objective corresponding to the expected information gain (EIG) in $\theta$ (Lindley, 1956; 1972),
+$\text{EIG}_{\theta}(x)$ $=\mathrm{H}[p(\theta)]-\mathbb{E}_{p(y;x)}[\mathrm{H}[p(\theta|y;x)]]$ (1)
+$=\mathrm{H}[p(y;x)]-\mathbb{E}_{p(\theta)}[\mathrm{H}[p(y|\theta;x)]],$ (2)
+where $\mathrm{H}$ denotes the Shannon entropy (i.e., $\mathrm{H}[p(\theta)]=-\mathbb{E}_{p(\theta)}[\log p(\theta)]$).
+We can thus equivalently think of the EIG as: a) the mutual information between $\theta$ and $y$, b) the expected information gain over possible data simulated from our model (where the information gain is defined as the reduction in entropy from our prior on $\theta$ to the posterior), or c) the expected reduction in entropy over data from observing $\theta$ simulated from our prior (Sebastiani & Wynn, 2000).
+Working with the EIG is highly suited to a sequential or adaptive design approach, wherein it is generally referred to as sequential BED or Bayesian Adaptive Design (Rainforth et al., 2024).
 Because the EIG is only a function of our underlying model, when we update the model as new data becomes available, our EIG design objective will naturally update as well.
-In particular, if $p(\theta,y_{t+1};h_{t},x_{t+1})$ represents our generative model after already observing the history $h_{t}$, then we can iteratively optimize the incremental EIG (MacKay,, 1992; Cavagnaro et al,, 2010),
-$\mathrm{EIG}_{\theta}(x_{t+1};h_{t})$ $=\mathbb{E}_{p(\theta,y_{t+1};h_{t},x_{t+1})}\left[\log\frac{p(\theta,y_{t+1};h_{t},x_{t+1})}{p\left(\theta;h_{t}\right)p(y_{t+1};h_{t},x_{t+1})}\right].$ (4)
-Here our updated model, $p(\theta,y_{t+1};h_{t},x_{t+1})$, conventionally comes from a Bayesian update of the original model, noting that the Savage axioms dictate that a rational agent should update its beliefs in a Bayesian manner (Savage,, 1954).
-However, in many cases, this is not practical and other non-Bayesian updates are performed instead.
-In particular, in many Bayesian active learning problems, the update actually corresponds to retraining the model with an updated dataset (Gal et al,, 2017; Bickford Smith et al,, 2023).
-It has recently been noted that this traditional sequential BED approach can sometimes be suboptimal in practice, as it only optimizes the EIG of the next observation, without planning ahead for the fact that design decisions taken at a given step can also influence the achievable EIGs from future steps (Foster,, 2021).
-A variety of policy–based BED approaches have subsequently been proposed to address this (Foster et al,, 2021; Ivanova et al,, 2021; Blau et al,, 2022; Huan & Marzouk,, 2016; Hedman et al,, 2025), while also removing the need to make model updates and conduct optimizations during the experiment itself.
-However, training these policies typically requires a very large number of simulations from the generative model, which makes them impractical for our setting, due to the expense of simulating full experimental rollouts from the LLM.
+Specifically, to derive the incremental EIG (Cavagnaro et al., 2010) for the $t$-th query, $\text{EIG}_{\theta}(x_{t};h_{t-1})$, we simply replace the joint $p(\theta,y;x)$ in the above formulation with the updated joint $p(\theta,y_{t};h_{t-1},x_{t})$, with all marginals a conditionals derived from this (e.g.
+$p(y;x)$ becomes $p(y_{t};h_{t-1},x_{t})$).
+Here this updated joint conventionally comes from a Bayesian update of the original model.
+However, in many cases, this is not practical and other non-Bayesian updates are performed instead, e.g.
+in active learning the update often actually corresponds to retraining the model with the new data (Gal et al., 2017; Bickford Smith et al., 2023).
 3 Sequential Bayesian Experimental Design for LLMs
-The sequential BED framework described in Section˜2.2 requires two core components to be specified by the user: an initial joint model $p(\theta,y;x)$ over hypotheses $\theta$ and outcomes $y$, given chosen experiment(s) $x$; and a procedure to derive an updated model $p(\theta,y_{t+1};h_{t},x_{t+1})$ after observing an experimental history $h_{t}$.
+The sequential BED framework described in §2.2 requires two core components to be specified by the user: a) an initial joint model $p(\theta,y;x)$ over hypotheses $\theta$ and outcomes $y$, given chosen experiment $x$, and b) a procedure to derive an updated model $p(\theta,y_{t};h_{t-1},x_{t})$ after observing $h_{t-1}$.
 In the LLM setting, there is significant flexibility in these critical design decisions.
-In particular, there are various ways that one can derive a suitable joint distribution from the LLM and the ability to learn in context provides opportunities for update methods that go beyond standard Bayesian model updates.
+In particular, there are many ways to derive a suitable joint distribution from the LLM and its ability to learn in-context provides opportunities for update methods that go beyond standard Bayesian model updates.
 Model Construction
 A major challenge in the LLM setting is that unlike conventional probabilistic models, in general, $p_{\mathrm{LLM}}(\theta)\,p_{\mathrm{LLM}}(y;[\theta,x])\neq p_{\mathrm{LLM}}(y;x)\,p_{\mathrm{LLM}}(\theta;[x,y])$.
-That is, we induce a different joint distribution if we first sample $\theta$ then sample $y$ with $\theta$ in context (which we refer to as the prior–likelihood pairing), than if we first sample $y$ then sample $\theta$ with $y$ in context (which we refer to as the data–estimation pairing).
-Moreover, in each pairing we may wish to deviate from the LLM’s directly induced distribution on one or both components.
-The success of using sequential BED with LLMs turns out to be critically dependent on which pairing we choose and the deviations we make from LLM’s directly induced distribution.
-We delay proper discussion of this complex issue until Section˜4, where we will see that the preferable setup can depend on the precise problem setting and, in particular, the relative complexity of spaces of $\theta$ and $y$.
-For now, we will focus on using the prior–likelihood pairing, as this transpires to generally be the advantageous setup when the space of $\theta$ is more complex than that of $y$, which we believe to be the more commonly occurring setting and the one we consider in our experiments.
+That is, we induce a different joint distribution if we first sample $\theta$ then sample $y$ with $\theta$ in context (which we refer to as the prior–likelihood pairing), than if we first sample $y$ then sample $\theta$ with $y$ in context (data–estimation pairing).
+Moreover, we can deviate from the distribution directly induced by the LLM on one or both variables.
+The success of using BED with LLMs turns out to be critically dependent on these choices.
+We delay proper discussion of this complex issue until §4, where we will see that the preferable setup can depend on problem setting and, in particular, the relative complexity of spaces of $\theta$ and $y$.
+For now, we will focus on using the prior–likelihood pairing; we will argue in §4 that this is the advantageous setup in many practical scenarios.
 While we will generally use the LLM’s directly induced distribution for the likelihood, we allow the prior to deviate from this in a problem–specific manner.
-As such, our initial joint model will be $p(\theta,y;x)=p(\theta)p_{\mathrm{LLM}}(y;[\theta,x])$, where the specification of the prior $p(\theta)$ can vary between problems.
+As such, our initial joint model will be $p(\theta,y;x)=p(\theta)p_{\mathrm{LLM}}(y;[\theta,x])$.
 Model Updating
-Optimally updating the joint model in this setting requires incorporating new observations in a way that both fully captures the available information from new data and is computationally tractable.
+Optimally updating the joint model in this setting requires incorporating new observations in a way that both fully captures the information from new data and is computationally tractable.
 At one extreme, we could perform full Bayesian updates via approximate inference, as in classical sequential BED.
-This approach, however, demands a prohibitively large number of LLM evaluations to accurately approximate the posterior, and it does not exploit the power of the LLM as a probabilistic generative model, where autoregressive sequential rollouts often lead to more nuanced and diverse behavior than repeated static likelihood queries.
-The other obvious choice is to rely on simple in context updates, such that we use $p(\theta;h_{t})=p_{\mathrm{LLM}}(\theta;h_{t})$.
-However, as we show later, this fails to sufficiently capture the information provided by the new data and thus undermines the sequential BED approach.
-As we will discuss in Section˜3.3, we therefore employ a strategy that is somewhere between the two, wherein we set up a rejection sampling procedure that draws samples from the in–context distribution, $p_{\mathrm{LLM}}(\theta;h_{t})$, but then filters out samples that are actually not compatible with previous observations.
-We refer to the resulting distribution as $p_{f}(\theta;h_{t})$.
-In our experiments and subsequent exposition of our approach, we will not update the likelihood model as more data is gathered, that is, our likelihood in the sequential setting will be $p_{\mathrm{LLM}}\left(y_{t+1};\left[\theta,x_{t+1}\right]\right)$ instead of $p_{\mathrm{LLM}}(y_{t+1};[h_{t},\theta,x_{t+1}])$.
-The main rationale of this choice is that for many problems our beliefs on $\theta$ capture all the required information to predict $y|x$, hence including the history is adding unnecessary context that could influence the LLM’s behavior in undesirable ways.
-However, it is important to note that $p_{\mathrm{LLM}}(y_{t+1};[h_{t},\theta,x_{t+1}])$ should be used instead for problems where $\theta$ will not capture all information from previous data, e.g.
-if $\theta$ is a binary value corresponding to whether we reject a null hypothesis, or is the answer to a particular other question of interest.
+However, this demands a prohibitively large number of LLM evaluations to accurately approximate the posterior, and it does not exploit the power of the LLM as a probabilistic generative model, where autoregressive sequential rollouts often lead to more nuanced and diverse behavior than repeated static likelihood queries.
+At the other extreme, simple in-context updating, $p(\theta;h_{t-1})=p_{\mathrm{LLM}}(\theta;h_{t-1})$, is cheap but, as we show later, fails to reliably capture information from new data, leading to inconsistent belief states and undermining the sequential BED approach.
+As we discuss in §3.3, we therefore employ a strategy that is somewhere between the two: drawing samples in a way that utilizes $p_{\mathrm{LLM}}(\theta;h_{t-1})$ while encouraging diversity, then filtering out samples that are actually not compatible with $h_{t-1}$ and renormalizing.
+We refer to the resulting distribution as $p_{f}(\theta;h_{t-1})$.
+We do not update our likelihood model $p_{\mathrm{LLM}}(\theta;h_{t-1})$; see Section˜A.1 for a discussion.
 BED-LLM
-We now introduce our specific principled algorithm for intelligent and adaptive information gathering in LLMs using sequential BED.
-We call our approach BED-LLM, which stands for Bayesian Experimental Design with Large Language Models.
-Here, the questions posed to the user [^2] More broadly, these could be more general queries such as document retrieval or function calling; however in our experiments we only consider questions asked of a user.
-will correspond to our designs, $x$, and the responses received will correspond to our outcomes, $y$.
-Using the LLM to derive a joint model over these outcomes and the target variables $\theta$ as described above, we can derive a precise form of the expected information gain, $\text{EIG}_{\theta}(x)$.
-We can then interleave choosing informative questions by optimising the incremental EIG defined in Equation˜4, and updating our underlying model based on the received question–response pairs.
-Specifically, BED-LLM iterates over the following key steps, where $t$ indexes the current turn.
+We now introduce our specific algorithmic approach, BED-LLM.
+Here, the queries will correspond to our designs, $x$ (assumed to be in form of questions posed to the user in the following for simplicity, but could also be, e.g., external function calls, document retrieval, web search, etc), and the responses received will correspond to our outcomes, $y$.
+Using the LLM to derive joint models over these outcomes and the target variables $\theta$ given histories, $h_{t-1}$ as described above, we can interleave choosing informative questions by optimizing the incremental EIG, $\text{EIG}_{\theta}(x_{t};h_{t-1})$, and updating our underlying model based on the received question–response pairs.
+Specifically, BED-LLM iterates over the following key steps, where $t$ indexes the current turn:
   1. 1.
-    Generate Candidate Questions: Propose a candidate set of $M$ diverse, multiple-choice questions, $\mathcal{X}^{\text{cand}}$, by appropriate sampling of the LLM based on the current conversational context (Section˜3.1).
+    Generate candidate questions3.1): Propose a candidate set of $M$ diverse, multiple-choice questions, $\mathcal{X}^{\text{cand}}$, by appropriate sampling of the LLM based on the conversational context $h_{t-1}$.
   2. 2.
-    Compute EIG Estimator: For each candidate question $x\in\mathcal{X}^{\text{cand}}$, estimate $\text{EIG}_{\theta}(x)$ as described in in Section˜3.2.
+    Compute EIG estimator3.2): For each candidate $x_{t}\in\mathcal{X}^{\text{cand}}$, estimate $\text{EIG}_{\theta}(x_{t};h_{t-1})$.
   3. 3.
@@
   4. 4.
-    Construct updated joint, $p(\theta,y_{t+1};h_{t},x_{t+1})=p_{f}(\theta;h_{t})p_{\mathrm{LLM}}\left(y_{t+1};\left[\theta,x_{t+1}\right]\right)$, using the new history (Section˜3.3) and return to Step 1 (unless a termination criterion like maximum number of questions has been achieved).
-Because BED-LLM uses the prior–likelihood pairing, our belief state on $\theta$ is simply given by $p_{f}(\theta;h_{t})$ at any given point in the process; our underlying aim is that the uncertainty of this distribution contracts as quickly as possible.
-In the remainder of this section, we provide details on these key individual steps.
-We note that some of the choices made are closely linked to our joint model construction; we will later provide extensive discussion on why this model construction was chosen and possible alternatives in Section˜4.
-3.1 Candidate Question Generation
-The first step of each BED-LLM iteration is to generate a finite pool of candidate questions, $\mathcal{X}^{\text{cand}}$, from which our chosen question will be selected.
-To ensure we ultimately choose good questions, we must generate a candidate set $\mathcal{X}^{\text{cand}}$ that is rich and diverse enough to include sufficiently informative questions in the first place.
-We consider two approaches for generating candidate questions:
-  • Unconstrained generation. Given the accumulated history, the LLM is simply asked to propose new questions by sampling from $p_{\mathrm{LLM}}(x_{t+1};h_{t})$ with appropriate prompting.
-  • Conditional generation.
-The LLM is given both the accumulated history and a generated set of hypotheses $\Theta^{\text{cand}}=\{\theta^{(n)}\}_{n=1}^{N}$, such that we sample from $p_{\mathrm{LLM}}(x_{t+1};[h_{t},\Theta^{\text{cand}}])$.
-Specifically, the LLM is prompted to propose questions whose answers would “slice” the hypothesis pool into roughly balanced subsets.
-For both strategies, we sample $M$ questions independently with a relatively high temperature (we use $T=1.3$ in our experiments), in order to encourage diversity.
+    Construct updated joint (Section˜3.3): $p(\theta,y_{t+1};h_{t},x_{t+1})=p_{f}(\theta;h_{t})p_{\mathrm{LLM}}(y_{t+1};[\theta,x_{t+1}])$, using the new history and return to Step 1 (unless a termination criterion has been achieved).
+Note our belief state on $\theta$ after the $t$-th turn is simply given by $p_{f}(\theta;h_{t})$.
+3.1 Generating Candidate Questions
+As it is not computationally feasible to directly optimize over the space of possible questions, we rely on using the LLM to propose diverse candidate questions, $\mathcal{X}^{\text{cand}}$, then select the best question from these.
+We consider two specific approaches: 1) Unconstrained generation.
+Given $h_{t-1}$, the LLM is simply asked to propose new questions by sampling from $p_{\mathrm{LLM}}(x_{t};h_{t-1})$ with appropriate prompting.
+2) Conditional generation.
+The LLM is given both $h_{t-1}$ and a generated set of hypotheses $\Theta^{\text{cand}}=\{\theta^{(n)}\}_{n=1}^{N}$, such that we sample from $p_{\mathrm{LLM}}(x_{t};[h_{t-1},\Theta^{\text{cand}}])$.
+Specifically, the LLM is prompted to propose questions that “slice” the hypothesis pool into roughly balanced subsets.
+For both strategies, we sample $M$ questions jointly with a relatively high temperature to encourage diversity.
 Conditional generation allows us to “guide” the LLM to propose highly informative questions.
-However, it risks over-fitting to the sampled set of hypotheses.
-In practice, we found that conditional generation works well for discrete, easily divisible hypothesis spaces such as those found in the 20–Questions setting, but tends to overfit in the case of more complex, overlapping hypotheses such as persona identification.
-We can avoid this overfitting by generating two hypothesis sets, one for generating questions and one for estimating EIG.
-However, this adds significant computational cost, leading us to prefer the unconstrained generation approach in such cases.
-We restrict our questions to be multiple-choice questions.
-This is because having a small space of single-token responses $y$ (e.g., yes/no, A/B/C/D) ensures that $p_{\mathrm{LLM}}\left(y_{t+1};\left[\theta,x_{t+1}\right]\right)$ provides a meaningful uncertainty over outcomes and that we can reliably calculate/estimate its entropy.
-We discuss this restriction at length in Section˜4.2, where we will see that limitations in LLM uncertainty quantification over unrestricted spaces generally require us to use a restricted space for $y$ (as done here) or $\theta$ (which can be tackled by instead using the data–estimation pairing).
+However, it risks overfitting to $\Theta^{\text{cand}}$.
+In practice, we find it is effective for discrete spaces (§6.1), but less so for spaces with complex, overlapping hypotheses (§6.2).
+We restrict questions to multiple-choice format to allow $p_{\mathrm{LLM}}\left(y_{t};\left[\theta,x_{t}\right]\right)$ to produce well-calibrated probabilities (see §4).
 3.2 Estimating EIG for each question
-We can estimate the EIG for a given question $x_{t+1}$ using the following Rao-Blackwellized estimator of Equation 3, independently introduced by Gal et al, (2017) and Rainforth, (2017):
-$\begin{split}\text{EIG}_{\theta}(x_{t+1};h_{t})\approx&\frac{1}{N}\sum_{n=1}^{N}\sum_{y_{t+1}\in\mathcal{Y}}p_{\mathrm{LLM}}(y_{t+1};[\theta^{(n)},x_{t+1}])\log p_{\mathrm{LLM}}(y_{t+1};[\theta^{(n)},x_{t+1}])\\ &-\sum_{y_{t+1}\in\mathcal{Y}}\hat{p}(y_{t+1};[h_{t},x_{t+1}])\log\hat{p}(y_{t+1};[h_{t},x_{t+1}]),\end{split}$ (5)
-where $\hat{p}(y_{t+1};[h_{t},x_{t+1}])\!:=\!\frac{1}{N}\sum_{n=1}^{N}p_{\mathrm{LLM}}(y_{t+1};[\theta^{(n)},x_{t+1}])$ and $\theta^{(n)}\sim p_{f}(\theta;h_{t})$ (see Section˜3.3).
-While this estimator is biased for finite NN, [^3] In general it is not possible to derive unbiased estimates for the EIG without using MLMC approaches (Goda et al,, 2022) that significantly increase the cost-normalized variance.
-it is still a singly intractable estimator that converges as $N\to\infty$ at the standard Monte Carlo rate $O(1/\sqrt{N})$ (Rainforth et al,, 2018).
-Note that the samples $\theta^{(n)}$ do not need to be independent for this to hold, provided that they can be shown to satisfy some appropriate form of ergodicity or decaying correlation (see, e.g., Billingsley, (2013)).
-Note that when constructing this estimator, one should directly utilize the LLM’s logits (whenever this is possible) to calculate the $p_{\mathrm{LLM}}(y_{t+1};[\theta^{(n)},x_{t+1}])$ terms.
-By the Rao-Blackwell theorem (Rao et al,, 1945), this will always produce lower variance estimators than purely sample–based estimators, like those employed in Hu et al, (2024) and Kobalczyk et al, (2025).
-3.2.1 EIG cannot be well approximated using predictive entropy
-Previous attempts to apply BED to the LLM setting with a prior–likelihood based model have all, either explicitly or implicitly, assumed that the expected likelihood entropy, $\mathbb{E}_{p(\theta;h_{t})}[\mathrm{H}[p(y_{t+1}|\theta;x_{t+1},h_{t})]]$, is constant over designs (Kobalczyk et al,, 2025; Mazzaccara et al,, 2024; Cooper et al,, 2025); for BED-LLM this would equate to ignoring the first line of the estimator in Equation˜5.
-Under this assumption, the EIG breaks down to simply the marginal predictive entropy, $\mathrm{H}[\mathbb{E}_{p(\theta;h_{t})}[p(y_{t+1}|\theta;x_{t+1},h_{t})]]$.
-Therefore, approaches that make this assumption are no longer information–based, and are instead making decisions simply based on predictive uncertainty.
-This is problematic as, in practice, this expected conditional uncertainty will vary between questions.
+To estimate the EIG based on Equation 2 for a given question $x_{t}$, we derive the following Rao-Blackwellized estimator based on the LLM’s predictive distribution:
+$\begin{split}\text{EIG}_{\theta}(x_{t};h_{t-1})\approx&\textstyle\frac{1}{N}\sum_{n=1}^{N}\sum_{y_{t}\in\mathcal{Y}}p_{\mathrm{LLM}}(y_{t};[\theta^{(n)},x_{t}])\log p_{\mathrm{LLM}}(y_{t};[\theta^{(n)},x_{t}])\\ &-\textstyle\sum_{y_{t}\in\mathcal{Y}}\hat{p}(y_{t};[h_{t-1},x_{t}])\log\hat{p}(y_{t};[h_{t-1},x_{t}]),\end{split}$ (3)
+where $\hat{p}(y_{t};[h_{t-1},x_{t}])\!:=\!\frac{1}{N}\sum_{n=1}^{N}p_{\mathrm{LLM}}(y_{t};[\theta^{(n)},x_{t}])$ and $\theta^{(n)}\!\sim\!p_{f}(\theta;h_{t-1})$, see §3.3.
+This estimator has been used in other BED contexts (Gal et al., 2017; Rainforth, 2017).
+Note that the samples do not need to be independent for this estimator to converge, provided they satisfy some appropriate form of ergodicity or decaying correlation (see, e.g., Billingsley (2013)).
+When constructing this estimator, we compute the $p_{\mathrm{LLM}}(y_{t};[\theta^{(n)},x_{t}])$ terms using the LLM’s logits whenever possible.
+By the Rao-Blackwell theorem, this always produces lower variance than purely sample–based estimators (Rao et al., 1945), like those employed in Hu et al.
+(2024) and Kobalczyk et al.
+(2025).
+Avoiding deterministic likelihood assumptions
+Previous attempts to apply information criteria to choosing queries in LLMs have generally assumed responses are deterministic given $(\theta,x_{t})$ (Cooper et al., 2025; Kobalczyk et al., 2025; Hu et al., 2024; Mazzaccara et al., 2024; Piriyakulkij et al., 2023).
+Under this assumption, the EIG simplifies to the marginal predictive entropy, $\mathrm{H}[p(y_{t};x_{t},h_{t-1})]$.
+This is problematic as, in practice, the expected likelihood entropy will vary with $x_{t}$.
 For example, if $\theta=$“Dog”, then the response to the question “Is it an animal?” should be close to deterministic, but the expected response to the question “Does it have black fur?” clearly is not.
-In general, $\mathbb{E}_{p(\theta;h_{t})}[\mathrm{H}[p(y_{t+1}|\theta;x_{t+1},h_{t})]]$ measures how easily the question can be answered once $\theta$ is known, and intuitively this will not be constant across questions for the vast majority of problems.
-This variability in the expected entropy of the likelihood can be crucial in determining good questions and omitting it provides no meaningful computational saving (as the same LLM evaluations are used for the top and bottom lines of Equation˜5, doing the full estimate of the EIG requires no additional LLM calls to be made).
-We, therefore, strongly advocate against making this assumption: including the expected entropy of the likelihood is essential in avoiding questions that are irrelevant, ambiguous, unclear, or simply unhelpful in our quest to learn about $\theta$.
-This recommendation will be strongly backed up by our experiments, where we will see that approximating the EIG with the marginal predictive entropy significantly degrades performance of the sequential BED approach.
-A further example of how predictive entropy can be detrimental is given in Figure˜1.
-One might instead be tempted to replace $\hat{p}(y_{t+1};[h_{t},x_{t+1}])$ with $p_{\mathrm{LLM}}(y_{t+1};[h_{t},x_{t+1}])$ in the EIG estimation, as the two essentially offer alternative predictive distributions for the outcome.
-We also advise against this though, noting that it again provides no meaningful computational benefits (unless one also assumes a deterministic likelihood, but this would then mean we no longer consider $\theta$ at all).
-A key reason for avoiding this substitution is that it would mean we are no longer estimating a true EIG: the inconsistency between the likelihood and the marginal data distribution means there is no longer a joint model where we are minimising our expected uncertainty in $\theta$.
-We also find that the LLM process of sampling $\theta$ from $p_{f}(\theta;h_{t})$ followed by $y$ from $p_{\mathrm{LLM}}\left(y_{t+1};\left[\theta,x_{t+1}\right]\right)$ tends to give a better uncertainty over responses than sampling $y$ directly from $p_{\mathrm{LLM}}(y_{t+1};[h_{t},x_{t+1}])$.
+In general, $\mathbb{E}_{p(\theta;h_{t-1})}[\mathrm{H}[p(y_{t}|\theta;x_{t},h_{t-1})]]$ measures how certainly the question can be answered once $\theta$ is known.
+Including it in our objective is essential in avoiding questions that are irrelevant, ambiguous, unclear, or simply unhelpful in our quest to learn about $\theta$.
+We provide an illustrative example of this in Section˜A.2.1 and empirical evidence for this in §6.
+Given that approximating the EIG with marginal predictive entropy does not provide meaningful computational savings (as it does not reduce the required number of LLM calls), we advise against making such deterministic likelihood assumptions.
 3.3 Prior Construction and Belief Updating
-The Savage axioms (Savage,, 1954) tell us that a rational agent should update its beliefs in a Bayesian manner.
+The Savage axioms (Savage, 1954) tell us that a rational agent should update its beliefs in a Bayesian manner.
 However, doing full Bayesian updates to our model as the history grows is generally impractical for computational reasons in the LLM setting, as it requires approximate inference and this, in turn, typically requires large numbers of expensive likelihood evaluations.
-Furthermore, the Savage axioms only hold if our (implied) prior truly represents our beliefs, but we find that $p_{\mathrm{LLM}}(\theta)$ is typically heavily over–collapsed to a small number of possible hypotheses and can struggle to convey the full range of possibilities even with careful prompting and a high temperature (c.f.
-Figure˜8).
+Furthermore, the Savage axioms only hold if our (implied) prior truly represents our beliefs, but we find that $p_{\mathrm{LLM}}(\theta)$ is typically heavily overconfident on a small number of possible hypotheses and can struggle to convey the full range of possibilities even with careful prompting and a high temperature (c.f.
+Figure˜5).
+A natural tractable alternative is to derive our beliefs through LLM in-context updates, that is, use $p_{\mathrm{LLM}}(\theta;h_{t-1})$, noting that this has been shown to behave differently to Bayesian updating (Falck et al., 2024; Kossen et al., 2024).
+However, we find that even state-of-the-art LLMs such as GPT-4o (OpenAI, 2024) often fail to incorporate history faithfully; they regularly sample hypotheses incompatible with past observations and exhibit premature overconfidence, with both issues becoming more pronounced as $h_{t-1}$ grows.
+We discuss reasons for these shortfalls in Section˜A.3.
+To avoid these shortfalls, we instead propose an approach that balances tractability and faithfulness.
+Although we will still use $p_{\mathrm{LLM}}(\theta;h_{t-1})$ as the basis for deriving our belief state over $\theta$ (i.e.
+our intermediate prior), we make various alterations to effectively incorporate historical information and ensure diversity.
+Our derived distribution, which we refer to as $p_{f}(\theta;h_{t-1})$, differs from $p_{\mathrm{LLM}}(\theta;h_{t-1})$ in two key ways.
+First, we filter the generated hypotheses according to whether they are compatible with the history $h_{t-1}$.
+We do this by using the LLM to zero-shot check the compatibility of each sampled $\theta$ with all the previous question–answer pairs in $h_{t-1}$ (using $p_{\mathrm{LLM}}(y_{i};[\theta,x_{i}])\forall i=1:t-1$) and then rejecting that sample if an incompatibility is found.
+Specifically, a sample is rejected if the likelihood of an observed answer falls below a predefined threshold, chosen to balance robustness to model uncertainty against the need to enforce strict historical coherence.
+To reduce the computational cost of generating and evaluating hypotheses, we further include a hypothesis-retention mechanism: any hypotheses from the previous turn which remain consistent with the most recent question and observation are retained in the hypothesis set without regeneration.
+Second, we make a number of modifications to promote diversity.
+Rather than generate candidates independently, we prompt the LLM to generate batches of candidates using a prompt encouraging diversity.
+After filtering these candidates as above and removing duplicates, we then impose a uniform distribution.
+Details of our exact setup for $p_{f}(\theta;h_{t-1})$ are given in Appendix˜E.
+4 On the Specification of $p(\theta,y_{t};h_{t-1},x_{t})$, and its Implications
+As we described in §3, successfully applying sequential BED in the LLM setting hinges upon how we specify, and update, the joint distribution $p(\theta,y_{t};h_{t-1},x_{t})$.
+In particular, as previously highlighted, there are two distinct ways to derive the joint model from our LLM: using a prior–likelihood pairing, $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}p(\theta;h_{t-1})p\left(y_{t};\left[\theta,x_{t}\right]\right)$}, or a data–estimation pairing, $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}p\left(y_{t};[h_{t-1},x_{t}]\right)p\left(\theta;[h_{t-1},x_{t},y_{t}]\right)$}.
+The first construction mirrors deriving our beliefs about $\theta$ from a conventional Bayesian posterior with a concrete prior and likelihood derived (at least partially) from the LLM, whereas the second has analogies to a marginal-posterior approach (Fong et al., 2023; Falck et al., 2024) in that it that samples hypothetical data and draws inferences on $\theta$ given hypothetical data using in-context learning.
+In our outlined BED-LLM approach, we adopted a prior–likelihood pairing.
+Below, we justify this decision and also discuss certain settings where the data–estimation setup might be preferable instead.
+Modeling flexibility   The most obvious relative merits of the prior–likelihood and data–estimation pairings are in the flexibility in how each term is chosen.
+The prior–likelihood pairing gives us greater flexibility to construct a prior set of beliefs over $\theta$ that is distinct to the LLM’s internal beliefs, as it allows us to directly control this prior by changing $p(\theta;h_{t-1})$, whereas the prior is only implicitly defined in the data–estimation pairing.
+In §3.3 we exploited this flexibility through our definition of $p_{f}(\theta;h_{t-1})$.
+On the other hand, the data–estimation pairing could provide some beneficial flexibility in specifying how the data itself is simulated through changing $p\left(y_{t};[h_{t-1},x_{t}]\right)$, which could, for example, be useful when we have access to external data simulators.
+Faithfulness of conditional distributions   While deviations from relying on direct LLM predictions are also in principle possible for the conditional models $p\left(y_{t};\left[\theta,x_{t}\right]\right)$ and $p\left(\theta;[h_{t-1},x_{t},y_{t}]\right)$, in practice, these will typically be more difficult and expensive to implement.
+This is first because these conditionals need to be instantiated for each sampled instance of the conditional variable ($\theta$ and $y_{t}$ respectively), rather than just needing us to set up a single marginal distribution.
+Second, to construct estimators for Equations (1) and (2), we require access to concrete probabilities for the conditional distributions (in order to calculate entropies), whereas we only needed to draw samples for the marginal distributions (in order to approximate expectations).
+As such, the conditionals need to be explicit distributions, or at least ones where the probability can be cheaply estimated, so they are more difficult to define through the output of some algorithmic procedure, especially in large spaces.
+When considering the conditional distributions, the decisive question on the relative merit of the two formulations is which conditional factor we are willing to trust the LLM to supply as a full probability distribution.
+Critically, we rely on how the LLM captures uncertainty in this full distribution—including, for example, tail behavior—not merely the fidelity of typical samples; the marginal factors, by contrast, only need to be sampled from.
+If we accept the LLM’s direct predictive distribution for $p\left(y_{t};\left[\theta,x_{t}\right]\right)$, then we are basing our notion of uncertainty around (and will need to calculate) $\mathrm{H}[p_{\mathrm{LLM}}\left(y_{t};\left[\theta,x_{t}\right]\right)]$, and if instead we place more faith in the LLM’s internal distribution for $p\left(\theta;[h_{t-1},x_{t},y_{t}]\right)$, then we are basing our uncertainty around $\mathrm{H}[p_{\mathrm{LLM}}(\theta;[h_{t-1},x_{t+1},y_{t+1}])]$.
+In essence, the choice between prior–likelihood and data–estimation pairings thus comes down to whether we believe the LLM will produce a more appropriate conditional uncertainty over $\theta$ or $y$, along with our ability to numerically estimate this uncertainty cheaply.
+This difference becomes particularly noticeable when the complexities of the spaces of $\theta$ and $y$ differ significantly.
+Our ability to draw sensible samples of either will generally be quite robust to these spaces being complex or high–dimensional; this is where LLMs tend to thrive, effectively generating highly complex outputs in an autoregressive manner.
+However, evaluating the entropy of a distribution becomes dramatically harder as the dimensionality or complexity increases (Acharya et al., 2019; Paninski, 2003), and the entropy of the predictive distribution of an LLM in such cases will not typically provide a sensible measure of uncertainty (Kadavath et al., 2022; Desai & Durrett, 2020).
+As such, the decision between joint formulations should predominantly be based on the complexity of the space of $\theta$ versus that of $y$: we should generally favor the prior–likelihood formulation if $\theta$ is more complex and the data–estimation formulation if $y$ is more complex.
+For the problems that we consider, the space of $y$ is less complex than that of $\theta$, indicating we should, in general, use the prior–likelihood formulation.
+However, in cases where this is not true, the data–estimation formulation may be preferable instead.
+We provide additional discussion on the impact of the chosen pairing on our entropy estimate, plus discussion on the choice of $\theta$, in Appendix˜B.
+Extracting the belief state   A further advantage of the prior–likelihood construction is that our belief state on $\theta$ can be extracted directly as $p_{f}(\theta;h_{t-1})$.
+With the data–estimation construction, we would have to estimate the marginal on $\theta$ by integrating $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}p\left(y_{t};[h_{t-1},x_{t}]\right)p\left(\theta;[h_{t-1},x_{t},y_{t}]\right)$} over the synthetic response $y_{t}$.
+Direct access to $p(\theta;h_{t-1})$ is also important to ensure that our current belief state is independent of the next question $x_{t}$, which is both intuitively desirable and theoretically required to be a valid BED approach (Lindley, 1972); data–estimation formulations will generally violate this.
+5 Related Work
+Several works have explored the baseline ability of LLMs to rapidly learn about a parameter of interest by asking questions (Zhang et al., 2024; Li et al., 2025b)—effectively our Naive baseline in §6.
+While these works demonstrate some ability to adaptively construct information-seeking questions, they often fail to extract important information (Li et al., 2025a).
+Some works have further specifically attempted to choose questions based on model-based uncertainty criteria (Piriyakulkij et al., 2023; Hu et al., 2024; Kobalczyk et al., 2025; Mazzaccara et al., 2024; Cooper et al., 2025).
+None of these works provide the same careful consideration of how the underlying joint model should be formulated, which underpins our own work, and they all assume deterministic likelihood models that mean their objectives correspond to a sample-based estimate of marginal predictive entropy in practice, as explained in Section˜3.2.
+In general, these previous works have also required restrictions on the space of allowable hypotheses, $\theta$, and typically require additional assumptions and/or approximations.
+More extensive discussion of related work is given in Appendix˜C.
+6 Experiments
+We now assess how well BED-LLM and alternative information-gathering approaches work in two practical scenarios: 20 Questions, a game in which the player has to guess a target entity and can ask up to 20 yes-no questions about the entity; and preference elicitation, a task in which the agent has to predict a user’s preference profile and can ask five multiple-choice questions to the user.
+Answerer
+We produce answers to the questioner LLM’s questions using a separate answerer LLM.
+The answerer is provided with a ground-truth $\theta^{*}$ (a target entity in 20 Questions or a user profile in preference elicitation) and processes individual questions from the questioner without access to any of the questioner’s context (i.e.
+$h_{t-1}$and the questioner’s prompts).
+We test two questioner-answerer setups, where the two are served by separate instances of the same LLM, or two different LLMs.
+The latter scenario is important because in practice, the answerer will often follow a different distribution than the questioner’s internal model for reasoning about responses, thereby forming a model misspecification.
+Baselines
+We compare BED-LLM against two existing baseline methods: Naive and Split.
+Naive involves prompting the questioner to directly generate an informative next question, without explicit hypotheses generation or computing a data-acquisition objective, and then sampling the question with temperature T=1T\!=\!1; this was explored by Zhang et al.
+(2024).
+Split involves choosing the question that most equally splits a sampled set of hypotheses $\Theta^{\text{cand}}$, which corresponds to maximizing the marginal predictive entropy $\mathrm{H}[p(y_{t};x_{t},h_{t-1})]$ in a model with a deterministic likelihood.
+As such, the methods of Cooper et al.
+(2025), Hu et al.
+(2024), Kobalczyk et al.
+(2025), Mazzaccara et al.
+(2024) and Piriyakulkij et al.
+(2023) can all be viewed as variants of this Split baseline.
+While Split is not applicable to the preference-elicitation scenario, where a deterministic likelihood is not viable, it is to our knowledge the state-of-the-art method for 20 Questions.
+We note that our own Split baseline implementation achieves dramatically better results than reported by, for example, Kobalczyk et al.
+(2025), so this constitutes a very strong baseline relative to previous work.
+On top of this, our implementation of Naive appears to signficantly improve over that of Zhang et al.
+(2024),
+6.1 20 Questions
+We consider three sets of 20 Questions problems: Animals, Celebrities, and Things (See Section˜F.1).
+Each problem set comprises 100 target entities $\{\theta^{*}_{i}\}_{i=1}^{100}$ from a given category.
+The space of possible $\theta$ is large and not explicitly defined or restricted: we do not tell the LLM this set of target entities, so the space of $\theta$ is bounded only by what the LLM can generate.
+We note that by comparison, many previous works have relied on restricted spaces for $\theta$ (Chan et al., 2025; Hu et al., 2024; Piriyakulkij et al., 2023; Wang et al., 2025).
+Success Rate (%)
+Dataset Model Naive Split BED-LLM Entropy Data–Est. ICL Beliefs Impl. Max.
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+Table 1: Success rate (%) for 20 Questions at the end of the game.
+Best result in bold.
+$\pm$ numbers show the standard error of the mean estimated using $\sqrt{p(1-p)/(n-1)}$ where $p$ is the success percentage and $n$ is the number of datapoints.
+This estimator is positively biased and thus conservative.
+Data–Est.
+is not possible to run for GPT models due to limited logprobs support in OpenAI API.
+[…]
+BED-LLM improves over Naive and Split baselines
+Our results in Table˜1 and Figure˜1 show BED-LLM significantly outperforming Naive and Split across all problems and LLMs.
+Particularly notable is that BED-LLM’s final success rate is typically more than double that of Naive, highlighting the big gains that can be achieved by using explicit EIG maximisation instead of implicit LLM reasoning.
+Figure 1: Success rate on 20 Questions: mean $\pm$ standard error across 100 targets per dataset.
+BED-LLM ablations
+In order to understand the importance of BED-LLM’s algorithmic components, we further evaluate four ablations on this problem: Entropy, Data–Estimation, ICL Beliefs and Implicit Maximization.
+Each of these differs from BED-LLM with respect to one algorithmic component.
+Entropy replaces the EIG data-acquisition objective with the marginal predictive entropy $\mathrm{H}[p(y_{t};x_{t},h_{t-1})]$ (Section˜3.2); this contrasts with Split in that it uses $p_{\mathrm{LLM}}(y_{t+1};[\theta,x_{t+1}])$ rather than a deterministic likelihood.
+Implicit Maximization involves sampling candidate questions and prompting the questioner to select the most informative one, without any explicit objective estimation (Section˜3.2).
+ICL Beliefs uses $\theta$ belief updates derived from simple in-context learning, namely $p_{\mathrm{LLM}}(\theta;h_{t-1})$ instead of $p_{f}(\theta;h_{t-1})$, testing the importance of our filtering mechanism (Section˜3.3).
+Data–Estimation uses a data–estimation pairing rather than BED-LLM’s prior–likelihood setup (see Appendix˜D).
+In Table˜1 we see that BED-LLM comfortably outperforms all alternative approaches.
+Notably, Entropy provides the strongest ablation baseline, with performance almost identical (sometimes superior) to Split.
+This shows that the use of a non-deterministic likelihood is beneficial in allowing us to target a proper EIG, rather than because it is particularly detrimental to the marginal predictive entropy itself.
+Prior–likelihood outperforms data–estimation
+Our analysis in §4 is validated by our results: BED-LLM’s prior-likelihood approach substantially outperforms Data–Estimation.
+Data–Estimation still outperforms Naive, but interestingly it performs worse than Entropy, highlighting the importance of estimating uncertainty in the $y$ space instead of $\theta$ space.
+These findings reinforce our claim that the choice of joint-model factorization is a critical algorithmic decision.
+Rejection sampling and explicit EIG maximization are key
+We also see the importance of two other aspects of BED-LLM.
+First, how we produce our beliefs over $\theta$ matters: deriving beliefs using simple in-context learning, as in ICL Beliefs, lead to massive performance drops.
+Second, while BED-LLM’s routines for sampling candidate questions and hypotheses are crucial, they alone are not sufficient: passing the samples to an LLM and prompting it to select the highest-EIG question, as in Implicit Maximization, works much less well than using the samples to explicitly maximize EIG.
+BED-LLM is robust to questioner–answerer mismatch
+Our results in Figure˜1 demonstrate that the benefit of BED-LLM persists even under model misspecification.
+This is important for applicability to real-world users, whose responses will follow a different distribution than the questioner LLM.
+6.2 Preference Elicitation
+Unlike 20 Questions, in which $\theta$ is a concrete entity and most reasonable questions have clear answers, many real-world information-gathering tasks involve more abstract targets and less predictable data generation.
+A key example is learning user preferences, where it may be difficult to explicitly define a concrete closed set of possible $\theta$, and it is also challenging for the LLM to develop appropriate uncertainty estimates.
+To study such a scenario, we evaluate BED-LLM on inferring users’ film preferences.
+Here the target we are trying to gather information about is somewhat abstract, and we have some flexibility in how we define $\theta$ in our joint model.
+Our chosen setup is to define $\theta$ to be a user profile, namely a paragraph of text describing the user’s film preferences, with our answerer model prompted to emulate a user with a given profile; see Appendix˜G for full details.
+We consider 200 different user profiles as the ground truth $\theta^{*}$, but as with 20 Questions this set is never given to the questioner.
+Because Split is not applicable as a baseline here (a deterministic likelihood assumption is clearly unreasonable), we benchmark with the similar Entropy approach instead.
+We also note that data–estimation setup is completely unviable here as well because of the large $\theta$ space.
+At each turn $t\in(0,1,\ldots,5)$ we use the question-answer history as context for generating a list of ten film recommendations.
+This list is then rated in its fit to the user profile using an LLM-as-judge setup (Trivedi et al., 2024; Zhu et al., 2025).
+Specifically, the answerer scores each film on a scale of 1 to 5 (in 0.5 increments), based on how well the film aligns with $\theta^{*}$; this score is output together with a brief justification to increase reliability.
+The films’ scores are not included in $h_{t-1}$.
+Our results in Figure˜2 show that, while Naive is often a strong baseline in this preference-elicitation scenario, BED-LLM is still able to provide a boost over both Naive and Entropy, producing higher-rated film recommendations.
+BED-LLM’s benefit is most clear in scenarios where the questioner belongs to a different model class to the answerer: here Naive’s performance is much less convincing.
+Figure 2: Mean rating across 10 film recommendations: mean $\pm$ standard error across 200 users.
+7 Conclusion
+In this work, we have shown how to effectively apply the framework of sequential Bayesian experimental design (BED) to the problem of interactive information gathering with LLMs.
+Specifically, we have introduced BED-LLM, which provides a specific, information–theoretic, sequential BED approach that makes a variety of carefully justified design choices in the joint-model factorization, belief updating, and EIG estimation.
+Particularly central to BED-LLM is the prior–likelihood pairing with filtering of hypotheses for consistency with the history.
+BED-LLM is notably the first work that uses both this prior–likelihood pairing without making a deterministic likelihood assumption that causes the EIG to simply to just marginal predictive entropy.
+Together, these innovations lead to substantial performance improvements compared to previous approaches.
+The results thus confirm that principled EIG–driven strategies can yield substantial gains for interactive, multi-turn, information gathering problems.
+Appendix Contents
+Appendix A Algorithmic considerations
+A.1 Updating the likelihood
+The success of BED-LLM hinges on our ability to update our joint distribution.
+As mentioned in §3, we choose not to update the likelihood model as more data is gathered, that is, our likelihood in the sequential setting will be $p_{\mathrm{LLM}}\left(y_{t};\left[\theta,x_{t}\right]\right)$ instead of $p_{\mathrm{LLM}}(y_{t+1};[h_{t-1},\theta,x_{t+1}])$.
+The main rationale of this choice is that for many problems our beliefs on $\theta$ capture all the required information to predict $y|x$, hence including the history is adding unnecessary context that could influence the LLM’s behavior in undesirable ways.
+However, it is important to note that $p_{\mathrm{LLM}}(y_{t+1};[h_{t-1},\theta,x_{t+1}])$ should be used instead for problems where $\theta$ will not capture all information from previous data, e.g.
+if $\theta$ is a binary value corresponding to whether we reject a null hypothesis, or is the answer to a particular other question of interest.
+A.2 Estimating EIG for each question
 Question 1 Question:
@@
 This thus demonstrates how the two criteria can select different questions.
-A natural tractable alternative is to simply derive our beliefs through in-context updates, that is, use $p_{\mathrm{LLM}}(\theta;h_{t})$, noting that this has been shown to behave differently to Bayesian updating in the LLM context (Falck et al,, 2024; Kossen et al,, 2024).
-Unfortunately, we find that the ability of current LLMs to fully incorporate the information from $h_{t}$, even when using large state-of-the-art LLM variants such as GPT-4o (OpenAI,, 2024) or Mistral Large (Mistral AI Team,, 2024), is surprisingly lacking.
-In particular, we find that all LLMs we evaluate regularly sample hypotheses $\theta$ that are incompatible with previous observations in $h_{t}$, especially as the length of the history grows.
-We also, again, find problems with over–collapsing, with LLMs often jumping to conclusions about the value of $\theta$ without appropriate evidence for this certainty in $h_{t}$.
-The shortfalls of in-context learning in such settings have also previously be noted by, for example, (Liu et al, 2024b, ; Zhang et al, 2025b, ; Zhang et al,, 2024).
+A.2.1 Predictive entropy is not a good approximation for EIG
+As discussed in §3.2, previous information-based query selection mechanisms have assumed that responses are deterministic given $\theta$ and $x$.
+This implies that the expected entropy of the likelihood, $\mathbb{E}_{p(\theta;h_{t-1})}[\mathrm{H}[p(y_{t+1}|\theta;x_{t+1},h_{t-1})]]$, is constant over designs, meaning that maximizing EIG is equivalent to maximizing the marginal predictive entropy, $\mathrm{H}[\mathbb{E}_{p(\theta;h_{t-1})}[p(y_{t}|\theta;x_{t},h_{t-1})]]$.
+In practice, the expected likelihood entropy can and will vary across designs.
+This variability in the expected likelihood entropy can be crucial in selecting good designs.
+Here, we walk through a concrete example where predictive entropy might differ significantly from EIG.
+Figure˜3 shows two candidate questions that could be asked to elicit film preference.
+Question 1 has high predictive entropy: in a randomly selected group of people, we would expect high variation in ice cream preference (regardless of the individual’s film preferences).
+However, since ice cream preference is unrelated to film preference, the answer would not help us narrow down our hypothesis space, and the EIG is zero.
+This is also supported by evidence in our experiments (§6).
+Both the Split baseline, and the Entropy ablation, assume a deterministic likelihood; in particular, the Entropy ablation uses the same estimator of the predictive entropy as BED-LLM.
+In both cases, we see the performance significantly degrades relative to using the full EIG.
+Further, omitting the expected likelihood entropy term provides no meaningful computational saving—the same LLM evaluations are used for the top and bottom lines of Equation˜3, hence doing the full estimate of the EIG requires no additional LLM calls to be made.
+A.2.2 EIG estimator
+One might be tempted to replace $\hat{p}(y_{t+1};[h_{t-1},x_{t+1}])$ with $p_{\mathrm{LLM}}(y_{t+1};[h_{t-1},x_{t+1}])$ in the EIG estimation in Equation˜3, as the two essentially offer alternative predictive distributions for the outcome.
+We also advise against this though, noting that it again provides no meaningful computational benefits (unless one also assumes a deterministic likelihood, but this would then mean we no longer consider $\theta$ at all).
+A key reason for avoiding this substitution is that it would mean we are no longer estimating a true EIG: the inconsistency between the likelihood and the marginal data distribution means there is no longer a joint model where we are minimising our expected uncertainty in $\theta$.
+We also find that the LLM process of sampling $\theta$ from $p_{f}(\theta;h_{t-1})$ followed by $y$ from $p_{\mathrm{LLM}}\left(y_{t};\left[\theta,x_{t}\right]\right)$ tends to give a better uncertainty over responses than sampling $y$ directly from $p_{\mathrm{LLM}}(y_{t+1};[h_{t-1},x_{t+1}])$.
+A.3 Prior construction and belief updating
+In Section˜3.3, we argued that naive in-context updating is not sufficient for updating our beliefs: We fail to fully incorporate the information from the history $h_{t}$, and we often have overconfident distributions.
+The shortfalls of in-context learning in such settings have also previously be noted by, for example, (Liu et al., 2024b; Zhang et al., 2025; 2024).
 We posit two reasons why they likely struggle in such settings.
 First, the information from the different examples in the history are generally highly distinct in these information-gathering settings (indeed, this is part of our aim in adaptively design informative questions), making it harder for the LLM to appropriately reconcile all the provided information than in many other uses of in–context learning.
-Second, $\theta$ will often represent a user–specific variable that cannot easily be predicted from any data other than the user’s responses to questions: it has been argued that much of the success of in–context learning in LLMs is down to improving problem specification and linking the requested task to data it has seen in its training, rather than truly “learning” from the provided examples (Min et al,, 2022; Kossen et al,, 2024), but the history in our setting is rarely helpful for this due to its user–specific nature.
-Addressing these shortfalls is critical to the success of BED-LLM.
-Although we will still use $p_{\mathrm{LLM}}(\theta;h_{t})$ as the basis for deriving our belief state over $\theta$ (i.e.
-our intermediate prior), we make various alterations to effectively incorporate historical information and ensure diversity.
-We refer to our derived distribution as $p_{f}(\theta;h_{t})$ and its key component is to filter the generated hypotheses according to whether they are compatible with the history $h_{t}$.
-We do this by using the LLM to zero-shot check the compatibility of each sampled $\theta$ with all the previous question–answer pairs in $h_{t}$ (using $p_{\mathrm{LLM}}(y;[\theta,x])$) and then rejecting that sample if an incompatibility is found.
-Specifically, a sample is rejected if the likelihood of an observed answer falls below a predefined threshold, chosen to balance robustness to model uncertainty against the need to enforce strict historical coherence.
-To reduce the computational cost of generating and evaluating hypotheses, we further include a hypothesis-retention mechanism: any hypotheses from the previous step t1t\!-\!1 which remain consistent with the most recent question and observation are retained in the hypothesis set for step $t$ without regeneration.
-We also make some further variations from $p_{\mathrm{LLM}}(\theta;h_{t})$ to try and combat the over–collapsing issue and encourage diversity (noting that our initial sampling acts like a proposal in a rejection sampler and invalid samples should still be removed by the filtering).
-First, we sample at a high temperature ($T=1.3$).
-Second, rather than generating candidates independently, we prompt the LLM to generate batches of candidates using a prompt that requests a “diverse and representative set of candidates”.
-We can sample multiple batches, with the previous batches in context to force the LLM to generate different samples.
-Finally, we assume a uniform distribution over the $\theta$ that make it through the rejection, such that we remove any duplicated samples and do not make any direct use of the LLM’s logits on $\theta$.
-We note that the precise prompting strategy used in $p_{\mathrm{LLM}}(\theta;h_{t})$ can also be quite important in achieving effective performance, with our prompts set up to try and maximize the diversity in the generated $\theta$ samples.
-Details on our exact setup for $p_{f}(\theta;h_{t})$ are in Appendix˜B.
-Under the aforementioned setup for $p_{f}(\theta;h_{t})$, the original prior $p(\theta)$ only affects the EIG at the first step as it has no direct bearing on $p_{f}(\theta;h_{t})$ for $t>0$.
-We, therefore, define our original prior to simply be equal to $p_{f}(\theta;h_{t})$ with $h_{t}=\emptyset$, such that no filtering is performed.
-We note though that in some cases it may be preferable to specify some customized $p(\theta)$, potentially then adjusting the filtering process and/or weighting the accepted samples in a non–uniform manner to ensure this prior information is not lost as more data is gathered.
-4 On the Specification of $p(\theta,y_{t+1};h_{t},x_{t+1})$, and its Implications
-As we described in Section˜3, adapting BED to the LLM setting hinges upon how we specify, and update, the joint distribution $p(\theta,y_{t+1};h_{t},x_{t+1})$.
-As we discuss in this section, this choice is critical to the success of BED-LLM and impacts how we model the current uncertainty about $\theta$; how easily and accurately we can estimate EIG; and how aligned the model’s actual information gain under a question is predicted by the EIG, which ultimately governs whether the questions we choose are actually good at reducing uncertainty about $\theta$.
-As discussed at the start of Section˜3, there are two distinct ways to derive the joint model from our LLM: using a prior–likelihood pairing, $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}p(\theta;h_{t})p\left(y_{t+1};\left[\theta,x_{t+1}\right]\right)$}, or a data–estimation pairing, $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}p\left(y_{t+1};[h_{t},x_{t+1}]\right)p\left(\theta;[h_{t},x_{t+1},y_{t+1}]\right)$}, [^4] One could in principle also consider prompting the LLM to generate $\theta$ and $y$ jointly as well, but we will omit these from our discussions on the basis that this will generally be impractical from an estimation perspective, while LLMs’ autoregressive nature means that this should be equivalent to one of the aforementioned joints, up to effects in prompting strategy and API handling etc.
-where each of the $p$ terms either corresponds directly to the LLM’s distribution over outputs given the provided context and suitable prompting (i.e.
-the corresponding $p_{\mathrm{LLM}}(\cdot;\cdot)$), some valid distribution derived from this, or an externally specified distribution.
-The first construction mirrors deriving our beliefs about $\theta$ from a conventional Bayesian posterior with a concrete prior and likelihood derived (at least partially) from the LLM, whereas the second is analogous to a one-sample martingale posterior that directly samples hypothetical data and draws inferences on $\theta$ given hypothetical data using in-context learning (Fong et al,, 2023; Falck et al,, 2024).
-In our outlined BED-LLM approach (c.f.
-Section˜3), we adopted a prior–likelihood pairing.
-Below, we will expand on some reasons to prefer this formulation and discuss certain settings where the data–estimation might be preferable.
-4.1 Modeling flexibility
-The most obvious relative merits of the prior–likelihood and data–estimation pairings are in the flexibility in how each term is chosen.
-The prior–likelihood pairing gives us greater flexibility to construct a prior set of beliefs over $\theta$ that is distinct to the LLM’s internal beliefs, as it allows us to directly control this prior by changing $p(\theta;h_{t})$, whereas the prior is only implicitly defined in the data–estimation pairing.
-In Section˜3.3 we exploited this flexibility through our definition of $p_{f}(\theta;h_{t})$, wherein we made various adjustments to $p_{\mathrm{LLM}}(\theta;h_{t})$ to encourage diversity and better incorporate information from the history.
-On the other hand, the data–estimation pairing could provide some beneficial flexibility in specifying how the data itself is simulated through changing $p\left(y_{t+1};[h_{t},x_{t+1}]\right)$.
-Most notably, in some applications like scientific discovery, we may have access to external simulators that can be run to predict our data, with the LLM used instead primarily to automate how we draw conclusions from such hypothetical data (Ghareeb et al,, 2025).
-In the cases considered in this paper, we do not have such external, validated simulators, and the flexibility in prior construction is more important for our needs.
-4.2 Faithfulness of conditional distributions
-In Section˜4.1, we considered flexibility in specifying the marginal distributions $p(\theta;h_{t})$ and $p\left(y_{t+1};[h_{t},x_{t+1}]\right)$.
-While deviations from relying on direct LLM predictions are also in principle possible for the conditional models $p\left(y_{t+1};\left[\theta,x_{t+1}\right]\right)$ and $p\left(\theta;[h_{t},x_{t+1},y_{t+1}]\right)$, in practice, these will typically be more difficult and expensive to implement.
-This is first because these conditionals need to be instantiated for each sampled instance of the conditional variable ($\theta$ and $y_{t+1}$ respectively), rather than just needing us to set up one marginal distribution that can then be repeatedly sampled from.
-Second, to construct estimators for Equations (2) and (3), we require access to concrete probabilities for the conditional distributions (in order to calculate entropies), whereas we only needed to draw samples for the marginal distributions (in order to approximate expectations).
-As such, the conditionals need to be explicit distributions, or at least distributions where the probability can be cheaply estimated (noting that such estimation will be nested and thus more expensive than in conventional settings (Rainforth et al,, 2018)), so they are more difficult to define through the output of some algorithmic procedure, especially if the number of possible realizations is large.
-When considering the conditional distributions, the decisive question on the relative merit of the two formulations is which conditional factor we are willing to trust the LLM to supply as a full probability distribution.
-Critically, we rely on how the LLM captures uncertainty in this full distribution—including, for example, tail behavior—not merely the fidelity of typical samples; the marginal factors, by contrast, only need to be sampled from.
-If we accept the LLM’s direct predictive distribution for $p\left(y_{t+1};\left[\theta,x_{t+1}\right]\right)$, then we are basing our notion of uncertainty around (and will need to calculate) $\mathrm{H}[p_{\mathrm{LLM}}\left(y_{t+1};\left[\theta,x_{t+1}\right]\right)]$, and if instead we place more faith in the LLM’s internal distribution for $p\left(\theta;[h_{t},x_{t+1},y_{t+1}]\right)$, then we are basing our uncertainty around $\mathrm{H}[p_{\mathrm{LLM}}(\theta;[h_{t},x_{t+1},y_{t+1}])]$.
-In essence, the choice between prior–likelihood and data–estimation pairings thus comes down to whether we believe the LLM will produce a more appropriate conditional uncertainty over $\theta$ or $y$, along with our ability to numerically estimate this uncertainty cheaply (noting the nesting means we need to repeatedly carry out this estimation for different $\theta$ and $y$ samples respectively).
-This difference becomes particularly noticeable when the complexities of the spaces of $\theta$ and $y$ differ significantly.
-Our ability to draw sensible samples of either will generally be quite robust to these spaces being complex or high–dimensional; this is where LLMs tend to thrive, effectively generating highly complex outputs in an autoregressive manner.
-However, evaluating the entropy of a distribution becomes dramatically harder as the dimensionality or complexity increases (Acharya et al,, 2019; Paninski,, 2003), and the entropy of the predictive distribution of an LLM in such cases will not typically provide a sensible measure of uncertainty (Kadavath et al,, 2022; Desai & Durrett,, 2020).
-As such, the decision between joint formulations should predominantly be based on the complexity of the space of $\theta$ versus that of $y$: we should generally favor the prior–likelihood formulation if $\theta$ is more complex and the data–estimation formulation if $y$ is more complex.
-For example, suppose $\theta$ represents the space of Python code a user wants to generate and $y$ is the answer to a simple yes/no question about that code’s desired behavior.
-Here, the LLM’s entropy over yes vs.
-no when given the code will be a much more meaningful uncertainty measure than its entropy over code when told the answer to the question.
-For the problems that we consider, the space of $y$ is less complex than that of $\theta$, indicating we should, in general, use the prior–likelihood formulation.
-However, in cases where this is not true, the data–estimation formulation may be preferable instead.
-We provide additional discussion on the impact of the chosen pairing on our entropy estimate, plus discussion on the choice of $\theta$, in Appendix˜A.
-4.3 Extracting the belief state
-A further advantage of the prior–likelihood construction is that our belief state on $\theta$ can be extracted directly, as it is simply $p_{f}(\theta;h_{t})$.
-With the data–estimation construction, we would have to estimate a marginal distribution from $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}p\left(y_{t+1};[h_{t},x_{t+1}]\right)p\left(\theta;[h_{t},x_{t+1},y_{t+1}]\right)$} by integrating out the synthetic response $p(\theta;h_{t})=\int{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}p\left(y_{t+1};[h_{t},x_{t+1}]\right)p\left(\theta;[h_{t},x_{t+1},y_{t+1}]\right)}\;\mathrm{d}y_{t+1}$.
-Direct access to $p(\theta;h_{t})$ is also important to ensure that our current belief state is independent of the next question $x_{t+1}$, as the marginal obtained from the integral above is unlikely to remain the same for all $x_{t+1}$ in the way we intuitively want it to.
-4.4 Alignment between EIG and belief updating procedure
+Second, $\theta$ will often represent a user–specific variable that cannot easily be predicted from any data other than the user’s responses to questions: it has been argued that much of the success of in–context learning in LLMs is down to improving problem specification and linking the requested task to data it has seen in its training, rather than truly “learning” from the provided examples (Min et al., 2022; Kossen et al., 2024), but the history in our setting is rarely helpful for this due to its user–specific nature.
+Appendix B Additional discussions on design choices
+B.1 An alternative view on the faithfulness of conditional distributions
+Another way of viewing the distinction between the prior–likelihood and data–estimation constructions is in which of the EIG forms, Equation˜1 or Equation˜2, we center our reasoning.
+For a given joint model, the two are, of course, mathematically equivalent.
+However, they give us different ways of thinking about what it means to maximize the EIG: reducing entropy in $\theta$ from seeing $y$, or reducing entropy in $y$ from seeing $\theta$.
+This, in turn, gives us a way to reason about how appropriate our joint model is.
+When we choose to use one of $p\left(y_{t};\left[\theta,x_{t}\right]\right)$ or $p\left(\theta;[h_{t-1},x_{t},y_{t}]\right)$, we are centering our reasoning around the entropy of this quantity making sense, while allowing the other entropy in the other form to be implicitly defined from the resulting joint distribution; because the two forms are equivalent, we know that if our explicit form is suitable/unsuitable, the implicit form will be as well.
+If, for example, we directly fix the form of $p\left(\theta;[h_{t-1},x_{t},y_{t}]\right)$ using our LLM’s predictive distribution, we are also directly relying on its expected entropy being a meaningful measure of design quality.
+If $\theta$ is high–dimensional and predominantly free–form, the resulting entropy produced by the LLM is unlikely to be meaningful and using the data–estimation pairing is unlikely to produce an effective strategy.
+However, if $y$ is instead quite constrained, the LLM can produce a meaningful entropy over it, and choosing a model based on the prior–likelihood pairing is likely to implicitly define a meaningful distribution, and thus entropy, on $\theta$.
+Conversely, if $\theta$ is constrained and $y$ is free form, the opposite will hold instead.
+B.2 Choice of $\theta$
+An important corollary of this reasoning is that it can be important to be careful in our choice of exactly what we take $\theta$ to be, especially if we are using the data–estimation formulation.
+In particular, it is essential for entropy in the space of $\theta$ to form a meaningful notion of uncertainty, even if this entropy is not being measured through the LLM’s predictive distribution of $\theta$ directly.
+Thus, while $\theta$ inherently represents what we are trying to learn about and should always be set up as such, if there is flexibility in how exactly we formulate it, we should be careful to choose a form that yields an appropriate uncertainty measure.
+For example, if the LLM is trying to clarify what code a user wishes it to generate, we could either choose $\theta$ to be the code itself or, following (Neiswanger et al., 2021; Bickford Smith et al., 2023), the output the code produces.
+Here the entropy over code outputs induced by our distribution on code is likely to be a much better measure of uncertainty than the entropy of the raw code itself, given that there are multiple ways one can code the same operation.
+B.3 Alignment between EIG and belief updating procedure
 Our ultimate goal is to minimize uncertainty in $\theta$, as measured by its entropy.
-With this in mind, we can use the expected uncertainty reduction framework of Bickford Smith et al, (2025) to provide insights into how well our EIG formulation and belief updating procedures align.
+With this in mind, we can use the expected uncertainty reduction framework of Bickford Smith et al.
+(2025) to provide insights into how well our EIG formulation and belief updating procedures align.
 To simplify discussions, for now we consider the setting where we choose a single question $x$ and obtain a response $y$.
-Following Bickford Smith et al, (2025), we can think of the “true” optimal design as selecting
+Following Bickford Smith et al.
+(2025), we can think of the “true” optimal design as selecting
 $x^{*}_{\mathrm{true}}=\mathrm{argmin}_{x}\mathbb{E}_{p_{\mathrm{true}}(y;x)}\left[\mathrm{H}[p(\theta;x,y)]\right],$ (4)
@@
 Thus, we can now see that our choice of joint model corresponds to different choices for approximating these quantities.
-Assuming that the LLM distribution is used directly for the conditional as per Section˜4.2, we thus have that:
+Assuming that the LLM distribution is used directly for the conditional as per Section˜4, we thus have that:
   • The prior-likelihood pairing equates to the approximations $p_{\mathrm{true}}(y;x)\approx\int p(\theta)p_{\mathrm{LLM}}(y;[\theta,x])d\theta$ and $p(\theta;x,y)\approx p(\theta)p_{\mathrm{LLM}}(y;[\theta,x])/\int p(\theta)p_{\mathrm{LLM}}(y;[\theta,x])d\theta$;
@@
 The appropriateness of each of these options, therefore, comes down to how faithful these approximations are respectively to the true data distribution, $p_{\mathrm{true}}(y;x)$, and how we actually derive our belief distribution on $\theta$ in practice once we have seen the new data.
-The former of these considerations is difficult to control for as we simply do not know the true response distribution and it is hard to say which approach will thus estimate it best (though we can refer to the discussion in Section˜4.1 to determine which best matches our beliefs about the true response distribution).
+The former of these considerations is difficult to control for as we simply do not know the true response distribution and it is hard to say which approach will thus estimate it best (though we can refer to the discussion in Section˜4 to determine which best matches our beliefs about the true response distribution).
 However, we do know upfront how we plan to derive our belief distribution on $\theta$ in practice, so we can use this to guide which joint model we formulate our EIG from.
-Namely, we see that using
-  • The prior–likelihood EIG pairing equates to assuming we will make a Bayesian update to our beliefs on $\theta$ using the likelihood $p_{\mathrm{LLM}}(y;[\theta,x])$;
-  • The data–estimation EIG pairing equates to assuming we will make an in-context update to our beliefs on $\theta$, as we are treating $p(\theta;x,y)$ as $p_{\mathrm{LLM}}(\theta;[x,y])$.
+Namely, we observe that: a) using the prior–likelihood EIG pairing equates to assuming we will make a Bayesian update to our beliefs on $\theta$ using the likelihood $p_{\mathrm{LLM}}(y;[\theta,x])$; b) using the data–estimation EIG pairing equates to assuming we will make an in-context update to our beliefs on $\theta$, as we are treating $p(\theta;x,y)$ as $p_{\mathrm{LLM}}(\theta;[x,y])$.
 Our preference between the pairings should therefore be guided in part by how we plan to update the model in practice.
@@
 For the settings we consider, we expect $p_{f}(\theta;[x,y])$ to generally be better approximated by a Bayesian update than an in–context update, aligning with our decision to use the prior–likelihood formulation.
-The reasons for this are that a) the filtering often removes a large proportion of the generated samples, especially at later experiment turns, with $p_{\mathrm{LLM}}(\theta;h_{t})$ not fully incorporating information from the history; b) the maintaining of the set of one consistent hypotheses from one turn to the next encourages a more Bayesian behavior, with samples persisting unless contradicted by a new likelihood term; and c) the typical premature over–collapse of $p_{\mathrm{LLM}}(\theta;h_{t})$ to a small number of hypotheses means it is typically unrepresentative of our beliefs.
+The reasons for this are that a) the filtering often removes a large proportion of the generated samples, especially at later experiment turns, with $p_{\mathrm{LLM}}(\theta;h_{t-1})$ not fully incorporating information from the history; b) the maintaining of the set of one consistent hypotheses from one turn to the next encourages a more Bayesian behavior, with samples persisting unless contradicted by a new likelihood term; and c) the typical premature overconfidence of $p_{\mathrm{LLM}}(\theta;h_{t-1})$ to a small number of hypotheses means it is typically unrepresentative of our beliefs.
 These theoretical benefits are perhaps secondary to the more practical benefits from the ease of constructing an appropriate model in the prior–likelihood formulation and avoiding direct uncertainty estimation in the space of $\theta$.
 Nonetheless, they help confirm that our choices have not induced unnecessary mismatch between the EIG formulation and our updating procedure.
-The analysis here can get a little more complicated when we consider the fact that “one-step optimal” decisions like Equation˜6 may no longer be optimal in the sequential decision–making setting.
-We provide more discussion on this in Section˜A.3.
-5 Related Work
-In this work, we consider the broad problem of sequential information gathering using LLMs—that is, adaptively asking natural-language questions to allow the LLM to rapidly learn about a parameter of interest $\theta$.
-Several works have explored the baseline ability of LLMs to ask such questions (Zhang et al,, 2024; Li et al, 2025b, )—effectively, the “naive" baseline in our experiments.
-While these works demonstrate some ability to adaptively construct information-seeking questions, they often fail to extract important information (Li et al, 2025a, ), leading us to more sophisticated approaches.
-Ours is also not the first paper to use a targeted model-based approach to information gathering with LLMs: several existing works have previously attempted to choose questions based on estimates of the expected information gain, marginal predictive entropy, or similar quantities (Piriyakulkij et al,, 2023; Hu et al,, 2024; Kobalczyk et al,, 2025; Wang et al,, 2025; Mazzaccara et al,, 2024; Chan et al,, 2025; Cooper et al,, 2025).
-However, as explained below, these existing methods all make design choices that significantly limit performance.
-Further, none of them provide comparable insights into how the joint model should be best formulated and updated from the underlying LLM; most of them even fail to directly link their approach and BED itself.
-A number of these previous works, including all that also employ a prior–likelihood pairing in their joint model, actually implicitly select queries using predictive entropy in practice (rather than EIG as claimed) because they assume a constant expected likelihood entropy (Kobalczyk et al,, 2025; Mazzaccara et al,, 2024; Cooper et al,, 2025).
-As we discussed in Section˜3.2.1 and show empirically in Section˜6, this simplification is not generally reasonable—as predictive entropy is not in general a good approximation to EIG—and leads to worse performance.
-We believe our work to be the first which uses a prior–likelihood model pairing without simplifying the EIG to just the predictive entropy.
-All of the other methods have instead relied on models derived from a data–estimation pairing, then either obtaining a sample-based estimate of $\mathbb{E}_{p\left(y_{t+1};[h_{t},x_{t+1}]\right)}\left[\mathrm{H}[p_{\mathrm{LLM}}\left(\theta;[h_{t},x_{t+1},y_{t+1}]\right)]\right]$ (Piriyakulkij et al,, 2023; Hu et al,, 2024; Wang et al,, 2025), or in the case of Chan et al, (2025), an approximation to this entropy based on the size of conformal sets.
-As discussed in Section˜4, our prior–likelihood model pairing provides significant benefits, particularly in the large or unbounded hypothesis spaces we explore in this paper.
-To get an approach which is based on the data–estimation model pairing to work, previous works have thus had to restrict the set of allowable hypotheses (i.e.
-$\theta$) to a small set, typically also restricting the space of allowing outcomes and further making a deterministic likelihood assumption (thereby again implicitly reducing the EIG to predictive entropy).
-These restrictions significantly limit applicability and performance.
-Rather than augmenting a frozen LLM with the ability to estimate utility functions, some works have instead aimed to post-train an LLM to improve its ability to ask questions (Zhang et al,, 2024; Wu et al,, 2025; Andukuri et al,, 2024).
-Most do not explicitly consider informativeness of questions: Zhang et al, (2024) and Wu et al, (2025) use reinforcement learning techniques to reward generations that quickly lead to the correct answer, and Andukuri et al, (2024) builds on Li et al, 2025b by fine-tuning on successful traces.
-Mazzaccara et al, (2024) do indirectly incorporate uncertainty: they use predictive entropy to identify informative questions, and then either fine-tune on the highest-entropy question, or perform DPO comparing the highest-entropy question with a lower-entropy question.
-We do not address fine-tuning in this work, focusing instead on exploring the correct way to formulate BED using LLMs.
-As discussed in Section˜4, a key challenge in adapting BED to the LLM setting is in aligning the expected information gain with the actual uncertainties extracted from the LLM after updating.
-Handa et al, (2024) take a different approach to this problem by using the LLM to generate features for an external conventional Bayesian joint model (in their case, a linear Bradley Terry model), rather than deriving their joint model more directly from the LLM itself.
-This can be a good choice when we already have a well-specified Bayesian model form for the problem at hand; however, this may be challenging in arbitrarily large and complex hypothesis spaces.
-In particular, their specific method is not applicable more widely beyond the preference learning context they consider.
-6 Experiments
-We now empirically evaluate our proposed BED-LLM method, compare it against relevant baselines, and analyze its properties through targeted ablation studies.
-We consider two interactive question–answering scenarios that test different types of sequential information acquisition: 1) 20–Questions: a deductive reasoning game requiring identification of a hidden target entity through binary (Yes/No) queries, and 2) Active Preference Elicitation: a task in which the goal is to infer a latent user preference vector via multiple–choice questions.
-To evaluate the different questioner methods (i.e., BED-LLM and our baselines), we further need some form of answerer that represents the user or external environment we are trying to learn about.
-Due to the complications of evaluating with real human participants, we instead conduct our tests with a simulated user based on a separate LLM instance that has been given a ground truth target $\theta^{*}$ (corresponding to the hidden target entity for 20–Questions and a user profile for the active preference elicitation).
-Note that this answerer LLM does not itself perform any learning—it does not keep previous question–response pairs in context.
-While we consider both deriving the questioner and answerer from the same LLM and from different LLMs, we emphasize that they are always using separate LLM instances with different prompts and contexts, thus no information is transferred between them except in the questions and responses themselves.
-We compare BED-LLM against two baselines:
-a) Entropy: selecting questions using the maximum predictive entropy, as done in Kobalczyk et al, (2025), Mazzaccara et al, (2024), and Cooper et al, (2025).
-For fairness and to ensure a strong baseline, we keep the entropy setup as close to BED-LLM as possible (including prompts, candidate question generation, and our joint model construction with hypothesis filtering), varying only in replacing the EIG estimator with an estimator of the marginal predictive entropy,
-$\mathrm{Entropy}_{\theta}(x_{t+1})=-\sum_{y_{t+1}\in\mathcal{Y}}\hat{p}(y_{t+1};[h_{t},x_{t+1}])\log\hat{p}(y_{t+1};[h_{t},x_{t+1}]).$ (7)
-We construct an estimator for Equation˜7 analogously to Equation˜5.
-We also note here that our version of the Entropy baseline produces noticeably stronger performance on the 20–Questions problem than Kobalczyk et al, (2025) report for their own “EIG” approach, demonstrating it efficacy.
-[^5] Though Kobalczyk et al, (2025) describe their approach as targeting the EIG, the estimator they use (given by their Algorithm 1) actually equates to a sample-based estimator of the marginal predictive entropy instead and is thus comparable with our Entropy baseline.
-As they do not provide code specifically for the 20–questions experiments or explicit results on their approaches’ success rate, we were unable to provide direct comparisons against their exact setup.
-However, we believe that it is still clear from the results they do present that their approach is performing worse than the Entropy baseline we do consider.
-First, the reported ranks they give indicate they are not beating the naive LLM approach with the same regularity as Entropy does.
-Second, in their rebuttal on OpenReview they explain that they only report ranks and not success rate because “only very few runs would result in a non-zero accuracy”, whereas we see that our Entropy baseline is regularly successful.
-b) Naive QA:we prompt the questioner LLM to directly produce an informative next question without explicit hypothesis generation, EIG computation, or filtering.
-The question is obtained by drawing a single sample at temperature $T=1$ from the questioner LLM.
-This approach, which was explored in Zhang et al, (2024), relies on the implicit reasoning capabilities of the LLM to select good questions.
-6.1 20–Questions
-We first evaluate BED-LLM in the classic “20–Questions” game, where the questioner must identify a hidden target $\theta^{*}$ by asking up to 20 Yes/No questions.
-Here, the hypothesis space of possible $\theta$ is large and not explicitly defined or restricted: it is bounded only by what the LLM can generate.
-As such, all previous work that has considered using BED-style decision with LLMs either a) cannot be directly applied in our precise problem setup because it requires the questioner to be given access to an explicit, restricted set of possible $\theta$ (Hu et al,, 2024; Piriyakulkij et al,, 2023; Wang et al,, 2025; Chan et al,, 2025), or b) is a variant of our Entropy baseline, due to using marginal predictive entropy in lieu of the EIG (Kobalczyk et al,, 2025; Mazzaccara et al,, 2024; Cooper et al,, 2025).
-We evaluate across three distinct datasets each containing a mix of 100 obscure and common targets corresponding to either Animals, Celebrities, or Things.
-Here, the dataset is just a list of different $\theta^{*}$ that will be individually provided to the answerer to instantiate different problems (e.g.
-we conduct a trial where $\theta^{*}=$“dog”, then one where $\theta^{*}=$“cat”, etc).
-The list of targets is never provided to the questioner model to restrict the set of possible hypotheses: the questioner is only prompted that is trying to identify an “animal”, “celebrity”, or “thing”.
-The datasets are: a) Animals: a set of animal species generated with OpenAI’s o3 model to ensure a diverse mix and balanced taxonomy, see Appendix˜D for details; b) Celebrities: a diverse set of public figures, as used by Zhang et al, (2024); c) Things: a collection of everyday and exotic entities drawn from the web corpus, as used by Zhang et al, (2024); it covers a wide range of categories, from plants and clothing to professions, events, and mythical creatures.
-We note that the best success rates achieved by Zhang et al, (2024) on Celebrities and Things are substantially below those of both BED-LLM and our Entropy baseline.
-Using our sample–then–filter process (c.f.
-Section˜3.3), we aim to sample at least $N=15$ hypotheses, repeating the cycle up to three times if needed (the exact number of hypotheses can be less than this as it may not be possible to generate sufficient valid hypotheses, especially in later experiment turns).
-The questioner generates $M=15$ candidate questions to test, $\mathcal{X}^{\text{cand}}$, using the “conditional generation” approach of Section˜3.1 when possible, but falling back on “unconditional generation” if insufficient candidate hypotheses have been generated.
-Evaluation
-[…]
-Table 1: Success rate (%) for 20–Questions at the end of the game.
-Best result in bold.
-Success rates are integers because each dataset contains exactly 100 entities.
-$\pm$ numbers represent the standard error of the mean estimate (this is estimated using $\sqrt{p(1-p)/(n-1)}$ where $p$ is the success percentage and $n$ is the number of datapoints, noting this estimator is positively biased and thus conservative).
-Success Rate (%) $\uparrow$
-Dataset Model Naive QA Entropy BED-LLM
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-Figure 2: Success rate against turn for the 20–Questions game.
-Error bars represent the standard error of the mean estimate across 100 target entities in each dataset.
-The same model used as both questioner and answerer.
-Figure 3: Success rates for the 20–Questions game with different models: the questioner Q and answerer A.
-Results
-First, we evaluate in the setting where the same LLM is used as both the questioner and the answerer.
-Table˜1 shows the success rates for the different datasets and a variety of different LLMs.
-Across all datasets and all LLMs, BED-LLM achieves the highest success rate, significantly outperforming the Entropy and Naive QA baselines.
-It is particularly notable how the success rate of our principled information-gain strategy is typically more than double that of Naive QA which relies solely on implicit LLM reasoning.
-Per-turn success rates are shown in Figure˜2.
-Over the first few turns, BED-LLM and Entropy perform similarly.
-However, after around 7-10 turns we consistently see BED-LLM pull away from Entropy.
-This may simply be because it takes some time for the effect of choosing good questions to take effect (the large space of possible $\theta$ means that no approach can reliably determine the answer in only a few steps), or it may be that the effect of using proper EIG estimates becomes more pronounced as the set of viable hypotheses decreases and it is harder to find useful questions.
-Different questioner and answerer
-In the experiments above, the same LLM served as both the questioner and the answerer.
-This setting implicitly assumes that the questioner model has access to the answerer’s response distribution, i.e.
-that there is no misspecification in the likelihood used by BED-LLM and Entropy or the internal reasoning of Naive QA.
-However, in practice, the answerer is typically a human whose responses will follow a different distribution to the questioner’s internal model for reasoning about responses.
-To test the robustness of the various methods to this model misspecification, we therefore also evaluate scenarios where the questioner and answerer are formulated with entirely different LLMs, instead of just different instances of the same LLM.
-As shown in Figure˜3, the benefit of BED-LLM persists even under this model misspecification; it consistently achieves higher success rates than both baselines across datasets.
-In contrast, Entropy exhibits a more severe drop in performance in this more realistic setting, often now performing more in line with Naive QA, which itself appears to be quite robust to the misspecification.
-Figure 4: Data–estimation ablation using Qwen2.5-72B-Instruct on Celebrities dataset.
-Data–estimation ablation
-As described in Section˜3, an important design choice is whether we choose the prior–likelihood pairing or the data–estimation pairing for BED-LLM.
-As an ablation for this critical decision, we also implement a variant of the alternative Data–Estimation method (see Appendix˜C for details).
-This gives an approach that is similar to those of Hu et al, (2024); Piriyakulkij et al, (2023); Wang et al, (2025); Chan et al, (2025), but which does not require us to assume a single fixed set of possible hypotheses throughout the sequential BED process.
-We evaluate this alternative approach on the Celebrities dataset using Qwen2.5-72B as both the questioner and answerer.
-Results are shown in Figure 4.
-Consistent with our analysis in Section˜4, BED-LLM substantially outperforms the Data–Estimation alternative.
-Data–Estimation still outperforms the Naive QA baseline, indicating that even this pairing benefits from explicit EIG–driven question selection over purely heuristic question generation, but interestingly does worse than Entropy, emphasizing the importance of estimating our uncertainty in the $y$ space instead of $\theta$ space.
-These findings reinforce our claim that the choice of joint–model factorization is a critical design decision for sequential BED in the LLM setting.
-6.2 Active Preference Elicitation
-In our 20–Questions problems, the generated hypotheses were all single, concrete entities (barring any failures of the LLMs) and most reasonable questions had clear–cut answers (e.g.
-“is Beyoncé a singer?”).
-In contrast, many problems, especially those related to eliciting information from a user, involve more abstract targets and/or more subjective and less predictable answers.
-For example, if we are learning about the characteristics or preferences of a user, then it may be difficult to explicitly define a concrete closed set of possible $\theta$, while it is also quite challenging for the LLM to accurately predict user responses and develop appropriate uncertainty estimates.
-To study such a scenario, we evaluate BED-LLM on inferring users’ film preferences.
-Here the target we are trying to gather information about is somewhat abstract, and we have some flexibility in how we define $\theta$ in our joint model.
-Our chosen setup is to define $\theta$ to be a natural language paragraph describing a user’s taste in films; we refer to this as a persona.
-To emulate a human user, our answerer is set up to be an LLM prompted to answer questions as if it were a person with a given persona $\theta^{*}$.
-To generate a dataset of ground-truth personas, we take a set of $200$ real user ratings from the MovieLens-100K dataset (Harper & Konstan,, 2015), then use an “oracle” LLM (namely, OpenAI’s o3 model) to produce a natural language persona that is consistent with each distinct set of ratings.
-As was the case for the 20–Questions problems, this dataset is never provided to the questioner and the set of allowed $\theta$ is not constrained.
-Because we need the LLM to be able to meaningfully capture uncertainty in the space of responses as discussed in Section˜4.2, we restrict questions to be multiple-choice.
-Specifically, the questioner is tasked with producing a question along with five possible responses A/B/C/D/E.
-We then define each $x_{t}$ to be the question coupled with the possible responses, and each $y_{t}$ to be one of the letters A/B/C/D/E to provide a restricted set of tokens over which we can measure entropy.
-Option E is further constrained to always be “none of the above” so that the answerer is not committed to choosing one of the directly generated choices if none are suitable.
-For BED-LLM and Entropy, we compare $M=8$ candidate questions at each turn and we aim to generate at least $N=5$ candidate hypotheses.
-We use the “unconstrained generation” approach of candidate question generation (c.f.
-Section˜3.1) as the user profiles can be quite diffuse and we are only generating a small number of possible hypotheses that can be quite easy to split.
-We note that data–estimation setup is not at all viable for this problem because the large number of tokens and varying dimensionality of each $\theta$ sample mean that $\mathrm{H}[p_{\mathrm{LLM}}(\theta;[h_{t},x_{t+1},y_{t+1}])]$ is not only infeasible to estimate, but also is not a meaningful measure of uncertainty.
-(a) Qwen2.5-72B
-(b) GPT-4o-mini
-(c) Llama-3.3-70B
-(d) GPT-4o
-Figure 5: Results for the active preference elicitation with the same LLM used for the questioner and answerer.
-Shown is the average ratings of 10 recommendations for each of 200 separate simulated users.
-Evaluation
-To again allow tracking of progress through the experiment, after each turn of the interaction $t$, the questioner generates ten film recommendations, conditioned on $h_{t}$.
-These recommendations are checked for consistency with prior questions and answers; if any are judged inconsistent then they are removed and additional recommendations are generated.
-The quality of the film recommendations is then assessed using an “LLM-as-judge” protocol (Zhu et al,, 2025; Trivedi et al,, 2024).
-Namely, the answerer evaluates each of the 10 films recommended by the questioner, conditioned on the hidden target user profile $\theta^{*}$.
-It scores each film on a scale of 1 to 5 (in 0.5 increments), based on how well the recommendation aligns with $\theta^{*}$ — this score is output together with a brief justification to increase reliability.
-We report the mean rating and standard error across 200 users, over 5 question–answer turns.
-Results
-In Figure˜5, we show the results using the same model for both questioner and answerer.
-For all models, BED-LLM establishes an advantage over other baselines by turn $t=3$, although the differences in performance between methods are smaller than for the 20–Questions game.
-On average, Entropy performs similarly to Naive QA.
-(a) Q: Qwen2.5-72B. A: GPT-4o-mini.
-(b) Q: GPT-4o-mini. A: Qwen2.5-72B.
-Figure 6: Results for Active Preference Elicitation with a different models: questioner Q and answerer A.
-Different questioner and answerer
-Like in the 20–Questions setting, we now also evaluate performance on Active Preference Elicitation in the misspecified model scenario by using a different LLM for the questioner and answerer.
-Figure˜6 presents the results.
-While BED-LLM and Entropy seem to be quite robust to the model misspecification, it is now Naive QA that seems to be brittle, exhibiting a drop in performance when using Qwen2.5-72B questioner and GPT-4o-mini answerer.
-7 Conclusion
-In this work, we have shown how to effectively apply the framework of sequential Bayesian experimental design (BED) to the problem of interactive information gathering with LLMs.
-Specifically, we have introduced BED-LLM, which provides a specific, information–theoretic, sequential BED approach that makes a variety of carefully justified design choices in the joint-model factorization, belief updating, and candidate question generation.
-Particularly central to BED-LLM is the prior–likelihood pairing with filtering of hypotheses for consistency with the history.
-BED-LLM is notably the first work that uses both this prior–likelihood pairing without resorting to using marginal predictive entropy in lieu of the expected information gain (EIG).
-Together, these innovations lead to substantial performance improvements compared to previous approaches.
-The results thus confirm that principled EIG–driven strategies can yield substantial gains for interactive, multi–turn, information gathering problems.
-Appendix A Additional discussions on design choices
-A.1 An alternative view on the faithfulness of conditional distributions
-Another way of viewing the distinction between the prior–likelihood and data–estimation constructions is in which of the EIG forms, Equation˜2 or Equation˜3, we center our reasoning.
-For a given joint model, the two are, of course, mathematically equivalent.
-However, they give us different ways of thinking about what it means to maximize the EIG: reducing entropy in $\theta$ from seeing $y$, or reducing entropy in $y$ from seeing $\theta$.
-This, in turn, gives us a way to reason about how appropriate our joint model is.
-When we choose to use one of $p\left(y_{t+1};\left[\theta,x_{t+1}\right]\right)$ or $p\left(\theta;[h_{t},x_{t+1},y_{t+1}]\right)$, we are centering our reasoning around the entropy of this quantity making sense, while allowing the other entropy in the other form to be implicitly defined from the resulting joint distribution; because the two forms are equivalent, we know that if our explicit form is suitable/unsuitable, the implicit form will be as well.
-If, for example, we directly fix the form of $p\left(\theta;[h_{t},x_{t+1},y_{t+1}]\right)$ using our LLM’s predictive distribution, we are also directly relying on its expected entropy being a meaningful measure of design quality.
-If $\theta$ is high–dimensional and predominantly free–form, the resulting entropy produced by the LLM is unlikely to be meaningful and using the data–estimation pairing is unlikely to produce an effective strategy.
-However, if $y$ is instead quite constrained, the LLM can produce a meaningful entropy over it, and choosing a model based on the prior–likelihood pairing is likely to implicitly define a meaningful distribution, and thus entropy, on $\theta$.
-Conversely, if $\theta$ is constrained and $y$ is free form, the opposite will hold instead.
-A.2 Choice of $\theta$
-An important corollary of this reasoning is that it can be important to be careful in our choice of exactly what we take $\theta$ to be, especially if we are using the data–estimation formulation.
-In particular, it is essential for entropy in the space of $\theta$ to form a meaningful notion of uncertainty, even if this entropy is not being measured through the LLM’s predictive distribution of $\theta$ directly.
-Thus, while $\theta$ inherently represents what we are trying to learn about and should always be set up as such, if there is flexibility in how exactly we formulate it, we should be careful to choose a form that yields an appropriate uncertainty measure.
-For example, if the LLM is trying to clarify what code a user wishes it to generate, we could either choose $\theta$ to be the code itself or, following (Neiswanger et al,, 2021; Bickford Smith et al,, 2023), the output the code produces.
-Here the entropy over code outputs induced by our distribution on code is likely to be a much better measure of uncertainty than the entropy of the raw code itself, given that there are multiple ways one can code the same operation.
-A.3 Expected uncertainty reduction in the sequential setting
-In Section˜4.4, we looked at whether the prior–likelihood pairing or the data–estimation pairing is better aligned with how we update our beliefs in practice.
 The picture here can get a somewhat more complicated once we move into the sequential BED setting.
@@
 The coherence of Bayesian updating means that the converse is unlikely to be true, so this provides further evidence towards using the prior–likelihood formulation.
-Appendix B Generating candidate hypotheses for BED-LLM
-A fundamental challenge for BED-LLM and the same method with the Entropy baseline is generating a sufficiently diverse set of candidate hypotheses from the LLM prior.
-As illustrated in Figure˜8, the raw distribution $p_{\mathrm{LLM}}(\theta)$ is highly over–collapsed, often concentrating mass on only a few high–likelihood hypotheses.
-Thus, it is not practical to directly use the LLM’s distribution as a prior $p(\theta)$.
-Instead, we jointly sample candidate $\theta$ and assume a uniform distribution over them.
-We can view this as sampling thetas from a mixture distribution.
-Instead, we employ a joint sampling strategy, in which the LLM is prompted to generate a list of $N$ hypotheses in a single rollout, which corresponds to drawing from the autoregressive list distribution
-$p_{\mathrm{LLM}}(\theta^{(1)}_{t},\ldots,\theta^{(N)}_{t};h_{t})\;=\;\prod_{n=1}^{N}p_{\mathrm{LLM}}(\theta_{t}^{(n)};[\theta_{t}^{(1:n-1)},h_{t}])$
-Each such joint sample is then filtered for history consistency using $p_{\mathrm{LLM}}(h_{t})$.
-This may be repeated.
-The surviving hypotheses from the current rollout are then merged with those retained from previous steps to form a cumulative candidate set $\Theta$.
-While one could in principle reweight candidates using importance sampling, in practice we choose to not rely on the model’s internal probabilities.
-Instead approximate the prior as a uniform distribution over this union
-$p_{f}(\theta;h_{t})\approx\frac{1}{\Theta}\sum_{\theta\in\Theta}\delta_{\theta}.$
-To further improve coverage and diversity, we incorporate several refinements in the generation of $p_{f}(\theta;h_{t})$.
-Due to the differences between different LLMs, the ways this is encouraged may differ, but we find in general that the following are effective:
-  • Repeat sampling and filtering: The sampling, filtering, and merging steps may be repeated to encourage diversity and ensure filtering correctness.
-  • Prompts: Prompts are designed to elicit stratified hypotheses by encouraging the LLM to consider different semantic features (e.g.
-age groups, genres, or categories) and implicitly diversify across them.
-  • Context ordering: During generation, we reverse the order of the question–answer pairs in $h_{t}$ to place the most recent question at the top of the context window (while retaining earlier exchanges), ensuring that specific constraints are prioritized and mitigating context drift.
-Finally, we note that different LLMs respond differently to strategies aiming to increase diversity: some benefit more from a higher temperature while others benefit from more repetitions of the sampling–filtering cycle.
-Appendix C Data–estimation method
-Our Data–Estimation method is based on a model derived from a data–estimation pairing (Section˜3).
-C.1 EIG Estimation
-Suppose our model is given by $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}p\left(y_{t+1};[h_{t},x_{t+1}]\right)p\left(\theta;[h_{t},x_{t+1},y_{t+1}]\right)$}.
-Here, it will clearly be beneficial to directly use Equation˜2 for estimating the EIG, as here we directly have access to all the required terms, other than $p(\theta)$ which can be simply ignored as it is not a function of $x$ so does not affect optimization of the question.
-If the possible values for $y$ are enumerable and we can evaluate $p\left(y_{t+1};[h_{t},x_{t+1}]\right)$ in closed–form, we can directly calculate the exact EIG (up to a constant) without requiring any estimation at all:
-$\text{EIG}_{\theta}(x)-\mathrm{Const}=-\sum_{y}p\left(y_{t+1};[h_{t},x_{t+1}]\right)\mathrm{H}\left[p\left(\theta;[h_{t},x_{t+1},y_{t+1}]\right)\right],$ (8)
-where the entropy $\mathrm{H}\left[p\left(\theta;[h_{t},x_{t+1},y_{t+1}]\right)\right]$ can be evaluated directly from the logits of the LLM, or if these are not available, estimated by sampling.
-If we cannot enumerate $y$ or evaluate $p\left(y_{t+1};[h_{t},x_{t+1}]\right)$, we can simply instead resort to Monte Carlo and use the estimator:
-$\text{EIG}_{\theta}(x)-\mathrm{Const}\approx-\frac{1}{N}\sum_{n=1}^{N}\mathrm{H}\left[p(\theta;[c,x,y_{n}])\right]\quad\mathrm{where}\quad y_{n}\sim p\left(y_{t+1};[h_{t},x_{t+1}]\right).$ (9)
-We provide an overview of how to do this in the LLM setting in Algorithm˜1.
+Appendix C Extended related work
+Information-based question answering with LLMs
+Several recent works have (explicitly or implicitly) looked at information gathering with LLMs.
+Most of these can be framed in a BED setting, with a deterministic likelihood (Piriyakulkij et al., 2023; Hu et al., 2024; Kobalczyk et al., 2025; Cooper et al., 2025), and can be seen as variants of our Split baseline.
+Piriyakulkij et al.
+(2023) use a deterministic 0/1 answer likelihood $p(a|x,q)$ via the LLM to prune items from a pre-enumerated finite set given a candidate question $q$.
+The question is selected by minimizing expected posterior entropy.
+They model user preferences with a binary ground truth, which would not be applicable in preference-elicitation scenarios with nebulous user profiles.
+Similarly,  Hu et al.
+(2024) use a deterministic likelihood to minimize entropy over a finite set $\Omega$ in a closed-world setting.
+Kobalczyk et al.
+(2025) target ambiguous task specifications in open-ended generation tasks by sampling a set of hypotheses (placing a uniform prior over them) and viewing each question as a deterministic partition over those samples, looking for questions that split the samples roughly evenly.
+Cooper et al.
+(2025) compute posterior entropy over a working set of top-k hypotheses (without filtering) through heuristic pruning.
+Wang et al.
+(2025) avoid the pitfall of deterministic likelihoods.
+They use a data–estimation framework to estimate EIG, focusing on scenarios where the target can be expressed as a predefined series of multiple-choice questions.
+Their approach relies on meta-training a predictive language model on historic question/answer pairs, and so is not directly comparable with BED-LLM which requires no additional training or data.
+Chan et al.
+(2025) do not model likelihoods or posterior beliefs, instead they rely on the expected size of conformal prediction sets as a surrogate uncertainty metric.
+This requires the use of an additional calibration dataset, and is confined to closed-world settings with a finite label set and pre-defined queries.
+Post-training LLMs for improved information gathering
+Rather than augmenting a frozen LLM with the ability to estimate utility functions, some works have instead aimed to post-train an LLM to improve its ability to ask questions (Zhang et al., 2024; Wu et al., 2025; Andukuri et al., 2024).
+Most do not explicitly consider informativeness of questions: Zhang et al.
+(2024) and Wu et al.
+(2025) use reinforcement learning techniques to reward generations that quickly lead to the correct answer, and Andukuri et al.
+(2024) builds on Li et al.
+(2025b) by fine-tuning on successful traces.
+Mazzaccara et al.
+(2024) do indirectly incorporate uncertainty, also using a deterministic likelihood: they use predictive entropy to identify informative questions, and then either fine-tune on the highest-entropy question, or perform DPO comparing the highest-entropy question with a lower-entropy question.
+We do not address fine-tuning in this work, focusing instead on exploring the correct way to formulate BED using LLMs.
+Combining LLMs with parametric models
+As discussed in Section˜4, a key challenge in adapting BED to the LLM setting is in aligning the expected information gain with the actual uncertainties extracted from the LLM after updating.
+Handa et al.
+(2024) take a different approach to this problem by using the LLM to generate features for an external conventional Bayesian joint model (in their case, a linear Bradley–Terry model), rather than deriving their joint model more directly from the LLM itself.
+This can be a good choice when the problem is well-bounded and we already have a well-specified Bayesian model form for the problem at hand; however, this may be challenging in arbitrarily large and complex hypothesis spaces.
+In particular, their specific method is not applicable more widely beyond the preference learning context they consider.
+BED
+It has been noted that the traditional sequential BED approach can sometimes be suboptimal in practice, as it only optimizes the EIG of the next observation, without planning ahead for the fact that design decisions taken at a given step can also influence the achievable EIGs from future steps (Foster, 2021).
+A variety of policy–based BED approaches have subsequently been proposed to address this (Foster et al., 2021; Ivanova et al., 2021; Blau et al., 2022; Huan & Marzouk, 2016; Hedman et al., 2025), while also removing the need to make model updates and conduct optimizations during the experiment itself.
+Our findings are complementary: by providing more faithful model factorizations, belief updates, and EIG estimators in the LLM setting, BED-LLM could supply stronger building blocks for policy-based methods, reducing variance, enhancing effectiveness, and improving the sample efficiency of policy training.
 Algorithm 1 Data–Estimation Selection at Turn $t$
-1:History $h_{t}$; candidate questions $\mathcal{X}_{\text{cand}}^{t}$; answer sets $\{\mathcal{Y}(x)\}$
+1:History $h_{t-1}$; candidate questions $\mathcal{X}_{\text{cand}}^{t}$; answer sets $\{\mathcal{Y}(x)\}$
 2:Selected question $x_{t}^{\star}$
 3:for each $x\in\mathcal{X}_{\text{cand}}^{t}$ do
-4:  Obtain predictive answer distribution $p_{\text{LLM}}(y;x,h_{t})$ for all $y\in\mathcal{Y}(x)$
+4:  Obtain predictive answer distribution $p_{\text{LLM}}(y;x,h_{t-1})$ for all $y\in\mathcal{Y}(x)$
 5:  for each $y\in\mathcal{Y}(x)$ do
-6:   Construct $\Theta_{t,y}\sim p_{f}(\theta;h_{t}\cup(x,y))$
-7:   Compute entropy $H_{y}\leftarrow\mathrm{H}\!\left[p_{\mathrm{LLM}}(\theta;h_{t},x,y)\right]$
-8:  end for
-9:  Compute $\text{EIG}(x)\leftarrow-\sum_{y\in\mathcal{Y}(x)}p_{\text{LLM}}(y;x,h_{t})\;H_{y}$
-10:end for
-11:Select $x_{t}^{\star}\leftarrow\arg\max_{x\in\mathcal{X}_{\text{cand}}^{t}}\text{EIG}(x)$
-12:return $x_{t}^{\star}$
-C.2 Generating candidate hypotheses
-To generate candidate values of $\theta$ for the data–estimation method, we used the prompt in Figure˜7.
-Figure˜8 shows an example of the distribution of samples obtained following two rounds in the 20–questions game.
-Note that the samples are highly concentrated on just a handful of answers.
-This lack of diversity shows that the model’s belief distribution is far more concentrated relative to the variability over valid hypotheses in the ground–truth task distribution, which negatively impacts the performance of the data–estimation method.
+6:   Compute entropy $H_{y}\leftarrow\mathrm{H}\!\left[p_{f}(\theta;h_{t-1}\cup(x,y))\right]$
+7:  end for
+8:  Compute $\text{EIG}(x)\leftarrow-\sum_{y\in\mathcal{Y}(x)}p_{\text{LLM}}(y;x,h_{t-1})\;H_{y}$
+9:end for
+10:Select $x_{t}^{\star}\leftarrow\arg\max_{x\in\mathcal{X}_{\text{cand}}^{t}}\text{EIG}(x)$
+11:return $x_{t}^{\star}$
+Appendix D Data–estimation method
+Our Data–Estimation method is based on a model derived from a data–estimation pairing (Section˜3).
+D.1 EIG Estimation
+Suppose our model is given by $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}p\left(y_{t};[h_{t-1},x_{t}]\right)p\left(\theta;[h_{t-1},x_{t},y_{t}]\right)$}.
+Here, it will clearly be beneficial to directly use Equation˜1 for estimating the EIG, as here we directly have access to all the required terms, other than $p(\theta)$ which can be simply ignored as it is not a function of $x$ so does not affect optimization of the question.
+If the possible values for $y$ are enumerable and we can evaluate $p\left(y_{t};[h_{t-1},x_{t}]\right)$ in closed–form, we can directly calculate the exact EIG (up to a constant) without requiring any estimation at all:
+$\text{EIG}_{\theta}(x)-\mathrm{Const}=-\sum_{y}p\left(y_{t};[h_{t-1},x_{t}]\right)\mathrm{H}\left[p\left(\theta;[h_{t-1},x_{t},y_{t}]\right)\right],$ (5)
+where the entropy $\mathrm{H}\left[p\left(\theta;[h_{t-1},x_{t},y_{t}]\right)\right]$ can be evaluated directly from the logits of the LLM, or if these are not available, estimated by sampling.
+If we cannot enumerate $y$ or evaluate $p\left(y_{t};[h_{t-1},x_{t}]\right)$, we can simply instead resort to Monte Carlo and use the estimator:
+$\text{EIG}_{\theta}(x)-\mathrm{Const}\approx-\frac{1}{N}\sum_{n=1}^{N}\mathrm{H}\left[p(\theta;[c,x,y_{n}])\right]\quad\mathrm{where}\quad y_{n}\sim p\left(y_{t};[h_{t-1},x_{t}]\right).$ (6)
+We provide an overview of how to do this in the LLM setting in Algorithm˜1.
+D.2 Generating candidate hypotheses
+To generate candidate values of $\theta$ for the data–estimation method, we use the prompt in Section˜D.2.
+Figure 4: Prompt for generating hypotheses (and evaluating their probability) for the data–estimation method.
 Return only the full name of one randomly selected famous person (living or deceased) consistent with the questions and answers above.
@@
 - No extra words, explanations, numbering, or punctuation beyond what’s in the name itself (hyphens/apostrophes allowed if part of the name).
-Figure 7: Prompt for generating hypotheses (and evaluating their probability) for the data–estimation method.
+Figure˜5 shows an example of the distribution of samples obtained following two rounds in the 20–questions game.
+Note that the samples are highly concentrated on just a handful of answers.
+This lack of diversity shows that the model’s belief distribution is far more concentrated relative to the variability over valid hypotheses in the ground–truth task distribution, which negatively impacts the performance of the data–estimation method.
 Is this person known for their contributions to science?
@@
   "Vincent van Gogh": 93,
+  "Salvador Dali": 44,
+  "Frida Kahlo": 37,
   "Georgia O Keeffe": 10,
-"Frida Kahlo": 37,
-"Salvador Dali": 44,
   "August Wilson": 8,
   "Auguste Rodin": 8
-Figure 8: An example of the sample distribution generated using the prompt in Figure˜7, conditioned on the two question/answer pairs at the top of this figure.
+Figure 5: An example of the sample distribution generated using the prompt in Section˜D.2, conditioned on the two question/answer pairs at the top of this figure.
 At this stage of the game, we independently sample 200 hypotheses from the LLM and record their frequencies.
@@
 This summary is for diagnostic purposes: Algorithm˜1 operates on the probabilities of individual samples and never instantiates such an aggregated summary.
-Appendix D Experimental details
-D.1 Generation of the Animals Dataset
-To create the Animals dataset, we prompted OpenAI o3 (OpenAI,, 2025, o3-2025-04-16) to generate a list of animals, using the prompt in Figure˜9.
-The resulting list is shown in Figure˜10.
+Appendix E Generating candidate hypotheses for BED-LLM
+Figure 6: Prompt for generating candidate hypotheses for the “Things” dataset.
+Similar prompts were used for “Celebrities” and “Animals”.
+You are playing a game of 20 Questions.
+Using all of the questions and answers so far:
+Generate up to {num_samples} candidate entities that satisfy every clue.
+Each candidate must be a single, self-contained entity (e.g., "Europa", "Bagpipe", "Diadem").
+List each entity on its own line - no numbering, punctuation, or extra text.
+Produce a varied set by identifying features not implied by the clues and diversifying along them.
+Do not repeat any entity.
+Return only the list of entities.
+A fundamental challenge for BED-LLM and its ablations is generating a sufficiently diverse set of candidate hypotheses from the LLM’s belief distribution, that are consistent with the previously-answered questions.
+Below, we detail the steps we take to construct our distribution over hypotheses.
+Candidate hypotheses are generated jointly, rather than independently.
+[…]
+$p_{\mathrm{LLM}}(\theta^{(1)}_{t},\ldots,\theta^{(N)}_{t};h_{t})\;=\;\prod_{n=1}^{N}p_{\mathrm{LLM}}(\theta_{t}^{(n)};[\theta_{t}^{(1:n-1)},h_{t}])\,.$
+We use a diversity-encouraging prompt
+We use a prompt designed to elicit stratified hypotheses by encouraging the LLM to consider different semantic features (e.g.
+age groups, genres, or categories) and implicitly diversify across them.
+An example prompt is shown in Appendix˜E.
+In our generation prompt, we reverse the order of the question–answer pairs in $h_{t}$ to place the most recent question at the top of the context window (while retaining earlier exchanges), ensuring that specific constraints are prioritized and mitigating context drift.
+For the 20 Questions experiments, we used a higher-than-normal temperature ($T=1.3$) to increase diversity of responses.
+For the preference elicitation experiments, we used $T=1$ to obtain more coherent responses.
+Candidates are filtered based on the history
+For each candidate, we use $p_{\mathrm{LLM}}(\theta;h_{t-1})$ to assess whether it is compatible with the previous question/answer pairs.
+We filter responses where the probability of the given answer falls below a certain threshold; in our experiments we set this threshold to 0.2.
+Valid candidates from previous generations are included
+We also filter the candidate hypotheses from the previous generation, based on the most recent question/answer pair, and include these in our candidate set.
+We repeat the generation process, keeping the previously generated and filtered samples in context to elicit new generations, either twice or three times if sufficient hypotheses have not been generated (noting the number of possible valid samples can be less than the number requested).
+We assume a uniform distribution over hypotheses
+While one could in principle reweight candidates using importance sampling, in practice we choose to not rely on the model’s internal probabilities.
+Instead we approximate the prior as a uniform distribution over this union
+$p_{f}(\theta;h_{t})\approx\frac{1}{|\Theta|}\sum_{\theta\in\Theta}\delta_{\theta}.$
+Finally, we note that different LLMs respond differently to strategies aiming to increase diversity: some benefit more from a higher temperature while others benefit from more repetitions of the sampling–filtering cycle.
+For fairness, in our experiments we have kept these parameters constant across models.
+Appendix F Experiment details for 20 Questions
+F.1 Problem sets
+We evaluate across three distinct problem sets—Animals, Celebrities, or Things—with each containing a mix of 100 obscure and common targets.
+Here, the problem set is just a list of different $\theta^{*}$ that will be individually provided to the answerer to instantiate different problems (e.g.
+we conduct a trial where $\theta^{*}=$“dog”, then one where $\theta^{*}=$“cat”, etc).
+The list of targets is never provided to the questioner model to restrict the set of possible hypotheses: the questioner is only prompted that is trying to identify an “animal”, “celebrity”, or “thing”.
+The problem sets are
+  • Animals: a set of animal species generated with OpenAI’s o3 model to ensure a diverse mix and balanced taxonomy.
+  • Celebrities: a diverse set of public figures, as used by Zhang et al. (2024).
+  • Things: a collection of everyday and exotic entities drawn from the web corpus, as used by Zhang et al.
+(2024); it covers a wide range of categories, from plants and clothing to professions, events, and mythical creatures.
+To create the Animals problem set, we prompted OpenAI o3 (OpenAI, 2025, o3-2025-04-16) to generate a list of animals, using the prompt in Section˜F.1.
+The resulting list is shown in Figure˜8.
 Alternative names (after $|$ character) were manually added.
-You are a zoologist.
-Please generate a list of 100 living animal species with very high taxonomic diversity, including diversity in phyla, classes, orders, and families.
-Present each animal on a different line.
-Figure 9: Prompt for Animals dataset generation.
+Figure 7: Prompt for Animals problem set generation. You are a zoologist.
+Please generate a list of 100 living animal species with very high taxonomic diversity, including diversity in phyla, classes, orders, and families. Present each animal on a different line.
 African elephant
@@
 Rosy boa
-Figure 10: Animals dataset (generated using OpenAI o3, with manual curation)
-D.2 LLM models details
-The open-weights LLM models used in Section˜6 were obtained from HuggingFace (HF) (Wolf et al,, 2020) and their repositories are: mistralai/Mistral-Large-Instruct-2411, Qwen/Qwen2.5-72B-Instruct, meta-llama/Llama-3.3-70B-Instruct.
-Since all models are instruction-finetuned, we used the latest chat templates provided as part of the HF implementation.
-All of the experiments using mistralai/Mistral-Large-Instruct-2411 were conducted by Oxford researchers on Oxford compute infrastructure.
-The OpenAI API models used were: gpt-4o-mini-2024-07-18, gpt-4o-2024-08-06.
-Appendix E Additional Personalization Experiments
-E.1 EIG estimates of questions chosen by different questioners
-Figure 11 demonstrates the progression of EIG estimates throughout the turns for BED-LLM and Naive QA.
-We see that at turn $t=1$, the questions chosen by both methods have similar EIG estimates, which explains the similar performance of BED-LLM and Naive QA at turn $t=1$.
-Figure 11: The average information gains for each strategy at each turn for the Qwen2.5-72B vs Qwen2.5-72B example.
-We see that the initial questions have similar information gains, which explains why the performance of the Naive QA method outperforms EIG at turn 1.
-Afterwards, the EIG information gains are much larger while the Baseline information gains decrease sharply.
+Figure 8: Animals problem set (generated using OpenAI o3, with manual curation)
+F.2 Evaluation
+[…]
+Figure 9: Full plots for 20 Questions Experiments.
+F.3 Algorithmic details
+Using our sample–then–filter process (see Section˜3.3), we aim to sample at least $N=15$ hypotheses, repeating the cycle up to three times if needed (the exact number of hypotheses can be less than this as it may not be possible to generate sufficient valid hypotheses, especially in later experiment turns).
+The questioner generates $M=15$ candidate questions to test, $\mathcal{X}^{\text{cand}}$, using the “conditional generation” approach of Section˜3.1 when possible, but falling back on “unconditional generation” if insufficient candidate hypotheses have been generated.
+Appendix G Experiment details for preference elicitation
+G.1 Problems
+To generate a set of ground-truth user profiles, we take a set of $200$ real user ratings from the MovieLens-100K dataset (Harper & Konstan, 2015), then use an “oracle” LLM (namely, OpenAI’s o3 model) to produce a paragraph of text that is consistent with each distinct set of ratings, using the prompt in Section˜G.3.
+As was the case for the 20 Questions problems, this problem set is never provided to the questioner and the set of allowed $\theta$ is not constrained.
+Because we need the LLM to be able to meaningfully capture uncertainty in the space of responses, we restrict questions to be multiple-choice.
+Specifically, the questioner is tasked with producing a question along with five possible responses A/B/C/D/E.
+We then define each $x_{t}$ to be the question coupled with the possible responses, and each $y_{t}$ to be one of the letters A/B/C/D/E to provide a restricted set of tokens over which we can measure entropy.
+Option E is further constrained to always be “none of the above” so that the answerer is not committed to choosing one of the directly generated choices if none are suitable.
+G.2 Evaluation
+To again allow tracking of progress through the experiment, after each turn of the interaction $t$, the questioner generates ten film recommendations, conditioned on $h_{t-1}$.
+These recommendations are checked for consistency with prior questions and answers; if any are judged inconsistent then they are removed and additional recommendations are generated.
+The quality of the film recommendations is then assessed using an “LLM-as-judge” protocol (Zhu et al., 2025; Trivedi et al., 2024).
+Namely, the answerer evaluates each of the 10 films recommended by the questioner, conditioned on the hidden target user profile $\theta^{*}$.
+It scores each film on a scale of 1 to 5 (in 0.5 increments), based on how well the recommendation aligns with $\theta^{*}$ — this score is output together with a brief justification to increase reliability.
+We report the mean rating and standard error across 200 users, over 5 question–answer turns.
+G.3 Algorithmic details
+For BED-LLM and Entropy, we compare $M=8$ candidate questions at each turn and we aim to generate at least $N=5$ candidate hypotheses.
+We use the “unconstrained generation” approach of candidate question generation (see Section˜3.1) as the user profiles can be quite diffuse and we are only generating a small number of possible hypotheses that can be quite easy to split.
+We note that data–estimation setup is not at all viable for this problem because the large number of tokens and varying dimensionality of each $\theta$ sample mean that $\mathrm{H}[p_{\mathrm{LLM}}(\theta;[h_{t-1},x_{t+1},y_{t+1}])]$ is not only infeasible to estimate, but also is not a meaningful measure of uncertainty.
+Figure 10: Prompt used to generate ground-truth user profiles for preference elicitation task.
+You will be given a user’s complete film rating history from the MovieLens dataset, provided as a dictionary structured by rating levels.
+Your task is to thoroughly analyze the user’s preferences across the entire range of their film ratings (from highest to lowest).
+Then, write a cohesive, descriptive paragraph (approximately 5–7 sentences) summarizing the user’s overall film taste profile.
+In your response, explicitly address:
+Favored Elements (inferred primarily from 4.5–5.0 ratings): Highlight the genres, narrative styles, themes, tones, historical eras, and emotional experiences that consistently resonate positively with this user.
+Avoid mentioning any specific film titles, characters, or explicit plot points.
+Neutral or Mixed Preferences (inferred from ratings around 2.5–4.0): Note if there are indications of genre overlap or conditional enjoyment, such as certain genres or styles they occasionally appreciate under specific circumstances.
+Disliked Elements (inferred primarily from 0.5–1.5 ratings): Clearly outline the genres, narrative characteristics, tones, or emotional impacts that the user consistently finds unappealing or poorly executed.
+Your paragraph must be precise, informative, nuanced, and balanced, effectively capturing the complexity and specificity of the user’s movie preferences.
+The resulting profile should be clear and detailed enough for a recommendation system to accurately predict the user’s likely enjoyment or dislike of other films based on their established patterns of taste.
+Proceed carefully, reasoning explicitly about the user’s overall rating patterns rather than relying exclusively on extreme ratings, to form a comprehensive, stable, and representative film preference profile.

### v3

@@
 This enables LLMs to act as effective multi-turn conversational agents and interactively interface with external environments.
-Our approach, which we call BED-LLM (Bayesian experimental design with large language models), is based on iteratively choosing questions or queries that maximize the expected information gain (EIG) about the task of interest given the responses gathered previously.
+Our approach, which we call BED-LLM (Bayesian experimental design with large language models), is based on iteratively choosing questions or queries that maximize the expected information gain (EIG) with respect to a variable of interest given the responses gathered previously.
 We show how this EIG can be formulated (and then estimated) in a principled way using a probabilistic model derived from the LLM’s predictive distributions and provide detailed insights into key decisions in its construction and updating procedure.
-We find that BED-LLM achieves substantial gains in performance across a wide range of tests based on the 20 Questions game and using the LLM to actively infer user preferences, compared to direct prompting of the LLM and other adaptive design strategies.
+We find that BED-LLM achieves substantial gains in performance across a wide range of tests based on the 20 Questions game and using the LLM to actively infer user preferences, compared to purely prompting-based design generation and other adaptive design strategies.
 1 Introduction
 Intelligent information gathering—the ability to ask the right questions at the right time—is fundamental to effective AI systems.
-However, despite their many successes, LLMs currently fall short on a crucial aspect of interactive intelligence: proactively seeking out information from a user or external environment in an intelligent and adaptive manner (Laban et al., 2025; Li et al., 2025c).
-For example, they have been shown to perform poorly on problems such as multi-turn guessing games (Bertolazzi et al., 2023; Zhang et al., 2024), task clarification (Chi et al., 2024), IT task automation (Jha et al., 2025), and multi-step tool use (Patil et al., 2025).
+However, despite their many successes, LLMs currently fall short on proactively seeking out information from a user or external environment in an adaptive manner (Laban et al., 2025; Li et al., 2025c).
+For example, they have been shown to perform poorly on multi-turn guessing games (Bertolazzi et al., 2023; Zhang et al., 2024), task clarification (Chi et al., 2024), IT task automation (Jha et al., 2025) and multi-step tool use (Patil et al., 2025).
 In particular, while modern LLMs are often capable of producing coherent and insightful questions (or other external queries) in a single-turn setting, they typically struggle to appropriately tailor their questions to previously gathered responses on interactive tasks (Bertolazzi et al., 2023; Patil et al., 2025).
@@
 They are also critical if we want to use LLMs in data gathering tasks or as automated agents in decision-making pipelines (Wu et al., 2025).
-In turn, these capabilities are essential across domains ranging from medical diagnosis (Hirosawa et al., 2024), troubleshooting (Jha et al., 2025), preference learning (Handa et al., 2024; Chakraborty et al., 2024; Ouyang et al., 2022), and tutoring systems (Kestin et al., 2025; Liu et al., 2024a), to conducting automated surveys (Aher et al., 2023; Lee et al., 2024; Jacobsen et al., 2025), and AI–driven scientific inquiries (Lu et al., 2024; Mandal et al., 2025).
-Note that in all these problems it is not enough for the LLM to generate full sets of suitable questions up front, we need it to be able to adaptively choose questions that are tailored to the already-collected user responses.
-We propose to address this challenge using the framework of sequential Bayesian experimental design (BED, Lindley (1956); MacKay (1992); Chaloner & Verdinelli (1995); Sebastiani & Wynn (2000); Rainforth et al.
-(2024)), which provides a model-based, information-theoretic mechanism for making adaptive design decisions, given a generative model of the experiment.
-Specifically, we show how the problem of interactive information gathering with LLMs can be formulated as a sequential experimental design problem with a model derived from the LLM, wherein we iterate between choosing queries based on maximizing their expected information gain (EIG) and updating our beliefs with the information from the received response.
-We call our approach BED-LLM and show how its success is critically dependent on our precise model formulation, belief updating procedure, and EIG estimation strategy.
+In turn, these capabilities are essential across domains ranging from medical diagnosis (Hirosawa et al., 2024), troubleshooting (Jha et al., 2025), preference learning (Chakraborty et al., 2024; Handa et al., 2024; Ouyang et al., 2022) and tutoring systems (Kestin et al., 2025; Liu et al., 2024a), to conducting automated surveys (Aher et al., 2023; Jacobsen et al., 2025; Lee et al., 2024) and AI-driven scientific inquiries (Lu et al., 2024; Mandal et al., 2025).
+Note that in all these problems it is not enough for the LLM to generate full sets of suitable questions up front: we need it to be able to adaptively choose questions that are tailored to the already-collected user responses.
+We propose to address this challenge using the framework of sequential Bayesian experimental design (BED; Chaloner & Verdinelli, 1995; Lindley, 1956; MacKay, 1992; Rainforth et al., 2024; Sebastiani & Wynn, 2000), which provides a model-based, information-theoretic mechanism for making adaptive design decisions, given a generative model of the experiment.
+Specifically, we show how the problem of interactive information gathering with LLMs can be formulated as a sequential experimental design problem with a model derived from the LLM, wherein we iterate between choosing queries by maximizing expected information gain (EIG) in a variable of interest and updating our beliefs with the information from the received response.
+We call our approach BED-LLM and show how its success is critically dependent on our precise model formulation, belief updating procedure and EIG estimation strategy.
 In particular, we show that it is essential to formulate the model with a precise distribution pairing that does not solely rely on in-context learning to update beliefs and uses the LLM’s uncertainties in the space of answers rather than the more complicated underlying hypothesis space we are trying to learn in.
@@
 Specifically, we first find that BED-LLM provides substantial improvements in the success rate for the 20 Questions problem across a variety of LLMs and target quantities.
-For example, when guessing celebrities with a small model, Llama-3.1-8B, we observe a 5.8x gain in success rate.
+On average, BED-LLM improves the success rate by 37.4 percentage points compared to direct prompting of the LLM, with the success rate more than doubling in over half of the setups considered and never decreasing.
 Second, we demonstrate noticeable improvements in using the LLM for movie recommendations, showing that these benefits hold even when the LLM’s predictive model differs from that of the answerer.
@@
 There are two natural ways to improve LLMs’ ability to gather information: modifying the model itself (e.g.
-via test-time- or post-training) or altering how the model is used at deployment time.
-We focus on the latter, since information–gathering tasks rarely provide task-specific data upfront (e.g.
-a user’s unknown preferences), and deployment–time methods avoid the cost and difficulty of finetuning an LLM altogether and are applicable to any existing LLM.
-However, we emphasize that improvements at the model level (e.g., Zhang et al., 2024) would be complementary to our approach.
+via finetuning) or altering how the model is used at deployment time.
+We focus on the latter, since information-gathering tasks rarely provide task-specific data upfront (e.g.
+a user’s unknown preferences), and deployment-time methods avoid the cost and difficulty of finetuning an LLM altogether and are applicable to any existing LLM.
+However, we emphasize that improvements at the model level (e.g.
+Zhang et al., 2024) would be complementary to our approach.
 To formalize the notion of information gathering, we need a concrete idea of what we wish to learn about.
@@
 We start with incomplete information about $\theta$, as represented by an initial belief distribution or prior, $p(\theta)$, but can refine these beliefs by making queries, $x\in\mathcal{X}$, to the user or some other external agent and receiving responses, $y\in\mathcal{Y}$, that are informative about $\theta$.
-Multiple such queries, $x_{1},\dots,x_{n}$, can be adaptively selected in a sequential decision-making process where we iteratively choose each $x_{t}$ based on the collected history $h_{t-1}:=(x_{i},y_{i})_{i=1:t-1}$.
+Multiple such queries, $x_{1},\dots,x_{T}$, can be adaptively selected in a sequential decision-making process where we iteratively choose each $x_{t}$ based on the collected history, $h_{t-1}:=(x_{i},y_{i})_{i=1}^{t-1}$.
 As our history grows, we will update our belief distribution to obtain $p(\theta;h_{t-1})$ via some model updating procedure.
@@
 Here, $h_{t-1}$ influences our distribution on $\theta$, but it is not derived via a joint distribution.
-In the LLM setting, there is considerable flexibility in how $p(\theta;h_{t-1})$ is constructed, as discussed in §3.3 and §4.
+In the LLM setting, there is considerable flexibility in how $p(\theta;h_{t-1})$ is constructed, as discussed in Section˜3.1 and Section˜4.
 While $p(\theta;h_{t-1})$ need not be explicitly defined, it provides the foundation for our information-theoretic method of query selection.
 For clarity of exposition, we focus on the case where the $x_{t}$ correspond to explicit questions asked to the user, but emphasize that the approach applies more broadly to other forms of external interaction by the LLM, such as retrieving documents or calling external functions.
-2.1 In–Context Updating of the belief distribution
+2.1 In-context updating of beliefs
 A natural and cheap way to incorporate the interaction history into the LLM is to include it in the context (Brown et al., 2020).
@@
 This is consistent with recent work that shows that in context updating does not treat all contextual information equally (Kossen et al., 2024; Liu et al., 2024b; Zhang et al., 2024).
-In §3.3, we introduce a more robust method for deriving $p(\theta;h_{t-1})$.
-2.2 Information–Theoretic Experimental Design
-The core of the BED framework is a joint generative model $p(\theta,y;x)$ over the target quantity $\theta$ and outcomes, $y$, given designs $x$.
-Most commonly, this is specified as a Bayesian model using a prior $p(\theta)$ and likelihood $p(y|\theta;x)$.
-In the general case, designs are then chosen to maximize the expectation of some utility function $U(\theta,y,x)$ under this model: we choose $x^{*}=\mathrm{argmax}_{x}\mathbb{E}_{p(\theta,y;x)}\left[U(\theta,y,x)\right]$.
-The most common choice is to take $U(\theta,y,x)=\log p(\theta,y;x)-\log p(\theta)\log p(y;x)$, where $p(\theta)$ and $p(y;x)$ are the marginal distributions on $\theta$ and $y$ implied by our joint model and we have assumed that our current beliefs on $\theta$ are independent of the design $x$.
+In Section˜3.1, we introduce a more robust method for deriving $p(\theta;h_{t-1})$.
+2.2 Information-theoretic experimental design
+The core of the BED framework is a joint generative model, $p(\theta,y;x)$, over the target quantity, $\theta$, and the outcome, $y$, corresponding to a design, $x$.
+Typically this is specified as a Bayesian model using a prior, $p(\theta)$, and a likelihood, $p(y|\theta;x)$.
+In the general case, designs are chosen to maximize the expectation of some utility function, $U(\theta,y,x)$, under this model: we choose $x^{*}=\mathop{\mathrm{arg\,max}}_{x}\mathbb{E}_{p(\theta,y;x)}\left[U(\theta,y,x)\right]$.
+The most common choice is to take $U(\theta,y,x)=\log p(\theta,y;x)-\log p(\theta)\log p(y;x)$, where $p(\theta)$ and $p(y;x)$ are the marginal distributions on $\theta$ and $y$ implied by our joint model and we have assumed that our current beliefs on $\theta$ are independent of the design, $x$.
 This leads to an objective corresponding to the expected information gain (EIG) in $\theta$ (Lindley, 1956; 1972),
-$\text{EIG}_{\theta}(x)$ $=\mathrm{H}[p(\theta)]-\mathbb{E}_{p(y;x)}[\mathrm{H}[p(\theta|y;x)]]$ (1)
+$\mathrm{EIG}_{\theta}(x)$ $=\mathrm{H}[p(\theta)]-\mathbb{E}_{p(y;x)}[\mathrm{H}[p(\theta|y;x)]]$ (1)
 $=\mathrm{H}[p(y;x)]-\mathbb{E}_{p(\theta)}[\mathrm{H}[p(y|\theta;x)]],$ (2)
-where $\mathrm{H}$ denotes the Shannon entropy (i.e., $\mathrm{H}[p(\theta)]=-\mathbb{E}_{p(\theta)}[\log p(\theta)]$).
-We can thus equivalently think of the EIG as: a) the mutual information between $\theta$ and $y$, b) the expected information gain over possible data simulated from our model (where the information gain is defined as the reduction in entropy from our prior on $\theta$ to the posterior), or c) the expected reduction in entropy over data from observing $\theta$ simulated from our prior (Sebastiani & Wynn, 2000).
-Working with the EIG is highly suited to a sequential or adaptive design approach, wherein it is generally referred to as sequential BED or Bayesian Adaptive Design (Rainforth et al., 2024).
+where $\mathrm{H}$ denotes the Shannon entropy (i.e.
+$\mathrm{H}[p(\theta)]=-\mathbb{E}_{p(\theta)}[\log p(\theta)]$).
+We can thus equivalently think of the EIG as (a) the mutual information between $\theta$ and $y$, (b) the expected reduction in entropy over $\theta$ (i.e.
+information gain in $\theta$) from observing data simulated from our model, or (c) the expected reduction in entropy over data from observing $\theta$ simulated from our prior (Sebastiani & Wynn, 2000).
+Working with the EIG is highly suited to a sequential or adaptive design approach, generally referred to as sequential BED or Bayesian adaptive design (Rainforth et al., 2024).
 Because the EIG is only a function of our underlying model, when we update the model as new data becomes available, our EIG design objective will naturally update as well.
-Specifically, to derive the incremental EIG (Cavagnaro et al., 2010) for the $t$-th query, $\text{EIG}_{\theta}(x_{t};h_{t-1})$, we simply replace the joint $p(\theta,y;x)$ in the above formulation with the updated joint $p(\theta,y_{t};h_{t-1},x_{t})$, with all marginals a conditionals derived from this (e.g.
+Specifically, to derive the incremental EIG (Cavagnaro et al., 2010) for the $t$-th query, $\mathrm{EIG}_{\theta}(x_{t};h_{t-1})$, we simply replace the joint in the above formulation, $p(\theta,y;x)$, with the updated joint, $p(\theta,y_{t};h_{t-1},x_{t})$, with all marginals and conditionals derived from this (e.g.
 $p(y;x)$ becomes $p(y_{t};h_{t-1},x_{t})$).
 Here this updated joint conventionally comes from a Bayesian update of the original model.
-However, in many cases, this is not practical and other non-Bayesian updates are performed instead, e.g.
-in active learning the update often actually corresponds to retraining the model with the new data (Gal et al., 2017; Bickford Smith et al., 2023).
-3 Sequential Bayesian Experimental Design for LLMs
-The sequential BED framework described in §2.2 requires two core components to be specified by the user: a) an initial joint model $p(\theta,y;x)$ over hypotheses $\theta$ and outcomes $y$, given chosen experiment $x$, and b) a procedure to derive an updated model $p(\theta,y_{t};h_{t-1},x_{t})$ after observing $h_{t-1}$.
-In the LLM setting, there is significant flexibility in these critical design decisions.
+However, in many cases, this is not practical and other non-Bayesian updates are performed instead.
+For example, in active learning the update often actually corresponds to retraining the model with the new data (Bickford Smith et al., 2023; Gal et al., 2017).
+3 Sequential Bayesian experimental design with LLMs
+The sequential BED framework described in Section˜2.2 requires two core components to be specified by the user: (a) an initial joint model, $p(\theta,y;x)$, over hypotheses, $\theta$, and outcomes, $y$, given a design, $x$; and (b) a procedure to derive an updated model, $p(\theta,y_{t};h_{t-1},x_{t})$, after observing $h_{t-1}$.
+In the LLM setting, there is significant flexibility in these critical methodological decisions.
 In particular, there are many ways to derive a suitable joint distribution from the LLM and its ability to learn in-context provides opportunities for update methods that go beyond standard Bayesian model updates.
@@
 A major challenge in the LLM setting is that unlike conventional probabilistic models, in general, $p_{\mathrm{LLM}}(\theta)\,p_{\mathrm{LLM}}(y;[\theta,x])\neq p_{\mathrm{LLM}}(y;x)\,p_{\mathrm{LLM}}(\theta;[x,y])$.
-That is, we induce a different joint distribution if we first sample $\theta$ then sample $y$ with $\theta$ in context (which we refer to as the prior–likelihood pairing), than if we first sample $y$ then sample $\theta$ with $y$ in context (data–estimation pairing).
+That is, we induce a different joint distribution if we first sample $\theta$ then sample $y$ with $\theta$ in context (which we refer to as the prior–likelihood pairing) versus if we first sample $y$ then sample $\theta$ with $y$ in context (data–estimation pairing).
 Moreover, we can deviate from the distribution directly induced by the LLM on one or both variables.
 The success of using BED with LLMs turns out to be critically dependent on these choices.
-We delay proper discussion of this complex issue until §4, where we will see that the preferable setup can depend on problem setting and, in particular, the relative complexity of spaces of $\theta$ and $y$.
-For now, we will focus on using the prior–likelihood pairing; we will argue in §4 that this is the advantageous setup in many practical scenarios.
-While we will generally use the LLM’s directly induced distribution for the likelihood, we allow the prior to deviate from this in a problem–specific manner.
+We delay proper discussion of this complex issue until Section˜4, where we will see that the preferable setup can depend on problem setting and, in particular, the relative complexity of spaces of $\theta$ and $y$.
+For now, we will focus on using the prior–likelihood pairing; we will argue in Section˜4 that this is the advantageous setup in many practical scenarios.
+While we will generally use the LLM’s directly induced distribution for the likelihood, we allow the prior to deviate from this in a problem-specific manner.
 As such, our initial joint model will be $p(\theta,y;x)=p(\theta)p_{\mathrm{LLM}}(y;[\theta,x])$.
@@
 Optimally updating the joint model in this setting requires incorporating new observations in a way that both fully captures the information from new data and is computationally tractable.
-At one extreme, we could perform full Bayesian updates via approximate inference, as in classical sequential BED.
-However, this demands a prohibitively large number of LLM evaluations to accurately approximate the posterior, and it does not exploit the power of the LLM as a probabilistic generative model, where autoregressive sequential rollouts often lead to more nuanced and diverse behavior than repeated static likelihood queries.
-At the other extreme, simple in-context updating, $p(\theta;h_{t-1})=p_{\mathrm{LLM}}(\theta;h_{t-1})$, is cheap but, as we show later, fails to reliably capture information from new data, leading to inconsistent belief states and undermining the sequential BED approach.
-As we discuss in §3.3, we therefore employ a strategy that is somewhere between the two: drawing samples in a way that utilizes $p_{\mathrm{LLM}}(\theta;h_{t-1})$ while encouraging diversity, then filtering out samples that are actually not compatible with $h_{t-1}$ and renormalizing.
+At one extreme, we could target full Bayesian updates via approximate inference, as in classical sequential BED.
+However, this demands a prohibitively large number of LLM evaluations to accurately approximate the posterior, and it does not exploit the power of the LLM as a probabilistic generative model, where autoregressive sequential rollouts often lead to more nuanced and diverse behavior than repeated static queries.
+At the other extreme, simple in-context updating, with $p(\theta;h_{t-1})=p_{\mathrm{LLM}}(\theta;h_{t-1})$, is cheap but, as we show later, fails to reliably capture information from new data, leading to inconsistent belief states and undermining the sequential BED approach.
+As we discuss in Section˜3.1, we therefore employ a strategy that is somewhere between the two: drawing samples in a way that utilizes $p_{\mathrm{LLM}}(\theta;h_{t-1})$ while encouraging diversity, then filtering out samples that are actually not compatible with $h_{t-1}$ and renormalizing.
 We refer to the resulting distribution as $p_{f}(\theta;h_{t-1})$.
-We do not update our likelihood model $p_{\mathrm{LLM}}(\theta;h_{t-1})$; see Section˜A.1 for a discussion.
+We do not update our likelihood model, $p_{\mathrm{LLM}}(\theta;h_{t-1})$; see Section˜F.3 for empirical comparison of updated vs.
+static likelihoods in our experiments, and a discussion.
 BED-LLM
 We now introduce our specific algorithmic approach, BED-LLM.
-Here, the queries will correspond to our designs, $x$ (assumed to be in form of questions posed to the user in the following for simplicity, but could also be, e.g., external function calls, document retrieval, web search, etc), and the responses received will correspond to our outcomes, $y$.
-Using the LLM to derive joint models over these outcomes and the target variables $\theta$ given histories, $h_{t-1}$ as described above, we can interleave choosing informative questions by optimizing the incremental EIG, $\text{EIG}_{\theta}(x_{t};h_{t-1})$, and updating our underlying model based on the received question–response pairs.
-Specifically, BED-LLM iterates over the following key steps, where $t$ indexes the current turn:
-  1. 1.
-    Generate candidate questions3.1): Propose a candidate set of $M$ diverse, multiple-choice questions, $\mathcal{X}^{\text{cand}}$, by appropriate sampling of the LLM based on the conversational context $h_{t-1}$.
-  2. 2.
-    Compute EIG estimator3.2): For each candidate $x_{t}\in\mathcal{X}^{\text{cand}}$, estimate $\text{EIG}_{\theta}(x_{t};h_{t-1})$.
-  3. 3.
-    Select and ask optimal question: Choose the question $x_{t}\in\mathcal{X}^{\text{cand}}$ that yields the highest estimated EIG.
-Pose $x_{t}$ to the user, observe response $y_{t}$, and update the history, $h_{t}=(h_{t-1},(x_{t},y_{t}))$.
-  4. 4.
-    Construct updated joint (Section˜3.3): $p(\theta,y_{t+1};h_{t},x_{t+1})=p_{f}(\theta;h_{t})p_{\mathrm{LLM}}(y_{t+1};[\theta,x_{t+1}])$, using the new history and return to Step 1 (unless a termination criterion has been achieved).
-Note our belief state on $\theta$ after the $t$-th turn is simply given by $p_{f}(\theta;h_{t})$.
-3.1 Generating Candidate Questions
-As it is not computationally feasible to directly optimize over the space of possible questions, we rely on using the LLM to propose diverse candidate questions, $\mathcal{X}^{\text{cand}}$, then select the best question from these.
-We consider two specific approaches: 1) Unconstrained generation.
-Given $h_{t-1}$, the LLM is simply asked to propose new questions by sampling from $p_{\mathrm{LLM}}(x_{t};h_{t-1})$ with appropriate prompting.
-2) Conditional generation.
-The LLM is given both $h_{t-1}$ and a generated set of hypotheses $\Theta^{\text{cand}}=\{\theta^{(n)}\}_{n=1}^{N}$, such that we sample from $p_{\mathrm{LLM}}(x_{t};[h_{t-1},\Theta^{\text{cand}}])$.
-Specifically, the LLM is prompted to propose questions that “slice” the hypothesis pool into roughly balanced subsets.
-For both strategies, we sample $M$ questions jointly with a relatively high temperature to encourage diversity.
-Conditional generation allows us to “guide” the LLM to propose highly informative questions.
-However, it risks overfitting to $\Theta^{\text{cand}}$.
-In practice, we find it is effective for discrete spaces (§6.1), but less so for spaces with complex, overlapping hypotheses (§6.2).
-We restrict questions to multiple-choice format to allow $p_{\mathrm{LLM}}\left(y_{t};\left[\theta,x_{t}\right]\right)$ to produce well-calibrated probabilities (see §4).
-3.2 Estimating EIG for each question
-To estimate the EIG based on Equation 2 for a given question $x_{t}$, we derive the following Rao-Blackwellized estimator based on the LLM’s predictive distribution:
-$\begin{split}\text{EIG}_{\theta}(x_{t};h_{t-1})\approx&\textstyle\frac{1}{N}\sum_{n=1}^{N}\sum_{y_{t}\in\mathcal{Y}}p_{\mathrm{LLM}}(y_{t};[\theta^{(n)},x_{t}])\log p_{\mathrm{LLM}}(y_{t};[\theta^{(n)},x_{t}])\\ &-\textstyle\sum_{y_{t}\in\mathcal{Y}}\hat{p}(y_{t};[h_{t-1},x_{t}])\log\hat{p}(y_{t};[h_{t-1},x_{t}]),\end{split}$ (3)
-where $\hat{p}(y_{t};[h_{t-1},x_{t}])\!:=\!\frac{1}{N}\sum_{n=1}^{N}p_{\mathrm{LLM}}(y_{t};[\theta^{(n)},x_{t}])$ and $\theta^{(n)}\!\sim\!p_{f}(\theta;h_{t-1})$, see §3.3.
-This estimator has been used in other BED contexts (Gal et al., 2017; Rainforth, 2017).
-Note that the samples do not need to be independent for this estimator to converge, provided they satisfy some appropriate form of ergodicity or decaying correlation (see, e.g., Billingsley (2013)).
-When constructing this estimator, we compute the $p_{\mathrm{LLM}}(y_{t};[\theta^{(n)},x_{t}])$ terms using the LLM’s logits whenever possible.
-By the Rao-Blackwell theorem, this always produces lower variance than purely sample–based estimators (Rao et al., 1945), like those employed in Hu et al.
-(2024) and Kobalczyk et al.
-(2025).
-Avoiding deterministic likelihood assumptions
-Previous attempts to apply information criteria to choosing queries in LLMs have generally assumed responses are deterministic given $(\theta,x_{t})$ (Cooper et al., 2025; Kobalczyk et al., 2025; Hu et al., 2024; Mazzaccara et al., 2024; Piriyakulkij et al., 2023).
-Under this assumption, the EIG simplifies to the marginal predictive entropy, $\mathrm{H}[p(y_{t};x_{t},h_{t-1})]$.
-This is problematic as, in practice, the expected likelihood entropy will vary with $x_{t}$.
-For example, if $\theta=$“Dog”, then the response to the question “Is it an animal?” should be close to deterministic, but the expected response to the question “Does it have black fur?” clearly is not.
-In general, $\mathbb{E}_{p(\theta;h_{t-1})}[\mathrm{H}[p(y_{t}|\theta;x_{t},h_{t-1})]]$ measures how certainly the question can be answered once $\theta$ is known.
-Including it in our objective is essential in avoiding questions that are irrelevant, ambiguous, unclear, or simply unhelpful in our quest to learn about $\theta$.
-We provide an illustrative example of this in Section˜A.2.1 and empirical evidence for this in §6.
-Given that approximating the EIG with marginal predictive entropy does not provide meaningful computational savings (as it does not reduce the required number of LLM calls), we advise against making such deterministic likelihood assumptions.
+Here, the queries will correspond to our designs, $x$ (assumed to be in form of questions posed to the user in the following for simplicity, but could also be external function calling, document retrieval, web search, etc), and the responses received will correspond to our outcomes, $y$.
+Using the LLM to derive joint models over these outcomes and the target variable, $\theta$, given the history, $h_{t-1}$, as described above, we can interleave choosing informative questions by optimizing the incremental EIG, $\mathrm{EIG}_{\theta}(x_{t};h_{t-1})$, and updating our underlying model based on the received question–response pairs.
+Specifically, BED-LLM iterates over the following key steps, where $t$ indexes the current turn (see also Figure˜1).
+  1. (A)
+    Extract beliefs (Section˜3.1).
+Use joint model $p(\theta,y_{t};h_{t-1},x_{t})=p(\theta)p_{\mathrm{LLM}}(y_{t};[\theta,x_{t}])$, where $h_{0}=\emptyset$, for $t=1$ and $p(\theta,y_{t};h_{t-1},x_{t})=p_{f}(\theta;h_{t-1})p_{\mathrm{LLM}}(y_{t};[\theta,x_{t}])$ for $t>1$.
+  2. (B)
+    Generate candidate questions (Section˜3.2).
+Propose a candidate set of $M$ diverse, multiple-choice questions, $\mathcal{X}^{\mathrm{cand}}$, by appropriate sampling of the LLM based on the history, $h_{t-1}$.
+  3. (C)
+    Estimate EIG (Section˜3.3). For each candidate, $x_{t}\in\mathcal{X}^{\mathrm{cand}}$, estimate $\mathrm{EIG}_{\theta}(x_{t};h_{t-1})$.
+  4. (D)
+    Select and ask the best question. Choose $x_{t}$ by EIG maximization and pose to the user.
+  5. (E)
+    Update the history. Observe response, $y_{t}$, and update the history to $h_{t}=(h_{t-1},(x_{t},y_{t}))$.
+Current history: $h_{2}=((\text{``Born in the 20th century?''},\text{``Yes''}),(\text{``Is this person male?''},\text{``Yes''}))$.
+(A) Extract beliefs (Section˜3.1).
+Construct hypothesis set $\Theta^{\mathrm{cand}}=\{\theta_{n}\}_{n=1}^{N}$ by sampling candidate hypotheses, $\theta\sim p_{\mathrm{LLM}}(\theta;h_{2})$, and rejecting any $\theta$ that is inconsistent with $h_{2}$.
+Use joint model $p(\theta,y;h_{2},x)=p_{f}(\theta;h_{2})p_{\mathrm{LLM}}(y;[\theta,x])$, where $p_{f}(\theta;h_{2})$ is uniform over $\Theta^{\mathrm{cand}}$.
+Barack Obama Steve Irwin Hugh Laurie Banksy Elvis Presley (B) Generate candidate questions (Section˜3.2).
+Sample $\tilde{x}_{1:M}\sim p_{\mathrm{LLM}}(\tilde{x}_{1:M};[h_{t-1},\Theta^{\mathrm{cand}}])$.
+$\tilde{x}_{1}$: “Was this person born in Antarctica?” $\tilde{x}_{2}$: “Was this person born in the 19th century?” $\tilde{x}_{3}$: “Is this person an artist?” $\tilde{x}_{4}$: “Is this person European?” $\tilde{x}_{5}$: “Does this person prefer thrash metal over death metal?” (C) Estimate EIG (Section˜3.3).
+Compute $\underbrace{\mathrm{H}[\hat{p}(y;[h_{2},x])]}_{\text{marginal entropy}}-\underbrace{\tfrac{1}{N}\textstyle\sum_{n=1}^{N}\mathrm{H}[p_{\mathrm{LLM}}(y;[\theta_{n},x])]}_{\text{expected conditional entropy}}\approx\mathrm{EIG}_{\theta}(x;h_{2})$.
+Question Marg.
+Exp.
+cond.
+EIG Intuition $\tilde{x}_{1}$ Born in Antarctica?
+$\approx 0$ $\approx 0$ $\approx 0$ Answer “No” for all $\theta_{n}$ $\tilde{x}_{2}$ Born in 19th century?
+$\approx 0$ $\approx 0$ $\approx 0$ Redundant given $h_{2}$ $\tilde{x}_{3}$ An artist?
+$0.56$ $0.41$ $0.15$ Uneven split $\tilde{x}_{4}$ European?
+$0.97$ $0.17$ $0.80$ Balanced split; crisp answers $\tilde{x}_{5}$ Thrash vs.
+death metal?
+$0.89$ $0.88$ $0.01$ Uncertain even given $\theta$ (D) Select and ask the best question.
+Choose $x_{3}=\tilde{x}_{4}$ by maximizing EIG.
+Output this to the user.
+(E) Update the history.
+Observe answer, $y_{3}$.
+Set $h_{3}=(h_{2},(x_{3},y_{3}))$.
+Figure 1: BED-LLM applied to the 20 Questions game involves repeatedly constructing a belief state, generating candidate questions, estimating and maximizing EIG to select a question, and interacting with the user to gather new data.
+The contrast between $\tilde{x}_{4}$ and $\tilde{x}_{5}$ illustrates the benefit of using EIG with a non-deterministic likelihood: $\tilde{x}_{5}$ has high marginal entropy (answers are uncertain), but its expected conditional entropy is equally high (answers are uncertain even given $\theta$), so nearly nothing is expected to be learned, and $\tilde{x}_{4}$ should be favoured as a result.
+Numerical values are illustrative.
 3.1 Prior construction and belief updating
@@
 However, doing full Bayesian updates to our model as the history grows is generally impractical for computational reasons in the LLM setting, as it requires approximate inference and this, in turn, typically requires large numbers of expensive likelihood evaluations.
-Furthermore, the Savage axioms only hold if our (implied) prior truly represents our beliefs, but we find that $p_{\mathrm{LLM}}(\theta)$ is typically heavily overconfident on a small number of possible hypotheses and can struggle to convey the full range of possibilities even with careful prompting and a high temperature (c.f.
-Figure˜5).
+Furthermore, the Savage axioms only hold if our (implied) prior truly represents our beliefs, but we find that $p_{\mathrm{LLM}}(\theta)$ is typically heavily overconfident on a small number of possible hypotheses and can struggle to convey the full range of possibilities even with careful prompting and a high temperature (see Figure˜5).
 A natural tractable alternative is to derive our beliefs through LLM in-context updates, that is, use $p_{\mathrm{LLM}}(\theta;h_{t-1})$, noting that this has been shown to behave differently to Bayesian updating (Falck et al., 2024; Kossen et al., 2024).
@@
 First, we filter the generated hypotheses according to whether they are compatible with the history $h_{t-1}$.
-We do this by using the LLM to zero-shot check the compatibility of each sampled $\theta$ with all the previous question–answer pairs in $h_{t-1}$ (using $p_{\mathrm{LLM}}(y_{i};[\theta,x_{i}])\forall i=1:t-1$) and then rejecting that sample if an incompatibility is found.
+We do this by using the LLM to check the compatibility of each sampled $\theta$ with all the previous question–answer pairs in $h_{t-1}$ (using $p_{\mathrm{LLM}}(y_{i};[\theta,x_{i}])$ for $i\in(1,2,\ldots,t-1)$) and then rejecting that sample if an incompatibility is found.
 Specifically, a sample is rejected if the likelihood of an observed answer falls below a predefined threshold, chosen to balance robustness to model uncertainty against the need to enforce strict historical coherence.
@@
 Details of our exact setup for $p_{f}(\theta;h_{t-1})$ are given in Appendix˜E.
+3.2 Generating candidate questions
+As it is not computationally feasible to directly optimize over the space of possible questions, we rely on using the LLM to propose diverse candidate questions, $\mathcal{X}^{\mathrm{cand}}$, then select the best question from these.
+We consider two specific approaches.
+(a) Unconstrained generation: given $h_{t-1}$, the LLM is simply asked to propose new questions by sampling from $p_{\mathrm{LLM}}(x_{t};h_{t-1})$ with appropriate prompting.
+(b) Conditional generation: the LLM is given both $h_{t-1}$ and a generated set of hypotheses, $\Theta^{\mathrm{cand}}=\{\theta_{n}\}_{n=1}^{N}$, such that we sample from $p_{\mathrm{LLM}}(x_{t};[h_{t-1},\Theta^{\mathrm{cand}}])$; specifically, the LLM is prompted for questions that “slice” the hypothesis pool into roughly balanced subsets.
+For both strategies, we sample $M$ questions jointly with a relatively high temperature to encourage diversity.
+Conditional generation allows us to “guide” the LLM to propose highly informative questions.
+However, it risks overfitting to $\Theta^{\mathrm{cand}}$.
+In practice, we find it is effective for discrete spaces (Section˜6.1), but less so for spaces with complex, overlapping hypotheses (Section˜6.2).
+We restrict questions to multiple-choice format to simplify uncertainty quantification (see Section˜4).
+3.3 Estimating EIG for each question
+To estimate the EIG based on Equation 2 for a given question $x_{t}$, we use the following Rao-Blackwellized estimator based on the LLM’s predictive distribution:
+$\begin{split}\mathrm{EIG}_{\theta}(x_{t};h_{t-1})\approx&\textstyle\frac{1}{N}\sum_{n=1}^{N}\sum_{y_{t}\in\mathcal{Y}}p_{\mathrm{LLM}}(y_{t};[\theta_{n},x_{t}])\log p_{\mathrm{LLM}}(y_{t};[\theta_{n},x_{t}])\\ &-\textstyle\sum_{y_{t}\in\mathcal{Y}}\hat{p}(y_{t};[h_{t-1},x_{t}])\log\hat{p}(y_{t};[h_{t-1},x_{t}]),\end{split}$ (3)
+where $\hat{p}(y_{t};[h_{t-1},x_{t}])\!:=\!\frac{1}{N}\sum_{n=1}^{N}p_{\mathrm{LLM}}(y_{t};[\theta_{n},x_{t}])$ and $\theta_{n}\!\sim\!p_{f}(\theta;h_{t-1})$ (see Section˜3.1).
+This estimator has been used in other BED contexts (Gal et al., 2017; Rainforth, 2017).
+Note that the samples do not need to be independent for this estimator to converge, provided they satisfy some appropriate form of ergodicity or decaying correlation (see, e.g., Billingsley (2013)).
+When constructing this estimator, we compute the $p_{\mathrm{LLM}}(y_{t};[\theta_{n},x_{t}])$ terms using the LLM’s logits whenever possible.
+By the Rao-Blackwell theorem, this always produces lower variance than purely sample-based estimators (Rao et al., 1945), like those employed in Hu et al.
+(2024) and Kobalczyk et al.
+(2025).
+Avoiding deterministic likelihood assumptions
+Previous attempts to apply information criteria to choosing queries in LLMs have generally assumed responses are deterministic given $(\theta,x_{t})$ (Cooper et al., 2025; Kobalczyk et al., 2025; Hu et al., 2024; Mazzaccara et al., 2024; Piriyakulkij et al., 2023).
+Under this assumption, the EIG simplifies to the marginal predictive entropy, $\mathrm{H}[p(y_{t};x_{t},h_{t-1})]$.
+This is problematic as, in practice, the expected likelihood entropy will vary with $x_{t}$.
+In general, $\mathbb{E}_{p(\theta;h_{t-1})}[\mathrm{H}[p(y_{t}|\theta;x_{t},h_{t-1})]]$ measures how certainly the question can be answered once $\theta$ is known.
+Including it in our objective is essential in avoiding questions that are irrelevant, ambiguous, unclear, or simply unhelpful in our quest to learn about $\theta$.
+Figure˜1 illustrates this concretely—question $\tilde{x}_{4}$ has high marginal entropy and low conditional entropy (answers are crisp given $\theta$), yielding high EIG, while $\tilde{x}_{5}$ has equally high marginal entropy but also high conditional entropy (answers are noisy even given $\theta$), so the terms cancel and nearly nothing is learned.
+An entropy-only method would score both similarly and risk wasting a turn; the same failure mode causes the Entropy ablation to provide no improvement over Prompt-Only in our preference elicitation experiments (Section˜6.2).
+Given that both terms of Equation˜3 are computed from the same LLM likelihood evaluations, retaining the full EIG requires no additional calls and provides no computational savings.
+Thus, we advise against making deterministic likelihood assumptions.
 4 On the specification of $p(\theta,y_{t};h_{t-1},x_{t})$, and its implications
-As we described in §3, successfully applying sequential BED in the LLM setting hinges upon how we specify, and update, the joint distribution $p(\theta,y_{t};h_{t-1},x_{t})$.
-In particular, as previously highlighted, there are two distinct ways to derive the joint model from our LLM: using a prior–likelihood pairing, $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}p(\theta;h_{t-1})p\left(y_{t};\left[\theta,x_{t}\right]\right)$}, or a data–estimation pairing, $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}p\left(y_{t};[h_{t-1},x_{t}]\right)p\left(\theta;[h_{t-1},x_{t},y_{t}]\right)$}.
+As we described in §3, successfully applying sequential BED in the LLM setting hinges upon how we specify, and update, the joint distribution, $p(\theta,y_{t};h_{t-1},x_{t})$.
+In particular, as previously highlighted, there are two distinct ways to derive the joint model from our LLM: using a prior–likelihood pairing, $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}p\!\left(\theta;h_{t-1}\right)p\!\left(y_{t};\left[\theta,x_{t}\right]\right)$}, or a data–estimation pairing, $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}p\!\left(y_{t};[h_{t-1},x_{t}]\right)p\!\left(\theta;[h_{t-1},x_{t},y_{t}]\right)$}.
 The first construction mirrors deriving our beliefs about $\theta$ from a conventional Bayesian posterior with a concrete prior and likelihood derived (at least partially) from the LLM, whereas the second has analogies to a marginal-posterior approach (Fong et al., 2023; Falck et al., 2024) in that it that samples hypothetical data and draws inferences on $\theta$ given hypothetical data using in-context learning.
@@
 Modeling flexibility   The most obvious relative merits of the prior–likelihood and data–estimation pairings are in the flexibility in how each term is chosen.
-The prior–likelihood pairing gives us greater flexibility to construct a prior set of beliefs over $\theta$ that is distinct to the LLM’s internal beliefs, as it allows us to directly control this prior by changing $p(\theta;h_{t-1})$, whereas the prior is only implicitly defined in the data–estimation pairing.
+The prior–likelihood pairing gives us greater flexibility to construct a prior set of beliefs over $\theta$ that is distinct to the LLM’s internal beliefs, as it allows us to directly control this prior by changing $p\!\left(\theta;h_{t-1}\right)$, whereas the prior is only implicitly defined in the data–estimation pairing.
 In §3.1 we exploited this flexibility through our definition of $p_{f}(\theta;h_{t-1})$.
@@
 This difference becomes particularly noticeable when the complexities of the spaces of $\theta$ and $y$ differ significantly.
-Our ability to draw sensible samples of either will generally be quite robust to these spaces being complex or high–dimensional; this is where LLMs tend to thrive, effectively generating highly complex outputs in an autoregressive manner.
+Our ability to draw sensible samples of either will generally be quite robust to these spaces being complex or high-dimensional; this is where LLMs tend to thrive, effectively generating highly complex outputs in an autoregressive manner.
 However, evaluating the entropy of a distribution becomes dramatically harder as the dimensionality or complexity increases (Acharya et al., 2019; Paninski, 2003), and the entropy of the predictive distribution of an LLM in such cases will not typically provide a sensible measure of uncertainty (Kadavath et al., 2022; Desai & Durrett, 2020).
@@
 For the problems that we consider, the space of $y$ is less complex than that of $\theta$, indicating we should, in general, use the prior–likelihood formulation.
-However, in cases where this is not true, the data–estimation formulation may be preferable instead.
-We provide additional discussion on the impact of the chosen pairing on our entropy estimate, plus discussion on the choice of $\theta$, in Appendix˜B.
+However, in cases where this is not true, the data–estimation formulation may be preferable instead, see Appendix˜B for further discussion on this and on the choice of $\theta$.
 Extracting the belief state   A further advantage of the prior–likelihood construction is that our belief state on $\theta$ can be extracted directly as $p_{f}(\theta;h_{t-1})$.
@@
 5 Related work
-Several works have explored the baseline ability of LLMs to rapidly learn about a parameter of interest by asking questions (Zhang et al., 2024; Li et al., 2025b)—effectively our Naive baseline in §6.
+Several works have explored the baseline ability of LLMs to rapidly learn about a parameter of interest by asking questions (Zhang et al., 2024; Li et al., 2025b)—effectively our Prompt-Only baseline in §6.
 While these works demonstrate some ability to adaptively construct information-seeking questions, they often fail to extract important information (Li et al., 2025a).
 Some works have further specifically attempted to choose questions based on model-based uncertainty criteria (Piriyakulkij et al., 2023; Hu et al., 2024; Kobalczyk et al., 2025; Mazzaccara et al., 2024; Cooper et al., 2025).
-None of these works provide the same careful consideration of how the underlying joint model should be formulated, which underpins our own work, and they all assume deterministic likelihood models that mean their objectives correspond to a sample-based estimate of marginal predictive entropy in practice, as explained in Section˜3.2.
+None of these works provide the same careful consideration of how the underlying joint model should be formulated, which underpins our own work, and they all assume deterministic likelihood models that mean their objectives correspond to a sample-based estimate of marginal predictive entropy in practice, as explained in Section˜3.3.
 In general, these previous works have also required restrictions on the space of allowable hypotheses, $\theta$, and typically require additional assumptions and/or approximations.
@@
 $h_{t-1}$ and the questioner’s prompts).
-We test two questioner-answerer setups, where the two are served by separate instances of the same LLM, or two different LLMs.
+We test two questioner–answerer setups, where the two are served by separate instances of the same LLM, or two different LLMs.
 The latter scenario is important because in practice, the answerer will often follow a different distribution than the questioner’s internal model for reasoning about responses, thereby forming a model misspecification.
 Baselines
-We compare BED-LLM against two existing baseline methods: Naive and Split.
-Naive involves prompting the questioner to directly generate an informative next question, without explicit hypotheses generation or computing a data-acquisition objective, and then sampling the question with temperature T=1T\!=\!1; this was explored by Zhang et al.
+We compare BED-LLM against three baselines.
+Prompt-Only prompts the questioner to directly generate an informative next question, without explicit hypothesis generation or a data-acquisition objective, and then sampling the question with temperature T=1T\!=\!1; this was explored by Zhang et al.
 (2024).
-Split involves choosing the question that most equally splits a sampled set of hypotheses $\Theta^{\text{cand}}$, which corresponds to maximizing the marginal predictive entropy $\mathrm{H}[p(y_{t};x_{t},h_{t-1})]$ in a model with a deterministic likelihood.
+CoT augments Prompt-Only with a ReAct-style (Yao et al., 2023b) chain-of-thought: the LLM first produces a thought (what has been established, what information would help most) then an action (the question).
+This tests whether structured reasoning alone can close the gap to BED-LLM.
+Split chooses the question that most equally splits a sampled set of hypotheses, $\Theta^{\mathrm{cand}}$, which corresponds to maximizing the marginal predictive entropy, $\mathrm{H}[p(y_{t};x_{t},h_{t-1})]$, in a model with a deterministic likelihood.
 As such, the methods of Cooper et al.
@@
 (2024) and Piriyakulkij et al.
-(2023) can all be viewed as variants of this Split baseline.
-While Split is not applicable to the preference-elicitation scenario, where a deterministic likelihood is not viable, it is to our knowledge the state-of-the-art method for 20 Questions.
-We note that our own Split baseline implementation achieves dramatically better results than reported by, for example, Kobalczyk et al.
-(2025), so this constitutes a very strong baseline relative to previous work.
-On top of this, our implementation of Naive appears to signficantly improve over that of Zhang et al.
-(2024),
+(2023) can all be viewed as variants of this objective.
+To the best of our knowledge, Split represented the previous state-of-the-art method for 20 Questions.
+We note that our own specific Split baseline implementation, which uses BED-LLM’s filtering mechanism, also achieves dramatically better results than reported by, for example, Kobalczyk et al.
+(2025), so this constitutes a strong baseline relative to previous work.
+On top of this, our implementation of Prompt-Only appears to significantly improve over that of Zhang et al.
+(2024).
+Additional evaluations and full plots are in Appendix˜F.
 6.1 20 Questions
-We consider three sets of 20 Questions problems: Animals, Celebrities, and Things (See Section˜F.1).
-Each problem set comprises 100 target entities $\{\theta^{*}_{i}\}_{i=1}^{100}$ from a given category.
-The space of possible $\theta$ is large and not explicitly defined or restricted: we do not tell the LLM this set of target entities, so the space of $\theta$ is bounded only by what the LLM can generate.
-We note that by comparison, many previous works have relied on restricted spaces for $\theta$ (Chan et al., 2025; Hu et al., 2024; Piriyakulkij et al., 2023; Wang et al., 2025).
+We consider three sets of 20 Questions problems: Animals, Celebrities, and Things (See Section˜H.1).
+Each problem set comprises 100 target entities, $\{\theta^{*}_{i}\}_{i=1}^{100}$.
+The space of possible $\theta$ is large and not explicitly defined or restricted: we do not tell the LLM this set of target entities, so the space is bounded only by what the LLM can generate; by comparison, many previous works have relied on restricted spaces for $\theta$ (Chan et al., 2025; Hu et al., 2024; Piriyakulkij et al., 2023; Wang et al., 2025).
+Figure 2: Success rate on 20 Questions: mean $\pm$ standard error across 100 targets per dataset.
 Success Rate (%)
-Dataset Model Naive Split BED-LLM Entropy Data–Est. ICL Beliefs Impl. Max.
+Model Prompt-Only Split CoT BED-LLM Entropy Data–Est. ICL Beliefs Impl. Max.
+[…]
+[…]
+[…]
 […]
@@
 This estimator is positively biased and thus conservative.
-Data–Est.
-is not possible to run for GPT models due to limited logprobs support in OpenAI API.
+Method Joint model Objective Belief updates
+BED-LLM Prior–likelihood Full EIG (Eq. 3) Filtered ($p_{f}$)
 […]
-BED-LLM improves over Naive and Split baselines
-Our results in Table˜1 and Figure˜1 show BED-LLM significantly outperforming Naive and Split across all problems and LLMs.
-Particularly notable is that BED-LLM’s final success rate is typically more than double that of Naive, highlighting the big gains that can be achieved by using explicit EIG maximisation instead of implicit LLM reasoning.
-Figure 1: Success rate on 20 Questions: mean $\pm$ standard error across 100 targets per dataset.
+[…]
+[…]
+[…]
+[…]
+[…]
+Table 2: Summary of how each method relates to BED-LLM’s three core algorithmic components. Each ablation (bottom section) modifies exactly one component of BED-LLM, highlighted in grey.
+[…]
+BED-LLM improves over Prompt-Only and Split baselines
+Our results in Figure˜2 and Table˜1 show BED-LLM significantly outperforming both baselines across all problems and LLMs.
+Particularly notable is that BED-LLM’s final success rate is typically more than double that of Prompt-Only, highlighting the substantial gains that can be achieved by using explicit EIG maximization.
 BED-LLM ablations
-In order to understand the importance of BED-LLM’s algorithmic components, we further evaluate four ablations on this problem: Entropy, Data–Estimation, ICL Beliefs and Implicit Maximization.
-Each of these differs from BED-LLM with respect to one algorithmic component.
-Entropy replaces the EIG data-acquisition objective with the marginal predictive entropy $\mathrm{H}[p(y_{t};x_{t},h_{t-1})]$ (Section˜3.2); this contrasts with Split in that it uses $p_{\mathrm{LLM}}(y_{t+1};[\theta,x_{t+1}])$ rather than a deterministic likelihood.
-Implicit Maximization involves sampling candidate questions and prompting the questioner to select the most informative one, without any explicit objective estimation (Section˜3.2).
-ICL Beliefs uses $\theta$ belief updates derived from simple in-context learning, namely $p_{\mathrm{LLM}}(\theta;h_{t-1})$ instead of $p_{f}(\theta;h_{t-1})$, testing the importance of our filtering mechanism (Section˜3.3).
-Data–Estimation uses a data–estimation pairing rather than BED-LLM’s prior–likelihood setup (see Appendix˜D).
+We further evaluate four ablations of BED-LLM to isolate the contribution of each of its core components; full descriptions are in Appendix˜G.
+Entropy replaces the full EIG objective with the marginal predictive entropy, $\mathrm{H}[p(y_{t};x_{t},h_{t-1})]$; it is similar to the Split baseline, but uses BED-LLM’s likelihood model instead of a deterministic one.
+Data–Estimation swaps BED-LLM’s prior–likelihood joint model for a data–estimation pairing (see Appendix˜D), testing the importance of BED-LLM measuring uncertainty in $y$ space instead of $\theta$ space (see Section˜3).
+ICL Beliefs omits our filtering procedure in the belief extraction (see Section˜3.1), just using the, often incoherent, raw in-context beliefs $p_{\mathrm{LLM}}(\theta;h_{t-1})$ instead of $p_{f}(\theta;h_{t-1})$.
+Finally, Implicit Maximization replaces explicit EIG estimation with LLM judgment, drawing on the Tree of Thoughts (ToT; Yao et al., 2023a) intuition of reasoning over hypothetical futures: the model is presented with the same candidate questions as BED-LLM, but is simply prompted to internally reason and select the question it judges most informative, providing a lightweight alternative to explicit EIG estimation.
 In Table˜1 we see that BED-LLM comfortably outperforms all alternative approaches.
-Notably, Entropy provides the strongest ablation baseline, with performance almost identical (sometimes superior) to Split.
-This shows that the use of a non-deterministic likelihood is beneficial in allowing us to target a proper EIG, rather than because it is particularly detrimental to the marginal predictive entropy itself.
+Notably, Entropy provides the strongest alternative, with performance slightly better than Split.
+The fact that Entropy performs much more similarly to Split than BED-LLM shows that the use of a non-deterministic likelihood is beneficial predominantly in allowing us to target a proper EIG, rather than due to changes in the marginal predictive entropy itself.
+The improvement of Implicit Maximization over Prompt-only is also notable, given how cheap it is to run (see Table˜3 for run time information).
 Prior–likelihood outperforms data–estimation
 Our analysis in §4 is validated by our results: BED-LLM’s prior-likelihood approach substantially outperforms Data–Estimation.
-Data–Estimation still outperforms Naive, but interestingly it performs worse than Entropy, highlighting the importance of estimating uncertainty in the $y$ space instead of $\theta$ space.
+Data–Estimation still outperforms Prompt-Only, but interestingly it performs worse than Entropy, highlighting the importance of estimating uncertainty in the $y$ space instead of $\theta$ space here.
 These findings reinforce our claim that the choice of joint-model factorization is a critical algorithmic decision.
 Rejection sampling and explicit EIG maximization are key
-We also see the importance of two other aspects of BED-LLM.
-First, how we produce our beliefs over $\theta$ matters: deriving beliefs using simple in-context learning, as in ICL Beliefs, lead to massive performance drops.
-Second, while BED-LLM’s routines for sampling candidate questions and hypotheses are crucial, they alone are not sufficient: passing the samples to an LLM and prompting it to select the highest-EIG question, as in Implicit Maximization, works much less well than using the samples to explicitly maximize EIG.
+We also see how we produce our beliefs over $\theta$ matters: deriving beliefs using simple in-context learning, as in ICL Beliefs, lead to massive performance drops.
+Further, while BED-LLM’s routines for sampling candidate questions and hypotheses are crucial, they alone are not sufficient: passing the samples to an LLM and prompting it to select the highest-EIG question, as in Implicit Maximization, works much less well.
 BED-LLM is robust to questioner–answerer mismatch
-Our results in Figure˜1 demonstrate that the benefit of BED-LLM persists even under model misspecification.
+Our results in Figure˜2 demonstrate that the benefit of BED-LLM persists even under model misspecification.
 This is important for applicability to real-world users, whose responses will follow a different distribution than the questioner LLM.
@@
 Unlike 20 Questions, in which $\theta$ is a concrete entity and most reasonable questions have clear answers, many real-world information-gathering tasks involve more abstract targets and less predictable data generation.
-A key example is learning user preferences, where it may be difficult to explicitly define a concrete closed set of possible $\theta$, and it is also challenging for the LLM to develop appropriate uncertainty estimates.
+A key example is learning user preferences, where it may be difficult to explicitly define a concrete closed set of possible $\theta$, or for the LLM to develop appropriate uncertainty estimates.
 To study such a scenario, we evaluate BED-LLM on inferring users’ film preferences.
-Here the target we are trying to gather information about is somewhat abstract, and we have some flexibility in how we define $\theta$ in our joint model.
-Our chosen setup is to define $\theta$ to be a user profile, namely a paragraph of text describing the user’s film preferences, with our answerer model prompted to emulate a user with a given profile; see Appendix˜G for full details.
-We consider 200 different user profiles as the ground truth $\theta^{*}$, but as with 20 Questions this set is never given to the questioner.
+Here the target $\theta$ is somewhat abstract, and we have more flexibility in how we define it in our joint model.
+Our chosen setup is to define $\theta$ to be a user profile, namely a paragraph of text describing the user’s film tastes with our answerer model prompted to emulate a user with a given profile; see Appendix˜I for full details.
+We consider 200 ground-truth profiles, $\{\theta^{*}_{i}\}_{i=1}^{200}$, which, as with 20 Questions, are never revealed to the questioner.
 Because Split is not applicable as a baseline here (a deterministic likelihood assumption is clearly unreasonable), we benchmark with the similar Entropy approach instead.
-We also note that data–estimation setup is completely unviable here as well because of the large $\theta$ space.
-At each turn $t\in(0,1,\ldots,5)$ we use the question-answer history as context for generating a list of ten film recommendations.
+We also note that data–estimation setup is completely unviable here due to the large $\theta$ space.
+At each turn, $t\in(0,1,\ldots,5)$, the questioner uses $h_{t-1}$ to generate ten film recommendations.
 This list is then rated in its fit to the user profile using an LLM-as-judge setup (Trivedi et al., 2024; Zhu et al., 2025).
@@
 The films’ scores are not included in $h_{t-1}$.
-Our results in Figure˜2 show that, while Naive is often a strong baseline in this preference-elicitation scenario, BED-LLM is still able to provide a boost over both Naive and Entropy, producing higher-rated film recommendations.
-BED-LLM’s benefit is most clear in scenarios where the questioner belongs to a different model class to the answerer: here Naive’s performance is much less convincing.
+Our results in Figure˜3 show that, while Prompt-Only seems to be a stronger baseline in this preference-elicitation scenario than for 20 Questions, BED-LLM is still able to provide a boost over both it and Entropy, producing higher-rated film recommendations.
+BED-LLM’s benefit is most clear in scenarios where the questioner belongs to a different model class to the answerer.
 Figure 3: Mean rating across 10 film recommendations: mean $\pm$ standard error across 200 users.
@@
 In this work, we have shown how to effectively apply the framework of sequential Bayesian experimental design (BED) to the problem of interactive information gathering with LLMs.
-Specifically, we have introduced BED-LLM, which provides a specific, information–theoretic, sequential BED approach that makes a variety of carefully justified design choices in the joint-model factorization, belief updating, and EIG estimation.
+Specifically, we have introduced BED-LLM, which provides a specific, information-theoretic, sequential BED approach that makes a variety of carefully justified design choices in the joint-model factorization, belief updating, and EIG estimation.
 Particularly central to BED-LLM is the prior–likelihood pairing with filtering of hypotheses for consistency with the history.
@@
 Together, these innovations lead to substantial performance improvements compared to previous approaches.
-The results thus confirm that principled EIG–driven strategies can yield substantial gains for interactive, multi-turn, information gathering problems.
+The results thus confirm that principled EIG-driven strategies can yield substantial gains for interactive, multi-turn information gathering problems.
+Acknowledgments
+DC is supported by the EPSRC CDT in Statistics and Machine Learning (EP/Y034813/1), an Oxford-Radcliffe Scholarship and an Oxford AI Fast Exploration grant.
+FBS is supported by the EPSRC Probabilistic AI Hub (EP/Y028783/1).
+TR is supported by the UK EPSRC grant EP/Y037200/1.
 Appendix contents
-Appendix A Algorithmic considerations
+Appendix A Discussion relating to Section˜3
 A.1 Updating the likelihood
 The success of BED-LLM hinges on our ability to update our joint distribution.
-As mentioned in §3, we choose not to update the likelihood model as more data is gathered, that is, our likelihood in the sequential setting will be $p_{\mathrm{LLM}}\left(y_{t};\left[\theta,x_{t}\right]\right)$ instead of $p_{\mathrm{LLM}}(y_{t+1};[h_{t-1},\theta,x_{t+1}])$.
+As mentioned in §3, we choose not to update the likelihood model as more data is gathered, that is, our likelihood in the sequential setting will be $p_{\mathrm{LLM}}\!\left(y_{t};\left[\theta,x_{t}\right]\right)$ instead of $p_{\mathrm{LLM}}(y_{t};[h_{t-1},\theta,x_{t}])$.
 The main rationale of this choice is that for many problems our beliefs on $\theta$ capture all the required information to predict $y|x$, hence including the history is adding unnecessary context that could influence the LLM’s behavior in undesirable ways.
-However, it is important to note that $p_{\mathrm{LLM}}(y_{t+1};[h_{t-1},\theta,x_{t+1}])$ should be used instead for problems where $\theta$ will not capture all information from previous data, e.g.
+See Section˜F.3 for a results comparison of static and updated likelihoods on the 20 Questions game.
+However, it is important to note that $p_{\mathrm{LLM}}(y_{t};[h_{t-1},\theta,x_{t}])$ should be used instead for problems where $\theta$ will not capture all information from previous data, e.g.
 if $\theta$ is a binary value corresponding to whether we reject a null hypothesis, or is the answer to a particular other question of interest.
@@
 This thus demonstrates how the two criteria can select different questions.
-A.2.1 Predictive entropy is not a good approximation for EIG
+A.2.1 Predictive entropy is not a good approximation of EIG
 As discussed in §3.3, previous information-based query selection mechanisms have assumed that responses are deterministic given $\theta$ and $x$.
@@
 This variability in the expected likelihood entropy can be crucial in selecting good designs.
-Here, we walk through a concrete example where predictive entropy might differ significantly from EIG.
-Figure˜3 shows two candidate questions that could be asked to elicit film preference.
+A concrete example helps highlight how predictive entropy can differ significantly from EIG.
+Figure˜4 shows two candidate questions that could be asked to elicit film preference.
 Question 1 has high predictive entropy: in a randomly selected group of people, we would expect high variation in ice cream preference (regardless of the individual’s film preferences).
@@
 One might be tempted to replace $\hat{p}(y_{t+1};[h_{t-1},x_{t+1}])$ with $p_{\mathrm{LLM}}(y_{t+1};[h_{t-1},x_{t+1}])$ in the EIG estimation in Equation˜3, as the two essentially offer alternative predictive distributions for the outcome.
-We also advise against this though, noting that it again provides no meaningful computational benefits (unless one also assumes a deterministic likelihood, but this would then mean we no longer consider $\theta$ at all).
+We also advise against this, though, noting that it again provides no meaningful computational benefits (unless one also assumes a deterministic likelihood, but this would then mean we no longer consider $\theta$ at all).
 A key reason for avoiding this substitution is that it would mean we are no longer estimating a true EIG: the inconsistency between the likelihood and the marginal data distribution means there is no longer a joint model where we are minimising our expected uncertainty in $\theta$.
@@
 A.3 Prior construction and belief updating
-In Section˜3.3, we argued that naive in-context updating is not sufficient for updating our beliefs: We fail to fully incorporate the information from the history $h_{t}$, and we often have overconfident distributions.
+In Section˜3.1, we argued that Prompt-Only in-context updating is not sufficient for updating our beliefs: We fail to fully incorporate the information from the history $h_{t}$, and we often have overconfident distributions.
 The shortfalls of in-context learning in such settings have also previously be noted by, for example, (Liu et al., 2024b; Zhang et al., 2025; 2024).
@@
 Second, $\theta$ will often represent a user–specific variable that cannot easily be predicted from any data other than the user’s responses to questions: it has been argued that much of the success of in–context learning in LLMs is down to improving problem specification and linking the requested task to data it has seen in its training, rather than truly “learning” from the provided examples (Min et al., 2022; Kossen et al., 2024), but the history in our setting is rarely helpful for this due to its user–specific nature.
-Appendix B Additional discussions on design choices
+Appendix B Discussion relating to Section˜4
 B.1 An alternative view on the faithfulness of conditional distributions
@@
 If, for example, we directly fix the form of $p\!\left(\theta;[h_{t-1},x_{t},y_{t}]\right)$ using our LLM’s predictive distribution, we are also directly relying on its expected entropy being a meaningful measure of design quality.
-If $\theta$ is high–dimensional and predominantly free–form, the resulting entropy produced by the LLM is unlikely to be meaningful and using the data–estimation pairing is unlikely to produce an effective strategy.
+If $\theta$ is high-dimensional and predominantly free-form, the resulting entropy produced by the LLM is unlikely to be meaningful and using the data–estimation pairing is unlikely to produce an effective strategy.
 However, if $y$ is instead quite constrained, the LLM can produce a meaningful entropy over it, and choosing a model based on the prior–likelihood pairing is likely to implicitly define a meaningful distribution, and thus entropy, on $\theta$.
@@
 B.3 Alignment between EIG and belief updating procedure
-Our ultimate goal is to minimize uncertainty in $\theta$, as measured by its entropy.
-With this in mind, we can use the expected uncertainty reduction framework of Bickford Smith et al.
+If our ultimate goal is to minimize uncertainty in $\theta$, as measured by its entropy, then we can use the expected uncertainty reduction framework of Bickford Smith et al.
 (2025) to provide insights into how well our EIG formulation and belief updating procedures align.
@@
 (2025), we can think of the “true” optimal design as selecting
-$x^{*}_{\mathrm{true}}=\mathrm{argmin}_{x}\mathbb{E}_{p_{\mathrm{true}}(y;x)}\left[\mathrm{H}[p(\theta;x,y)]\right],$ (4)
+$x^{*}_{\mathrm{true}}=\mathop{\mathrm{arg\,min}}_{x}\mathbb{E}_{p_{\mathrm{true}}(y;x)}\left[\mathrm{H}[p(\theta;x,y)]\right],$ (4)
 where $p_{\mathrm{true}}(y;x)$ is the true response distribution and $p(\theta;x,y)$ is our belief state after the experiment.
@@
 Thus, we can now see that our choice of joint model corresponds to different choices for approximating these quantities.
-Assuming that the LLM distribution is used directly for the conditional as per Section˜4, we thus have that:
-  • The prior-likelihood pairing equates to the approximations $p_{\mathrm{true}}(y;x)\approx\int p(\theta)p_{\mathrm{LLM}}(y;[\theta,x])d\theta$ and $p(\theta;x,y)\approx p(\theta)p_{\mathrm{LLM}}(y;[\theta,x])/\int p(\theta)p_{\mathrm{LLM}}(y;[\theta,x])d\theta$;
-  • The data–estimation pairing equates to directly specifying a model for $p_{\mathrm{true}}(y;x)$ and then using the approximation $p(\theta;x,y)\approx p_{\mathrm{LLM}}(\theta;[x,y])$.
+Assuming that the LLM distribution is used directly for the conditional as per Section˜4, we thus have that
+  • the prior–likelihood pairing corresponds to using the approximations $p_{\mathrm{true}}(y;x)\approx\int p(\theta)p_{\mathrm{LLM}}(y;[\theta,x])d\theta$ and $p(\theta;x,y)\approx p(\theta)p_{\mathrm{LLM}}(y;[\theta,x])/\int p(\theta)p_{\mathrm{LLM}}(y;[\theta,x])d\theta$; and
+  • the data–estimation pairing corresponds to directly specifying a model for $p_{\mathrm{true}}(y;x)$ and then using the approximation $p(\theta;x,y)\approx p_{\mathrm{LLM}}(\theta;[x,y])$.
 The appropriateness of each of these options, therefore, comes down to how faithful these approximations are respectively to the true data distribution, $p_{\mathrm{true}}(y;x)$, and how we actually derive our belief distribution on $\theta$ in practice once we have seen the new data.
@@
 However, we do know upfront how we plan to derive our belief distribution on $\theta$ in practice, so we can use this to guide which joint model we formulate our EIG from.
-Namely, we observe that: a) using the prior–likelihood EIG pairing equates to assuming we will make a Bayesian update to our beliefs on $\theta$ using the likelihood $p_{\mathrm{LLM}}(y;[\theta,x])$; b) using the data–estimation EIG pairing equates to assuming we will make an in-context update to our beliefs on $\theta$, as we are treating $p(\theta;x,y)$ as $p_{\mathrm{LLM}}(\theta;[x,y])$.
+Namely, we observe that (a) using the prior–likelihood EIG pairing equates to assuming we will make a Bayesian update to our beliefs on $\theta$ using the likelihood $p_{\mathrm{LLM}}(y;[\theta,x])$, and (b) using the data–estimation EIG pairing equates to assuming we will make an in-context update to our beliefs on $\theta$, as we are treating $p(\theta;x,y)$ as $p_{\mathrm{LLM}}(\theta;[x,y])$.
 Our preference between the pairings should therefore be guided in part by how we plan to update the model in practice.
-In particular, if we plan to make pure Bayesian updates, then the prior-likelihood formulation will tend to yield an EIG that is more faithful to our updating procedure, while if we only make simple in-context updates, the data–estimation formulation will tend to yield a more faithful EIG instead.
-The update we use in practice, namely taking $p(\theta;x,y)=p_{f}(\theta;[x,y])$ as outlined in Section˜3.3, can be seen as being somewhere between the in-context and Bayesian updating: we initially sample from $p_{\mathrm{LLM}}(\theta;[x,y])$, but then perform filtering and other steps.
+In particular, if we plan to make pure Bayesian updates, then the prior–likelihood formulation will tend to yield an EIG that is more faithful to our updating procedure, while if we only make simple in-context updates, the data–estimation formulation will tend to yield a more faithful EIG instead.
+The update we use in practice, namely taking $p(\theta;x,y)=p_{f}(\theta;[x,y])$ as outlined in Section˜3.1, can be seen as being somewhere between the in-context and Bayesian updating: we initially sample from $p_{\mathrm{LLM}}(\theta;[x,y])$, but then perform filtering and other steps.
 The relative extent to which it resembles each will be problem–dependent and again be linked to how much we trust the LLM to capture uncertainty in the space of $\theta$ vs.
 $y$.
-For the settings we consider, we expect $p_{f}(\theta;[x,y])$ to generally be better approximated by a Bayesian update than an in–context update, aligning with our decision to use the prior–likelihood formulation.
+For the settings we consider, we expect $p_{f}(\theta;[x,y])$ to generally be better approximated by a Bayesian update than an in-context update, aligning with our decision to use the prior–likelihood formulation.
 The reasons for this are that a) the filtering often removes a large proportion of the generated samples, especially at later experiment turns, with $p_{\mathrm{LLM}}(\theta;h_{t-1})$ not fully incorporating information from the history; b) the maintaining of the set of one consistent hypotheses from one turn to the next encourages a more Bayesian behavior, with samples persisting unless contradicted by a new likelihood term; and c) the typical premature overconfidence of $p_{\mathrm{LLM}}(\theta;h_{t-1})$ to a small number of hypotheses means it is typically unrepresentative of our beliefs.
@@
 Cooper et al.
-(2025) compute posterior entropy over a working set of top-k hypotheses (without filtering) through heuristic pruning.
+(2025) compute posterior entropy over a working set of top $k$ hypotheses (without filtering) through heuristic pruning.
 Wang et al.
@@
 Our findings are complementary: by providing more faithful model factorizations, belief updates, and EIG estimators in the LLM setting, BED-LLM could supply stronger building blocks for policy-based methods, reducing variance, enhancing effectiveness, and improving the sample efficiency of policy training.
-Algorithm 1 Data–Estimation Selection at Turn $t$
-1:History $h_{t-1}$; candidate questions $\mathcal{X}_{\text{cand}}^{t}$; answer sets $\{\mathcal{Y}(x)\}$
-2:Selected question $x_{t}^{\star}$
-3:for each $x\in\mathcal{X}_{\text{cand}}^{t}$ do
-4:  Obtain predictive answer distribution $p_{\text{LLM}}(y;x,h_{t-1})$ for all $y\in\mathcal{Y}(x)$
-5:  for each $y\in\mathcal{Y}(x)$ do
-6:   Compute entropy $H_{y}\leftarrow\mathrm{H}\!\left[p_{f}(\theta;h_{t-1}\cup(x,y))\right]$
-7:  end for
-8:  Compute $\text{EIG}(x)\leftarrow-\sum_{y\in\mathcal{Y}(x)}p_{\text{LLM}}(y;x,h_{t-1})\;H_{y}$
+Algorithm 1 Data–Estimation sequential information gathering
+1:LLM $p$, history $h_{0}=\varnothing$, budget $T$, num. hypothesis samples $M$
+2:for $t=1,\ldots,T$ do
+3:  Generate candidates: sample $\mathcal{X}_{\mathrm{cand}}=\{x_{t}^{(1)},\ldots,x_{t}^{(K)}\}$ from $p(x_{t};h_{t-1})$
+4:  for each candidate $x_{t}\in\mathcal{X}_{\mathrm{cand}}$ do
+5:   for each $y\in\mathcal{Y}$ do // enumerate answer options
+6:     $w_{y}\leftarrow p(y;[h_{t-1},x_{t}])$ // answer prob. via LLM logits
+7:     Sample $\{\theta_{m}^{(y)}\}_{m=1}^{M}\sim p(\theta;[h_{t-1},x_{t},y])$
+8:     Estimate $\hat{H}_{y}\leftarrow\widehat{\mathrm{H}}\!\left[p(\theta;[h_{t-1},x_{t},y])\right]$ using $\{\theta_{m}^{(y)}\}$
 9:   end for
-10:Select $x_{t}^{\star}\leftarrow\arg\max_{x\in\mathcal{X}_{\text{cand}}^{t}}\text{EIG}(x)$
-11:return $x_{t}^{\star}$
+10:   $\hat{H}_{\mathrm{cond}}\leftarrow\sum_{y\in\mathcal{Y}}w_{y}\,\hat{H}_{y}$ // Term 1: marginal entropy of $\theta$ (design-dependent)
+11:    Pool: $\{\theta_{m}\}_{m=1}^{M^{\prime}}\leftarrow\bigcup_{y\in\mathcal{Y}}\{\theta_{m}^{(y)}\}$ // samples from mixture
+12:   for $m=1,\ldots,M^{\prime}$ do
+13:     $\ell_{m}\leftarrow\log\sum_{y\in\mathcal{Y}}w_{y}\cdot p(\theta_{m};[h_{t-1},x_{t},y])$
+14:   end for
+15:    $\hat{H}_{\mathrm{marg}}\leftarrow-\frac{1}{M^{\prime}}\sum_{m=1}^{M^{\prime}}\ell_{m}$
+16:   $\widehat{\mathrm{EIG}}(x_{t})\leftarrow\hat{H}_{\mathrm{marg}}-\hat{H}_{\mathrm{cond}}$
+17:  end for
+18:  Select: $x_{t}^{*}\leftarrow\arg\max_{x_{t}\in\mathcal{X}_{\mathrm{cand}}}\widehat{\mathrm{EIG}}(x_{t})$
+19:  Ask and observe: pose $x_{t}^{*}$, receive $y_{t}$, set $h_{t}\leftarrow(h_{t-1},(x_{t}^{*},y_{t}))$
+20:end for
 Appendix D Data–estimation method
 Our Data–Estimation method is based on a model derived from a data–estimation pairing (Section˜3).
+$p(\theta,y_{t};h_{t-1},x_{t})=p\!\left(y_{t};[h_{t-1},x_{t}]\right)p\!\left(\theta;[h_{t-1},x_{t},y_{t}]\right).$ (5)
 D.1 EIG estimation
-Suppose our model is given by $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}p\left(y_{t};[h_{t-1},x_{t}]\right)p\left(\theta;[h_{t-1},x_{t},y_{t}]\right)$}.
-Here, it will clearly be beneficial to directly use Equation˜1 for estimating the EIG, as here we directly have access to all the required terms, other than $p(\theta)$ which can be simply ignored as it is not a function of $x$ so does not affect optimization of the question.
-If the possible values for $y$ are enumerable and we can evaluate $p\left(y_{t};[h_{t-1},x_{t}]\right)$ in closed–form, we can directly calculate the exact EIG (up to a constant) without requiring any estimation at all:
-$\text{EIG}_{\theta}(x)-\mathrm{Const}=-\sum_{y}p\left(y_{t};[h_{t-1},x_{t}]\right)\mathrm{H}\left[p\left(\theta;[h_{t-1},x_{t},y_{t}]\right)\right],$ (5)
-where the entropy $\mathrm{H}\left[p\left(\theta;[h_{t-1},x_{t},y_{t}]\right)\right]$ can be evaluated directly from the logits of the LLM, or if these are not available, estimated by sampling.
-If we cannot enumerate $y$ or evaluate $p\left(y_{t};[h_{t-1},x_{t}]\right)$, we can simply instead resort to Monte Carlo and use the estimator:
-$\text{EIG}_{\theta}(x)-\mathrm{Const}\approx-\frac{1}{N}\sum_{n=1}^{N}\mathrm{H}\left[p(\theta;[c,x,y_{n}])\right]\quad\mathrm{where}\quad y_{n}\sim p\left(y_{t};[h_{t-1},x_{t}]\right).$ (6)
-We provide an overview of how to do this in the LLM setting in Algorithm˜1.
-D.2 Generating candidate hypotheses
-To generate candidate values of $\theta$ for the data–estimation method, we use the prompt in Section˜D.2.
+Using the form of the EIG given by Equation˜1, we have
+$\mathrm{EIG}_{\theta}(x)=\underbrace{\mathrm{H}\!\left[p(\theta;h_{t-1},x)\right]}_{\text{marginal entropy of }\theta}\;-\;\underbrace{\mathbb{E}_{p\!\left(y_{t};[h_{t-1},x_{t}]\right)}\!\left[\mathrm{H}\left[p\!\left(\theta;[h_{t-1},x_{t},y_{t}]\right)\right]\right]}_{\text{expected posterior entropy}}.$ (6)
+Crucially, the marginal on $\theta$ implied by this pairing is obtained by integrating out $y$:
+$p(\theta;h_{t-1},x)=\sum_{y_{t}}p\!\left(y_{t};[h_{t-1},x_{t}]\right)p\!\left(\theta;[h_{t-1},x_{t},y_{t}]\right),$ (7)
+which is a mixture distribution whose components and weights both depend on $x$.
+Unlike in the prior–likelihood pairing, where the belief state $p_{f}(\theta;h_{t-1})$ is constructed independently of the candidate design, the marginal entropy $\mathrm{H}[p(\theta;h_{t-1},x)]$ here varies across candidate questions and cannot be dropped from the optimization.
+This is a concrete manifestation of the issue identified in Section˜3: the data–estimation pairing does not provide a design–independent belief state on $\theta$.
+Discrete answer space.
+When the possible values for $y$ are discrete (and finite) and we can evaluate $p\!\left(y_{t};[h_{t-1},x_{t}]\right)$ in closed form, both terms of Equation˜6 can be computed directly.
+The second term (expected posterior entropy) is:
+$\mathbb{E}_{p\!\left(y_{t};[h_{t-1},x_{t}]\right)}\!\left[\mathrm{H}\!\left[p\!\left(\theta;[h_{t-1},x_{t},y_{t}]\right)\right]\right]=\sum_{y_{t}}p\!\left(y_{t};[h_{t-1},x_{t}]\right)\mathrm{H}\!\left[p\!\left(\theta;[h_{t-1},x_{t},y_{t}]\right)\right],$ (8)
+where each posterior entropy $\mathrm{H}[p\!\left(\theta;[h_{t-1},x_{t},y_{t}]\right)]$ can be evaluated from the LLM’s logits or, if these are unavailable, estimated by sampling.
+The first term (marginal entropy) requires evaluating $p(\theta;h_{t-1},x)$ as the mixture in Equation˜7, then computing its entropy.
+In practice, this entropy cannot be obtained in closed form and must be estimated.
+A natural approach is to draw samples $\theta_{m}\sim p(\theta;h_{t-1},x)$ (by first sampling $y_{m}\sim p\!\left(y_{t};[h_{t-1},x_{t}]\right)$, then $\theta_{m}\sim p\!\left(\theta;[h_{t-1},x_{t},y_{t}]\right)[,y_{m}]$) and using a plug-in or nearest-neighbor entropy estimator, or evaluating $\log p(\theta_{m};h_{t-1},x)=\log\sum_{y}p\!\left(y_{t};[h_{t-1},x_{t}]\right)\,p(\theta_{m};[h_{t-1},x,y])$ exactly by enumerating over $y$.
+Monte Carlo estimator.
+When $y$ cannot be enumerated, both terms must be estimated by Monte Carlo. Drawing $N$ samples $y_{n}\sim p\!\left(y_{t};[h_{t-1},x_{t}]\right)$, the full estimator takes the form:
+$\mathrm{EIG}_{\theta}(x;h_{t-1})\approx\;\mathrm{H}\!\left[\frac{1}{N}\sum_{n=1}^{N}p(\theta;[h_{t-1},x,y_{n}])\right]\;-\;\frac{1}{N}\sum_{n=1}^{N}\mathrm{H}\!\left[p(\theta;[h_{t-1},x,y_{n}])\right].$ (9)
+Whereas Equation˜3 performs this computation over the answer space $\mathcal{Y}$ (which is small and enumerable for multiple-choice questions), Equation˜9 requires it over the hypothesis space $\Theta$.
+For the first term, evaluating the mixture entropy requires computing $p(\theta;[h_{t-1},x,y_{n}])$ for a set of sampled $\theta$ values across all $N$ mixture components — an $O(M\times N)$ cost in autoregressive probability evaluations over the full hypothesis text.
+D.2 Implementation
+We provide an overview of the Data–Estimation procedure in Algorithm˜1.
+In our experiments, the answer space $\mathcal{Y}$ is enumerable (multiple-choice options), so we use Equation˜8 for the second term.
+For the first term, we estimate the marginal entropy by sampling $\theta$ values from the mixture in Equation˜7 and evaluating $\log p(\theta;h_{t-1},x)$ by enumerating over $y$; that is, we use:
+$\mathrm{H}\!\left[p(\theta;h_{t-1},x)\right]\approx-\frac{1}{M}\sum_{m=1}^{M}\log\!\left(\sum_{y}p\!\left(y_{t};[h_{t-1},x_{t}]\right)\,p(\theta_{m};[h_{t-1},x,y])\right),$ (10)
+where $\theta_{m}$ are drawn from $p(\theta;h_{t-1},x)$.
+This requires $M\times|\mathcal{Y}|$ evaluations of $p(\theta_{m};[h_{t-1},x,y])$ per candidate question, each involving a full forward pass over the hypothesis text.
+D.3 Generating candidate hypotheses
+To generate candidate values of $\theta$ for the data–estimation method, we use the prompt in Section˜D.3.
+Figure˜5 shows an example of the distribution of samples obtained following two rounds in the 20–questions game.
+Note that the samples are highly concentrated on just a handful of answers.
+This lack of diversity shows that the model’s belief distribution is far more concentrated relative to the variability over valid hypotheses in the ground–truth task distribution, which negatively impacts the performance of the data–estimation method.
+Conversation History ($\mathcal{D}$) Q: Is this person known for their contributions to science?   A: No. Q: Is this person known for their contributions to the arts?    A: Yes.
+Sampled Hypothesis Distribution
+Vincent van Gogh 93
+Salvador Dali 44
+Frida Kahlo 37
+Georgia O’Keeffe 10
+August Wilson 8
+Auguste Rodin 8
+Figure 5: An empirical hypothesis distribution generated by the LLM, conditioned on the conversation history (top).
+By independently sampling $N=200$ hypotheses, we observe significant mode collapse; the probability mass is heavily concentrated on a few prominent figures (e.g., Vincent van Gogh).
+This overconfidence negatively impacts the performance of the data-estimation baseline.
+Note that this aggregated view is purely diagnostic; Algorithm˜1 operates directly on the probabilities of individual samples.
 Figure 6: Prompt for generating hypotheses (and evaluating their probability) for the data–estimation method.
@@
 - No extra words, explanations, numbering, or punctuation beyond what’s in the name itself (hyphens/apostrophes allowed if part of the name).
-Figure˜5 shows an example of the distribution of samples obtained following two rounds in the 20–questions game.
-Note that the samples are highly concentrated on just a handful of answers.
-This lack of diversity shows that the model’s belief distribution is far more concentrated relative to the variability over valid hypotheses in the ground–truth task distribution, which negatively impacts the performance of the data–estimation method.
-Is this person known for their contributions to science?
-No.
-Is this person known for their contributions to the arts?
-Yes.
-  "Vincent van Gogh": 93,
-  "Salvador Dali": 44,
-  "Frida Kahlo": 37,
-  "Georgia O Keeffe": 10,
-  "August Wilson": 8,
-  "Auguste Rodin": 8
-Figure 5: An example of the sample distribution generated using the prompt in Section˜D.2, conditioned on the two question/answer pairs at the top of this figure.
-At this stage of the game, we independently sample 200 hypotheses from the LLM and record their frequencies.
-Note that the distribution exhibits strong mode collapse, with most of the mass highly concentrated on just a few answers, which negatively impacts the performance of the data–estimation method.
-This summary is for diagnostic purposes: Algorithm˜1 operates on the probabilities of individual samples and never instantiates such an aggregated summary.
+Section˜D.3 shows an example of in-context belief updating failing to respect the conversation history.
+The response to the first question indicates that the target is male; however, after the second question, the LLM generates two female candidate hypotheses.
+Figure 7: An example of in-context belief tracking with GPT-4o-mini proposing hypotheses inconsistent with the history. Q1: Is the person male?
+Answerer: Yes.
+Q2: Is this person often associated with civil rights or social justice?
+Answerer: Yes.
+Sampled hypotheses: [’James Baldwin’, ’A.
+Philip Randolph’, ’Angela Davis’, ’Malcolm X’, ’W.E.B.
+Du Bois’, ’Desmond Tutu’, ’Cesar Chavez’, ’Rosa Parks’, ’Martin Luther King Jr.’, ’Langston Hughes’, ’John Lewis’, ’Frederick Douglass’, ’Nelson Mandela’, ’Thurgood Marshall’, ’Bayard Rustin’]
+Hypotheses rejected with filtering: [’Angela Davis’, ’Rosa Parks’]
 Appendix E Generating candidate hypotheses for BED-LLM
@@
 Below, we detail the steps we take to construct our distribution over hypotheses.
-Candidate hypotheses are generated jointly, rather than independently.
+Candidate hypotheses are generated jointly, rather than independently
 […]
@@
 We repeat the generation process, keeping the previously generated and filtered samples in context to elicit new generations, either twice or three times if sufficient hypotheses have not been generated (noting the number of possible valid samples can be less than the number requested).
-We assume a uniform distribution over hypotheses
+We define a uniform prior over hypotheses
 While one could in principle reweight candidates using importance sampling, in practice we choose to not rely on the model’s internal probabilities.
-Instead we approximate the prior as a uniform distribution over this union
+Instead we define the prior as a uniform distribution:
 $p_{f}(\theta;h_{t})\approx\frac{1}{|\Theta|}\sum_{\theta\in\Theta}\delta_{\theta}.$
@@
 For fairness, in our experiments we have kept these parameters constant across models.
-Appendix F Experiment details for 20 Questions
-F.1 Problem sets
+Appendix F Additional experimental results
+F.1 20 Questions and Preference Elicitation
+Figure 9: Success rate on 20 Questions with Qwen2.5-72B, extended to 50 turns: mean $\pm$ standard error across 100 targets per dataset.
+BED-LLM remains the best-performing method at every turn.
+The performance gap to baselines narrows in later turns as the task saturates — most methods eventually solve the easier instances — but BED-LLM consistently retains its advantage.
+Figure 10: Analysis of the empirical convergence of the Rao-Blackwellized EIG estimator using GPT-4o-mini.
+We plot EIG estimates for 5 candidate questions at turn 5 of the 20 Questions game with $\theta^{*}=$“Saoirse Ronan”, as a function of the number of samples in the hypothesis set $\Theta$.
+Figure˜10 shows the EIG estimates for five candidate questions as the number of hypothesis samples $|\Theta|$ increases, at turn 5 of a representative 20 Questions game.
+The estimates converge rapidly with the number of samples, and more importantly, the ranking of candidate questions stabilizes well before the estimates themselves have fully converged.
+By $|\Theta|=10$ the relative ordering of all five candidates already matches the ordering observed at $|\Theta|=85$.
+This means that the question selection decision—which is what matters algorithmically, since we only need the $\arg\max$, not the absolute EIG value—is robust to using a small hypothesis set.
+This justifies BED-LLM’s use of a moderate sample budget without sacrificing decision quality.
+Figure 11: Success rate on 20 Questions. Mean $\pm$ standard error across 100 targets. BED-LLM beats all other methods across all datasets and models evaluated.
+Figure 12: Full plots for 20 Questions Experiments with different questioner and answerer.
+Figure 13: Mean rating across 10 film recommendations: mean $\pm$ standard error across 200 users. BED-LLM beats all other methods across all datasets and models evaluated.
+F.2 Wall-clock times
+Method Runtime
+Data–Estimation 1d 17h 13m 34s
+ICL Beliefs 3h 10m 39s
+Entropy 2h 46m 19s
+Split 2h 30m 35s
+BED-LLM 2h 28m 43s
+Implicit Maximization 45m 10s
+CoT 35m 28s
+Prompt-Only 26m 29s
+Table 3: Wall-clock runtimes for 20 Questions for Qwen2.5-72B on the entire Animals problem set.
+Table˜3 shows that Data–Estimation is over an order of magnitude slower than all other methods: its full-sequence likelihood evaluations per candidate question dominate the cost.
+Among the remaining methods, BED-LLM’s runtime is comparable to—and even slightly lower than—Entropy and Split, despite computing the full EIG; this is because the additional conditional-entropy term in Equation˜3 reuses the same likelihood evaluations and adds negligible overhead.
+Prompt-Only and CoT are the cheapest methods, as they require only a single LLM generation call per turn with no hypothesis sampling or objective computation.
+Implicit Maximization falls in between: it samples hypotheses and candidate questions but replaces explicit EIG estimation with a single LLM selection call.
+ICL Beliefs is slightly more expensive than BED-LLM because the absence of filtering leads to lower-quality hypotheses, requiring a much larger number of questions before the game terminates.
+F.3 Effect of sequentially updating the BED-LLM likelihood
+As discussed in Section˜F.3, we choose to use a static likelihood, $p_{\mathrm{LLM}}\!\left(y_{t};\left[\theta,x_{t}\right]\right)$, in the model factorization for BED-LLM.
+Alternatively, we could condition the likelihood on the full interaction history, i.e.
+$p_{\mathrm{LLM}}(y_{t+1};[h_{t-1},\theta,x_{t+1}])$.
+However, this can lead to undesirable effects (e.g.
+context-induced calibration shifts).
+We provide results in Table˜4 for the 20 Questions game where we update the likelihood, and demonstrate that it leads to consistently worse performance than BED-LLM, supporting our rationale in Section˜F.3 to keep the likelihood static.
+Success Rate (%)
+Dataset Model BED-LLM + static likelihood BED-LLM + likelihood updating
+Animals GPT-4o-mini 88 $\pm$ 3.3 85 $\pm$ 3.6
+GPT-4o 93 $\pm$ 2.6 86 $\pm$ 3.5
+Llama-3.1-8B 63 $\pm$ 4.9 67 $\pm$ 4.7
+Llama-3.3-70B 79 $\pm$ 4.1 80 $\pm$ 4.0
+Qwen2.5-72B 95 $\pm$ 2.2 89 $\pm$ 3.1
+Celebrities GPT-4o-mini 72 $\pm$ 4.5 63 $\pm$ 4.9
+GPT-4o 86 $\pm$ 3.5 83 $\pm$ 3.8
+Llama-3.1-8B 58 $\pm$ 5.0 56 $\pm$ 5.0
+Llama-3.3-70B 55 $\pm$ 5.0 44 $\pm$ 5.0
+Qwen2.5-72B 84 $\pm$ 3.7 70 $\pm$ 4.6
+Things GPT-4o-mini 49 $\pm$ 5.0 43 $\pm$ 5.0
+GPT-4o 64 $\pm$ 4.8 57 $\pm$ 5.0
+Llama-3.1-8B 26 $\pm$ 4.4 24 $\pm$ 4.3
+Llama-3.3-70B 55 $\pm$ 5.0 54 $\pm$ 5.0
+Qwen2.5-72B 62 $\pm$ 4.9 61 $\pm$ 4.9
+Table 4: BED-LLM success rates across datasets and models.
+$\pm$ numbers show the standard error of the mean estimated using $\sqrt{p(1-p)/(n-1)}$ where $p$ is the success percentage and $n$ is the number of datapoints.
+This estimator is positively biased and thus conservative.
+Appendix G Further description of methods and ablations
+Prompt-Only: A direct prompting baseline where the model is asked to immediately generate the next question given the dialogue history, no candidate generation, and no belief modeling (Appendix˜G).
+This corresponds to the simplest and most common way users currently interact with LLM agents for question-asking tasks, and therefore serves as a natural “standard LLM prompting” baseline.
+After the LLM has generated its question, it waits for the user response and generates a new question based on the new history.
+Figure 14: Prompt for the Prompt-Only method.
+Your task is to ask a series of questions to deduce the identity of the ’person’ if task == ’celebrities’ else ’animal’ if task == ’animals’ else ’entity’ that I’m thinking of with as few queries as possible.
+Only ask questions that can be answered by ’yes’, ’no’ or ’maybe’.
+Do not ask for hint. Make your question standalone with no linebreaker.
+Output format: <question>Your question here?</question>
+Now start asking a question.
+CoT: A direct prompting baseline where the model is asked to reason and provide action (Appendix˜G) loosely following the ReAct framework (Yao et al., 2023b).
+Figure 15: Prompt for the CoT method.
+Your task is to ask a series of questions to deduce the identity of the ’person’ if task == ’celebrities’ else ’animal’ if task == ’animals’ else ’entity’ that I’m thinking of with as few queries as possible.
+Only ask questions that can be answered by ’yes’, ’no’ or ’maybe’.
+Do not ask for hints. The actual question you ask must be standalone with no linebreakers.
+Use the following format for every response:
+Thought: <brief reasoning about what to ask next>
+Action: ASK["<a single yes/no/maybe question with no linebreaks>"]
+Now start by producing your first Thought and Action.
+Split: Essentially a version of our method in which we cast previous approaches, such as Kobalczyk et al.
+(2025): uses the same candidate question generating method as BED-LLM, assumes a deterministic likelihood over hypotheses (i.e., each hypothesis deterministically predicts a single answer to each question), and Evaluates the question on the sampled hypotheses, and selects the question that maximally balances the hypotheses, i.e., splits the current hypothesis set into subsets whose sizes are as close to equal as possible.
+Implicit Maximization (IM).
+The Implicit Maximization baseline is a lightweight, reasoning-driven method inspired by the Tree-of-Thoughts (ToT, Yao et al., 2023a) framework.
+ToT methods explicitly expand a search tree of intermediate reasoning steps (“thoughts”) before selecting an action.
+Implicit Maximization provides a computationally efficient, collapsed variant of this idea:
+  • Rather than explicitly branching over thoughts, the model is prompted to internally deliberate over several possible next questions and their consequences.
+  • The branching and evaluation occur inside the model’s chain-of-thought, rather than through externally enumerated tree expansion.
+  • The model then outputs the question it judges to be most informative, performing an amortized, ToT-style lookahead within a single LLM call.
+Thus, IM captures the central intuition of Tree-of-Thoughts—reasoning over hypothetical futures—while avoiding the substantial computational cost of explicit tree search.
+It serves as a strong inference-time reasoning comparator to BED-LLM.
+Appendix H Experiment details for 20 Questions
+H.1 Problem sets
 We evaluate across three distinct problem sets—Animals, Celebrities, or Things—with each containing a mix of 100 obscure and common targets.
@@
 (2024); it covers a wide range of categories, from plants and clothing to professions, events, and mythical creatures.
-To create the Animals problem set, we prompted OpenAI o3 (OpenAI, 2025, o3-2025-04-16) to generate a list of animals, using the prompt in Section˜F.1.
-The resulting list is shown in Figure˜8.
+To create the Animals problem set, we prompted OpenAI o3 (OpenAI, 2025, o3-2025-04-16) to generate a list of animals, using the prompt in Section˜H.1.
+The resulting list is shown in Figure˜17.
 Alternative names (after $|$ character) were manually added.
@@
 Figure 17: Animals problem set (generated using OpenAI o3, with manual curation)
-F.2 Evaluation
+H.2 Evaluation
 […]
-Figure 9: Full plots for 20 Questions Experiments.
-F.3 Algorithmic details
-Using our sample–then–filter process (see Section˜3.3), we aim to sample at least $N=15$ hypotheses, repeating the cycle up to three times if needed (the exact number of hypotheses can be less than this as it may not be possible to generate sufficient valid hypotheses, especially in later experiment turns).
-The questioner generates $M=15$ candidate questions to test, $\mathcal{X}^{\text{cand}}$, using the “conditional generation” approach of Section˜3.1 when possible, but falling back on “unconditional generation” if insufficient candidate hypotheses have been generated.
-Appendix G Experiment details for preference elicitation
-G.1 Problems
-To generate a set of ground-truth user profiles, we take a set of $200$ real user ratings from the MovieLens-100K dataset (Harper & Konstan, 2015), then use an “oracle” LLM (namely, OpenAI’s o3 model) to produce a paragraph of text that is consistent with each distinct set of ratings, using the prompt in Section˜G.3.
+H.3 Algorithmic details
+Using our sample–then–filter process (see Section˜3.1), we aim to sample at least $N=15$ hypotheses, repeating the cycle up to three times if needed (the exact number of hypotheses can be less than this as it may not be possible to generate sufficient valid hypotheses, especially in later experiment turns).
+The questioner generates $M=15$ candidate questions to test, $\mathcal{X}^{\text{cand}}$, using the “conditional generation” approach of Section˜3.2 when possible, but falling back on “unconditional generation” if insufficient candidate hypotheses have been generated.
+Appendix I Experiment details for preference elicitation
+I.1 Problems
+To generate a set of ground-truth user profiles, we take a set of $200$ real user ratings from the MovieLens-100K dataset (Harper & Konstan, 2015), then use an “oracle” LLM (namely, OpenAI’s o3 model) to produce a paragraph of text that is consistent with each distinct set of ratings, using the prompt in Section˜I.3.
 As was the case for the 20 Questions problems, this problem set is never provided to the questioner and the set of allowed $\theta$ is not constrained.
@@
 Option E is further constrained to always be “none of the above” so that the answerer is not committed to choosing one of the directly generated choices if none are suitable.
-G.2 Evaluation
+I.2 Evaluation
 To again allow tracking of progress through the experiment, after each turn of the interaction $t$, the questioner generates ten film recommendations, conditioned on $h_{t-1}$.
@@
 We report the mean rating and standard error across 200 users, over 5 question–answer turns.
-G.3 Algorithmic details
+I.3 Algorithmic details
 For BED-LLM and Entropy, we compare $M=8$ candidate questions at each turn and we aim to generate at least $N=5$ candidate hypotheses.
-We use the “unconstrained generation” approach of candidate question generation (see Section˜3.1) as the user profiles can be quite diffuse and we are only generating a small number of possible hypotheses that can be quite easy to split.
+We use the “unconstrained generation” approach of candidate question generation (see Section˜3.2) as the user profiles can be quite diffuse and we are only generating a small number of possible hypotheses that can be quite easy to split.
 We note that data–estimation setup is not at all viable for this problem because the large number of tokens and varying dimensionality of each $\theta$ sample mean that $\mathrm{H}[p_{\mathrm{LLM}}(\theta;[h_{t-1},x_{t+1},y_{t+1}])]$ is not only infeasible to estimate, but also is not a meaningful measure of uncertainty.
@@
 Proceed carefully, reasoning explicitly about the user’s overall rating patterns rather than relying exclusively on extreme ratings, to form a comprehensive, stable, and representative film preference profile.
+Appendix J Code